### v2

# Inference-Time Alignment for Diffusion Models via Variationally Stable Doob’s Matching

Abstract
Inference-time alignment for diffusion models aims to adapt a pre-trained reference diffusion model toward a target distribution without retraining the reference score network, thereby preserving the generative capacity of the reference model while enforcing desired properties at the inference time. A central mechanism for achieving such alignment is guidance, which modifies the sampling dynamics through an additional drift term. In this work, we introduce variationally stable Doob’s matching, a novel framework for provable guidance estimation grounded in Doob’s $h$-transform. Our approach formulates guidance as the gradient of logarithm of an underlying Doob’s $h$-function and employs gradient-regularized regression to simultaneously estimate both the $h$-function and its gradient, resulting in a consistent estimator of the guidance. Theoretically, we establish non-asymptotic convergence rates for the estimated guidance. Moreover, we analyze the resulting controllable diffusion processes and prove non-asymptotic convergence guarantees for the generated distributions in the 2-Wasserstein distance. Finally, we show that variationally stable guidance estimators are adaptive to unknown low dimensionality, effectively mitigating the curse of dimensionality under low-dimensional subspace assumptions.
Keywords: Controllable generative learning, inference-time alignment, Doob’s $h$-transform, convergence rate
1  Introduction
Diffusion models (Sohl-Dickstein et al., 2015; Ho et al., 2020; Song and Ermon, 2019; Song et al., 2021) have emerged as powerful generative tools for sampling from data distributions, achieving remarkable success across diverse domains, including text-to-image and text-to-video generation (Ramesh et al., 2021), Bayesian inverse problems (Chung et al., 2023; Song et al., 2023; Chen et al., 2025; Chang et al., 2025b), and scientific applications (Bao et al., 2024; Li et al., 2025; Si and Chen, 2025; Ding et al., 2024; Uehara et al., 2025a). Recent years have witnessed the development of large-scale diffusion models pre-trained on vast datasets. Despite the robust capabilities of these reference models in capturing the training distribution, the target distributions of downstream generative tasks rarely align perfectly with this reference distribution. For example, in conditional generative learning (Dhariwal and Nichol, 2021; Ho and Salimans, 2021), reference diffusion models generate samples from a mixture of distributions, whereas downstream tasks require sampling from specific constituent distributions. Similarly, in posterior sampling contexts (Chung et al., 2023; Song et al., 2023; Chen et al., 2025; Chang et al., 2025b; Purohit et al., 2025; Martin et al., 2025), the reference distribution serves as the prior, while the target is the posterior defined by tilting the prior with a measurement likelihood. Furthermore, downstream tasks frequently impose additional constraints, such as human preferences or safety considerations (Domingo-Enrich et al., 2024, 2025; Uehara et al., 2025b; Kim et al., 2025; Sabour et al., 2025; Denker et al., 2025; Ren et al., 2025), which must be satisfied without compromising the model’s generative quality.
To bridge the gap between reference and target distributions, researchers have proposed numerous alignment methods (Xu et al., 2023; Lee et al., 2023b; Fan et al., 2023; Clark et al., 2024; Domingo-Enrich et al., 2024, 2025; Uehara et al., 2024b, 2025b). These strategies generally fall into two categories: fine-tuning and inference-time alignment. Fine-tuning approaches involve retraining the reference score network via supervised learning (Lee et al., 2023b), reinforcement learning (Fan et al., 2023; Black et al., 2024; Clark et al., 2024; Uehara et al., 2024c), or classifier-free fine-tuning (Ho and Salimans, 2021; Zhang et al., 2023; Yuan et al., 2023). Despite its conceptual simplicity, fine-tuning presents significant limitations. First, it often requires a substantial collection of high-quality samples from the target distribution, which may be unavailable in practical scenarios like posterior sampling. Second, the computational cost of retraining score networks can be prohibitive, particularly for large-scale models with billions of parameters (Uehara et al., 2025b). Third, fine-tuning is vulnerable to over-optimization (Gao et al., 2023; Rafailov et al., 2024; Kim et al., 2025), where the network overfits to limited target samples or preferences, causing it to “forget” the valuable prior information encoded in the reference model. This degradation undermines the fundamental advantage of leveraging pre-trained models.
In contrast, inference-time alignment (Uehara et al., 2025b; Kim et al., 2025; Sabour et al., 2025; Denker et al., 2025; Ren et al., 2025; Pachebat et al., 2025) eliminates the need to retrain the reference diffusion model. These methods offer substantial computational advantages and preserve the generative capacity of the underlying model. The core technique is guidance (Jiao et al., 2025), which incorporates target information by introducing an additional drift term to the reference diffusion model. Within the framework of Doob’s $h$-transform (Rogers and Williams, 2000; Särkkä and Solin, 2019; Chewi, 2025), the score function for the target tilted distribution decomposes into the sum of the reference score and a guidance term, where the guidance is defined as the gradient of the log-Doob’s $h$-function (Heng et al., 2024; Tang and Xu, 2024; Denker et al., 2024, 2025). This relationship has also been investigated through the lens of classifier guidance (Dhariwal and Nichol, 2021), stochastic optimal control (Han et al., 2024; Tang and Zhou, 2025), and Bayes’ rule (Chung et al., 2023; Song et al., 2023). The primary challenge lies in accurately estimating this guidance.
Guidance estimation methods can be categorized into two main approaches: approximation and learning. The approximation approach, exemplified by diffusion posterior sampling (Chung et al., 2023) and loss-guided diffusion (Song et al., 2023), relies on heuristic approximations that often lead to inconsistencies with the underlying mathematical formulation. The learning approach (Dhariwal and Nichol, 2021; Tang and Xu, 2024; Denker et al., 2024, 2025) attempts to learn the necessary components by deep neural networks. Classifier guidance (Dhariwal and Nichol, 2021) learns Doob’s $h$-function via a classifier, but this is effective primarily for discrete labels. Furthermore, the convergence of the plug-in gradient estimator of the classifier is not guaranteed, potentially undermining guidance reliability (Mou, 2025, Section 3.2.2). To mitigate this, Tang and Xu (2024) estimate Doob’s $h$-function and its gradient using separate neural networks, increasing training complexity. While Denker et al. (2024) directly learn the guidance, their method requires samples from the target tilted distribution. Denker et al. (2025) attempt to address this data requirement via iterative retraining.
To address these limitations, we introduce variationally stable Doob’s $h$-matching, a novel framework for provable guidance estimation in inference-time alignment. We propose a gradient-regularized regression method that simultaneously estimates Doob’s $h$-function and its gradient, yielding a consistent estimator of the guidance. When combined with the pre-trained reference score, our method enables efficient sampling from the target distribution without the need for computationally expensive fine-tuning or access to target distribution samples.
1.1  Contributions
Our main contributions are summarized as follows:
  1. (i)
    We introduce variationally stable Doob’s matching, a novel guidance estimation framework for controllable diffusion models grounded in Doob’s $h$-transform. The Doob $h$-function is estimated via a least-squares regression approach augmented with a gradient regularization, and the plug-in gradient of the logarithm of the resulting $h$-function estimator yields an estimator for the Doob’s guidance. Additionally, this method is derivative-free, meaning it does not require access to the gradient of the weight function between the target tilted distribution and the reference distribution.
  2. (ii)
    We establish non-asymptotic convergence rates for variationally stable Doob’s matching, showing that the proposed method guarantees convergence of both the $h$-function estimator and its gradient. These results provide rigorous theoretical guarantees for Doob’s guidance estimation (Theorem 5.3). Moreover, we derive non-asymptotic convergence rates for the induced controllable diffusion models, thereby establishing rigorous guarantees for the generated distributions in the $2$-Wasserstein distance (Theorem 5.6). Additionally, we obtain convergence rates that depend only on the intrinsic dimension, thereby mitigating the curse of dimensionality under low-dimensional subspace assumptions (Theorem 5.9 and Corollary 5.10).
1.2  Organization
The remainder of this paper is organized as follows. In Section 2, we provide a brief introduction to diffusion models. In Section 3, we propose the stochastic dynamics of controllable diffusion models within the framework of Doob’s $h$-transform, and we present a practical algorithm to simultaneously estimate Doob’s $h$-function and its gradient. Section 5 establishes non-asymptotic error bounds for both the estimation of the $h$-function and the induced controllable diffusion models. Finally, concluding remarks are provided in Section 6. Detailed proofs of theoretical results are deferred to the appendix.
2  Preliminaries on Diffusion Models
2.1  Forward and time-reversal process
We consider the diffusion model for a reference distribution $p_{0}$. The forward process of the reference diffusion model is defined by the Ornstein–Uhlenbeck process:
(2.1) $\mathrm{d}\mathbf{X}_{t}=-\mathbf{X}_{t}\,\mathrm{d}t+\sqrt{2}\,\mathrm{d}\mathbf{B}_{t},\quad t\in(0,T),\penalty 10000\ \mathbf{X}_{0}\sim p_{0},$
where $\mathbf{B}_{t}$ is a $d$-dimensional standard Brownian motion, and $T>0$ is the terminal time. The transition distribution of the forward process can be expressed as:
(2.2) $(\mathbf{X}_{t}|\mathbf{X}_{0}=\mathbf{x}_{0})\sim\mathcal{N}(\mu_{t}\mathbf{x}_{0},\sigma_{t}^{2}\mathbf{I}_{d}),$
where the mean and variance coefficients are given, respectively, as $\mu_{t}=\exp(-t)$ and $\sigma_{t}^{2}=1-\exp(-2t)$. The forward process (2.1) is commonly referred to as the variance-preserving (VP) SDE (Song et al., 2021) as $\mu_{t}^{2}+\sigma_{t}^{2}=1$ for each $t\in(0,T)$. Denote by $p_{t}$ the marginal density of $\mathbf{X}_{t}$ for each $t\in(0,T)$, which satisfies
(2.3) $p_{t}(\mathbf{x}_{t})=\int\varphi_{d}(\mathbf{x}_{t};\mu_{t}\mathbf{x}_{0},\sigma_{t}^{2}\mathbf{I}_{d})p_{0}(\mathbf{x}_{0})\,\mathrm{d}\mathbf{x}_{0},$
where $\varphi_{d}(\cdot;\mu_{t}\mathbf{x}_{0},\sigma_{t}^{2}\mathbf{I}_{d})$ denotes the density function of the Gaussian distribution ${\mathcal{N}}(\mu_{t}\mathbf{x}_{0},\sigma_{t}^{2}\mathbf{I}_{d})$. The corresponding time-reversal process (Anderson, 1982) of (2.1) is characterized by:
(2.4) $\,\mathrm{d}\mathbf{X}_{t}^{\leftarrow}$ $=\Big(\mathbf{X}_{t}^{\leftarrow}+2\overbrace{\nabla\log p_{T-t}(\mathbf{X}_{t}^{\leftarrow})}^{\text{base score}}\Big)\,\mathrm{d}t+\sqrt{2}\,\mathrm{d}\mathbf{B}_{t},\quad t\in(0,T),$
$\mathbf{X}_{0}^{\leftarrow}$ $\sim p_{T}.$
It has been established that the path measure of the time-reversal process $(\mathbf{X}_{t}^{\leftarrow})_{0\leq t\leq T}$ corresponds exactly to the reverse of the forward process $(\mathbf{X}_{t})_{0\leq t\leq T}$ (Anderson, 1982).
2.2  Path measure and filtration
We formally define the probability space for the time-reversal process (2.4). Let $\Omega\coloneq C([0,T],{\mathbb{R}}^{d})$ be the space of continuous functions mapping $[0,T]$ to ${\mathbb{R}}^{d}$, equipped with the topology of uniform convergence. Let $\mathcal{F}$} be the Borel $\sigma$-algebra on $\Omega$. We define the canonical process $\mathbf{X}^{\leftarrow}$ on $\Omega$ via the coordinate mapping $\mathbf{X}_{t}^{\leftarrow}(\omega)=\omega(t)$ for all $\omega\in\Omega$. The natural filtration is given by ${\mathbb{F}}\coloneq({\mathcal{F}}_{t})_{0\leq t\leq T}$, where ${\mathcal{F}}_{t}\coloneq\sigma(\mathbf{X}_{s}^{\leftarrow}\mid 0\leq s\leq t)$ is the $\sigma$-algebra generated by the path up to time $t$. We denote by $\mathbb{P}$} the probability measure on $(\Omega,{\mathcal{F}})$ induced by the law of the solution to the SDE (2.4) with initial distribution $\mathbf{X}_{0}^{\leftarrow}\sim p_{T}$. Consequently, the filtered probability space is denoted as $(\Omega,{\mathcal{F}},{\mathbb{F}},{\mathbb{P}})$, and $\mathbf{B}_{t}$ is an $\mathbb{F}$}-Brownian motion under $\mathbb{P}$}.
2.3  Training phase: score matching
In generative learning, the exact reference score $\nabla\log p_{t}$ in the time-reversal process (2.4) is intractable. One can estimate the reference score using samples from the reference density $p_{0}$ via standard techniques such as implicit score matching (Hyvärinen, 2005), sliced score matching (Song et al., 2020), and denoising score matching (Vincent, 2011). Let $\widehat{\mathbf{s}}:(0,T)\times{\mathbb{R}}^{d}\to{\mathbb{R}}^{d}$ denote an estimator for the prior score, that is,
(2.5) $\|\widehat{\mathbf{s}}(t,\cdot)-\nabla\log p_{T-t}\|_{L^{2}(p_{T-t})}\leq\varepsilon,$
for a small tolerence $\varepsilon\ll 1$. Considerable research has established theoretical guarantees for this score estimation (Tang and Yang, 2024; Oko et al., 2023; Fu et al., 2024; Ding et al., 2025a), leveraging standard techniques from non-parametric regression with deep neural networks (Bauer and Kohler, 2019; Schmidt-Hieber, 2020; Kohler and Langer, 2021; Jiao et al., 2023).
2.4  Inference phase: sampling
Given a reference score estimator $\widehat{\mathbf{s}}$ in (2.5), the inference phase of diffusion models aims to generate samples by simulating the time-reversal process with estimated score. Since the explicit solution of the time-reversal process is intractable, we employ an exponential integrator (Hochbruck and Ostermann, 2005, 2010; Lu et al., 2022a; Zhang and Chen, 2023). This approach is well-suited for solving the time-reversal process due to the semi-linearity of the drift term of the SDE in (2.4).
Let $K\in{\mathbb{N}}$ denote the number of discretization steps, and let $T_{0}>0$ be an early-stopping time. We define a sequence of uniform time points $t_{k}\coloneq kh$ for $k=0,\dots,K$, where the step size is $h\coloneq(T-T_{0})/K$. In each time sub-interval, the exponential integrator approximates the score function by its value at the left endpoint:
(2.6) $\,\mathrm{d}\widehat{\mathbf{X}}_{t}^{\leftarrow}$ $=(\widehat{\mathbf{X}}_{t}^{\leftarrow}+2\widehat{\mathbf{s}}(kh,\widehat{\mathbf{X}}_{kh}^{\leftarrow}))\,\mathrm{d}t+\sqrt{2}\,\mathrm{d}\mathbf{B}_{t},\quad t\in[kh,(k+1)h),$
$\widehat{\mathbf{X}}_{0}^{\leftarrow}$ $\sim{\mathcal{N}}({\mathbf{0}},\mathbf{I}_{d}),$
where $0\leq k\leq K-1$. The resulting linear approximation to the original semi-linear SDE has the following explicit solution:
$\widehat{\mathbf{X}}_{(k+1)h}^{\leftarrow}=\exp(h)\widehat{\mathbf{X}}_{kh}^{\leftarrow}+2\phi^{2}(h)\widehat{\mathbf{s}}(kh,\widehat{\mathbf{X}}_{kh}^{\leftarrow})+\phi(2h)\bm{\xi}_{k},\quad 0\leq k\leq K-1,$
where $\phi(z)\coloneq\sqrt{\exp(z)-1}$, and $\bm{\xi}_{0},\ldots,\bm{\xi}_{K-1}$ are i.i.d. standard Gaussian random variables.
Remark 2.1 (Initialization).
Note that the true initial distribution of the time-reversal process (2.4) is $p_{T}$, rather than the ${\mathcal{N}}({\mathbf{0}},\mathbf{I}_{d})$ used in (2.6). We adopt the standard normal distribution because sampling from ${\mathcal{N}}({\mathbf{0}},\mathbf{I}_{d})$ is significantly more computationally tractable. This approximation is justified by the fact that $p_{T}$ converges to ${\mathcal{N}}({\mathbf{0}},\mathbf{I}_{d})$ exponentially in KL-divergence as $T\to\infty$ (Bakry et al., 2014; Vempala and Wibisono, 2019); thus, the Gaussian initialization is valid for a sufficiently large terminal time $T$.
3  Controllable Diffusion Models and Doob’s Transform
In this section, we propose controllable diffusion models for sampling from a target distribution, defined as the reference distribution tilted by a weight function. We utilize the theory of measure change for diffusion processes on the filtered space $(\Omega,{\mathcal{F}},{\mathbb{F}},{\mathbb{P}})$.
3.1  Problem setup
We assume access to a pre-trained reference diffusion model (2.6) that generates samples approximately from the reference distribution $p_{0}$. We aim to sample from a tilted distribution $q_{0}$, defined by reweighting the reference distribution with a known weight function $w:{\mathbb{R}}^{d}\to{\mathbb{R}}_{\geq 0}$:
(3.1) $q_{0}(\mathbf{x})\coloneq\frac{w(\mathbf{x})p_{0}(\mathbf{x})}{Z},\quad\text{where}\penalty 10000\ Z\coloneq\int w(\mathbf{x})p_{0}(\mathbf{x})\,\mathrm{d}\mathbf{x}<\infty.$
Our goal is to derive a new diffusion process that generates samples from the tilted distribution $q_{0}$ directly by introducing a drift correction to the pre-trained reference diffusion model (2.6).
This problem encompasses a wide range of application scenarios.
Example 1 (Bayesian inverse problems).
Bayesian inverse problems play a critical role in scientific computing (Stuart, 2010; Kantas et al., 2014; Ding et al., 2024), image science (Chung et al., 2023; Mardani et al., 2024; Purohit et al., 2025; Chang et al., 2025b), and medical imaging (Song et al., 2022). In Bayesian inverse problems, we aim to recover an unknown signal $\mathbf{X}_{0}\in{\mathbb{R}}^{d}$ from noisy measurements $\mathbf{Y}\in{\mathbb{R}}^{m}$, which are linked by the following measurement model:
(3.2) $\mathbf{Y}={\mathcal{A}}(\mathbf{X}_{0})+\mathbf{n},$
where $\mathcal{A}:{\mathbb{R}}^{d}\to{\mathbb{R}}^{m}$ is a known measurement operator, and $\mathbf{n}\in{\mathbb{R}}^{m}$ represents a measurement noise with a known distribution. The Bayesian approach incorporates prior knowledge about $\mathbf{X}_{0}$ in the form of a prior distribution $p_{0}$. Given observed measurements $\mathbf{Y}=\mathbf{y}$, the goal of Bayesian inverse problems is to sample from the posterior distribution:
(3.3) $q_{0}(\mathbf{x}):=p_{\mathbf{X}_{0}|\mathbf{Y}}(\mathbf{x}|\mathbf{y})=\frac{w(\mathbf{x})p_{0}(\mathbf{x})}{Z},$
where $w(\mathbf{x})\coloneq p_{\mathbf{Y}|\mathbf{X}_{0}}(\mathbf{y}|\mathbf{x})$ is a likelihood determined by the measurement model (3.2), and $Z$ is a partition function to ensure $q_{0}$ is a valid probability density. For example, for a Gaussian noise $\mathbf{n}\sim N({\mathbf{0}},\sigma^{2}\mathbf{I}_{d})$, it holds that
(3.4) $w(\mathbf{x})=(2\pi\sigma^{2})^{-\frac{d}{2}}\exp\Big(-\frac{1}{2\sigma^{2}}\|\mathbf{y}-\mathcal{A}(\mathbf{x})\|_{2}^{2}\Big).$
In Bayesian inverse problems, one typically has a reference diffusion model pre-trained on the prior distribution, and aims to sample from the posterior distribution (3.3) without retraining the reference model.
Example 2 (Reward-guided generation).
In the reward-guided generation (Domingo-Enrich et al., 2025; Uehara et al., 2025b; Kim et al., 2025; Sabour et al., 2025; Denker et al., 2025; Ren et al., 2025), human preferences and constraints can be encoded into a reward function $r:{\mathbb{R}}^{d}\to{\mathbb{R}}$. For instance, in text-to-image generation, the reward function $r$ quantifies how well the generated data aligns with the input prompt. In practice, such reward function can be learned from the human feedback or preference data (Stiennon et al., 2020; Ouyang et al., 2022; Lee et al., 2023b). For the sake of simplicity, we assume throughout this work that the reward function has already been given. A naive approach to reward-guided generation is to maximize the expected reward $\max_{\pi\in{\mathcal{P}}}{\mathbb{E}}_{\pi}[r]$, where $\mathcal{P}$} is the set of probability measures on ${\mathbb{R}}^{d}$. However, solely maximizing the expected reward may lead to over-optimization and degenerate solutions (Kim et al., 2025). To mitigate this, entropy regularization (Uehara et al., 2024a; Tang and Zhou, 2025) is incorporated into the objective functional, yielding the following optimization problem:
(3.5) $q_{0}\coloneq\operatorname*{arg\,max}_{\pi\in{\mathcal{P}}}{\mathbb{E}}_{\pi}[r]-\alpha\mathrm{KL}(\pi\|p_{0}),$
where $\alpha>0$ is a regularization parameter, and $p_{0}$ is the density of a reference distribution, i.e., the distribution corresponding to the pre-trained reference diffusion model. This objective comprises two components: the expected reward, which captures human preferences, and the KL-divergence term, which prevents the distribution from deviating excessively from the reference diffusion model. The closed-form solution to this optimization problem (3.5) is given by (Rafailov et al., 2023):
(3.6) $q_{0}(\mathbf{x})=\frac{w(\mathbf{x})p_{0}(\mathbf{x})}{Z},\quad w(\mathbf{x}):=\exp\Big(\frac{r(\mathbf{x})}{\alpha}\Big),$
where $Z$ is the partition function to ensure $q_{0}$ is a valid probability density. In reward-guided generation, the central objective is to incorporate preferences exclusively during the inference phase, thereby avoiding the substantial computational cost of retraining the large-scale reference diffusion model.
Example 3 (Transfer learning for diffusion models).
Diffusion models have achieved remarkable success in image generation. However, their performance critically depends on the availability of large-scale training data. In the scenarios of few-shot generation, training diffusion models solely on limited samples typically results in poor generative performance. To retain strong generative capability under data scarcity, a common strategy is to transfer expressive diffusion models pre-trained on large datasets to the target domain (Ouyang et al., 2024; Wang et al., 2024; Zhong et al., 2025; Bahram et al., 2026). Formally, transfer learning aims to adapt a model pre-trained on a large-scale source distribution to a target distribution of much smaller size and diversity of samples. However, directly fine-tuning a large pre-trained diffusion model using only limited target samples often leads to severe overfitting (Wang et al., 2024). To address this issue, transfer learning approaches for diffusion models typically train a lightweight guidance network on the limited target data and combine it with the pre-trained reference score network, yielding a modified diffusion model capable of sampling from the target distribution (Ouyang et al., 2024; Zhong et al., 2025; Bahram et al., 2026). Concretely, Ouyang et al. (2024) estimates the density ratio between the target and source distributions,
(3.7) $w(\mathbf{x}):=\frac{q_{0}(\mathbf{x})}{p_{0}(\mathbf{x})},$
using limited samples from the target dsitribution. For simplicity, we assume throughout this work that the density ratio is known. Under this assumption, the objective of transfer learning for diffusion models is to sample from the target distribution $q_{0}$ by leveraging the pre-trained reference diffusion model together with the estimated density ratio.
3.2  Controllable diffusion models with Doob’s $h$-transform
In this subsection, we achieve inference-time alignment by incorporating guidance into the reference diffusion models via Doob’s $h$-transform (Rogers and Williams, 2000; Särkkä and Solin, 2019; Heng et al., 2024; Chewi, 2025). We begin by constructing a target path measure $\mathbb{Q}$} on the filtered probability space $(\Omega,{\mathcal{F}},{\mathbb{F}})$. We assume that $\mathbb{Q}$} is absolutely continuous with respect to the reference path measure $\mathbb{P}$} defined in Section (2.2). By the Radon-Nikodym theorem, this relationship is characterized by the existence of an $\mathbb{F}$}-adapted process $(L_{t})_{0\leq t\leq T}$ with $L_{t}\geq 0$, such that for every $t\in[0,T]$:
(3.8) $L_{t}\coloneq\frac{\,\mathrm{d}{\mathbb{Q}}}{\,\mathrm{d}{\mathbb{P}}}\Big|_{{\mathcal{F}}_{t}}.$
The process $L_{t}$ is known as the Radon-Nikodym derivative process, representing the likelihood ratio between the target and reference measures conditioned on the filtration ${\mathcal{F}}_{t}$.
We impose two boundary conditions on the target path measure $\mathbb{Q}$}: (i) the marginal distribution of the initial state $\mathbf{X}_{0}^{\leftarrow}$ under $\mathbb{Q}$} must match that of $\mathbb{P}$}; and (ii) the marginal distribution of the terminal state $\mathbf{X}_{T}^{\leftarrow}$ under $\mathbb{Q}$} must coincide with the tilted distribution $q_{0}$ defined in (3.1). To enforce these constraints, we specify the Radon-Nikodym derivatives of $\mathbb{Q}$} with respect to $\mathbb{P}$} at the initial and terminal times as follows:
(3.9) $L_{0}=\frac{\,\mathrm{d}{\mathbb{Q}}}{\,\mathrm{d}{\mathbb{P}}}\Big|_{{\mathcal{F}}_{0}}\equiv 1,\quad\text{and}\quad L_{T}=\frac{\,\mathrm{d}{\mathbb{Q}}}{\,\mathrm{d}{\mathbb{P}}}\Big|_{{\mathcal{F}}_{T}}=\frac{\,\mathrm{d}{\mathbb{Q}}}{\,\mathrm{d}{\mathbb{P}}}=\frac{w(\mathbf{X}_{T}^{\leftarrow})}{{\mathbb{E}}^{{\mathbb{P}}}[w(\mathbf{X}_{T}^{\leftarrow})]},$
where the denominator serves as the normalizing constant ensuring that $\mathbb{Q}$} is a valid probability measure, i.e., ${\mathbb{E}}^{{\mathbb{P}}}[L_{T}]=1$.
The rest of our derivation proceeds in two steps. First, we characterize the stochastic dynamics of $\mathbf{X}_{t}^{\leftarrow}$ under the target measure $\mathbb{Q}$}. Second, by invoking the weak uniqueness of solutions to stochastic differential equations (Øksendal, 2003, Lemma 5.3.1), we construct a controllable diffusion process under the reference path measure $\mathbb{P}$} whose terminal distribution coincides with the target tilted distribution $q_{0}$.
Dynamics under the target path measure
Girsanov’s theorem (Øksendal, 2003, Theorem 8.6.8) establishes a fundamental correspondence between a drift shift in the driving Brownian motion and the dynamics of the associated Radon-Nikodym derivative process. Accordingly, we first characterize the evolution of the time-dependent likelihood ratio $L_{t}$.
Proposition 3.1 (Doob’s $h$-function).
The Radon-Nikodym derivative process $L_{t}$, as defined in (3.8), admits the following representation:
$L_{t}={\mathbb{E}}^{{\mathbb{P}}}\Big[\frac{\,\mathrm{d}{\mathbb{Q}}}{\,\mathrm{d}{\mathbb{P}}}\Big|{\mathcal{F}}_{t}\Big]=\frac{h^{*}(t,\mathbf{X}_{t}^{\leftarrow})}{{\mathbb{E}}^{{\mathbb{P}}}[w(\mathbf{X}_{T}^{\leftarrow})]},$
where $h^{*}:[0,T]\times{\mathbb{R}}^{d}\to{\mathbb{R}}$, referred to as Doob’s $h$-function, is defined as the conditional expectation of the terminal weight:
(3.10) $h^{*}(t,\mathbf{x})\coloneq{\mathbb{E}}^{{\mathbb{P}}}[w(\mathbf{X}_{T}^{\leftarrow})\mid\mathbf{X}_{t}^{\leftarrow}=\mathbf{x}].$
Furthermore, the log-likelihood ratio satisfies the following SDE:
$\mathrm{d}(\log L_{t})=\nabla\log h^{*}(t,\mathbf{X}_{t}^{\leftarrow})^{\top}\sqrt{2}\,\mathrm{d}\mathbf{B}_{t}-\|\nabla\log h^{*}(t,\mathbf{X}_{t}^{\leftarrow})\|_{2}^{2}\,\mathrm{d}t.$
The proof of Proposition 3.1 is provided in Appendix A. With the dynamics of $\log L_{t}$ established, we derive the stochastic dynamics of $\mathbf{X}_{t}^{\leftarrow}$ under the target measure $\mathbb{Q}$} via Girsanov’s theorem (Øksendal, 2003, Theorem 8.6.6).
Proposition 3.2.
Let the reference process $\mathbf{X}_{t}^{\leftarrow}$ satisfy the SDE (2.4) under the path measure $\mathbb{P}$}, and let $\mathbb{Q}$} be the target path measure defined by (3.9). Assume that the Novikov condition holds:
(3.11) ${\mathbb{E}}^{{\mathbb{P}}}\Big[\exp\Big(\int_{0}^{T}\|\nabla\log h^{*}(s,\mathbf{X}_{s}^{\leftarrow})\|_{2}^{2}\,\mathrm{d}s\Big)\Big]<\infty.$
Define a process $(\widetilde{\mathbf{B}}_{t})_{1\leq t\leq T}$ by
(3.12) $\mathrm{d}\widetilde{\mathbf{B}}_{t}=\,\mathrm{d}\mathbf{B}_{t}-\sqrt{2}\nabla\log h^{*}(t,\mathbf{X}_{t}^{\leftarrow})\,\mathrm{d}t,$
where $\mathbf{B}_{t}$ is a Brownian motion under $\mathbb{P}$}. Then $\widetilde{\mathbf{B}}_{t}$ is a standard Brownian motion under $\mathbb{Q}$}. Further, under the path measure $\mathbb{Q}$}, the reference process $\mathbf{X}_{t}^{\leftarrow}$ in (2.4) evolves according to:
(3.13) $\mathrm{d}\mathbf{X}_{t}^{\leftarrow}=(\mathbf{X}_{t}^{\leftarrow}+2\nabla\log p_{T-t}(\mathbf{X}_{t}^{\leftarrow})+2\nabla\log h^{*}(t,\mathbf{X}_{t}^{\leftarrow}))\,\mathrm{d}t+\sqrt{2}\,\mathrm{d}\widetilde{\mathbf{B}}_{t}.$
The proof is deferred to Appendix A. By the construction in (3.9), the law of $\mathbf{X}_{T}^{\leftarrow}$ under $\mathbb{Q}$} coincides with the target tilted distribution $q_{0}$.
Remark 3.3 (Novikov condition).
The condition in (3.11) ensures that the exponential local martingale defined by the drift shift is a true martingale (Karatzas and Shreve, 1998, Corollary 5.13), which is sufficient for the Radon-Nikodym derivative to be well-defined and for Girsanov’s theorem to apply.
Controllable diffusion process under the reference measure
Although Proposition 3.2 constructs stochastic dynamics driven by Brownian motion under $\mathbb{Q}$} that achieve the desired terminal distribution $q_{0}$, practical implementation necessitates an SDE driven by Brownian motion under the reference measure $\mathbb{P}$}.
To address this, we construct a surrogate process $\mathbf{Z}_{t}^{\leftarrow}$ on the reference probability space $(\Omega,{\mathcal{F}},{\mathbb{F}},{\mathbb{P}})$ that adopts the drift derived for $\mathbb{Q}$}:
(3.14) $\mathrm{d}\mathbf{Z}_{t}^{\leftarrow}=\Big(\mathbf{Z}_{t}^{\leftarrow}+2\underbrace{\nabla\log p_{T-t}(\mathbf{Z}_{t}^{\leftarrow})}_{\text{base score}}+2\underbrace{\nabla\log h^{*}(t,\mathbf{Z}_{t}^{\leftarrow})}_{\text{Doob's guidance}}\Big)\,\mathrm{d}t+\sqrt{2}\,\mathrm{d}\mathbf{B}_{t}.$
Since the process $\mathbf{Z}_{t}^{\leftarrow}$ driven by the $\mathbb{P}$}-Brownian motion (3.14) satisfies the same SDE as $\mathbf{X}_{t}^{\leftarrow}$ driven by the $\mathbb{Q}$}-Brownian motion (3.13), the weak uniqueness property of SDE solutions guarantees (Øksendal, 2003, Lemma 5.3.1) that the law of $\mathbf{Z}_{t}^{\leftarrow}$ under $\mathbb{P}$} is identical to the law of $\mathbf{X}_{t}^{\leftarrow}$ under $\mathbb{Q}$} for all $t\in[0,T]$. Thus, the law of $\mathbf{Z}_{T}^{\leftarrow}$ under $\mathbb{P}$} coincides the target tilted distribution $q_{0}$. For a detailed formal statement, see Øksendal (2003, Theorem 8.6.8).
4  Variationally Stable Doob’s Matching
We have thus far established a controllable diffusion process (3.14) capable of generating samples from the target tilted distribution $q_{0}$. However, the Doob’s $h$-guidance, required by (3.14), remains intractable. This subsection proposes a variationally stable Doob’s matching method to address the estimation of the Doob’s guidance.
4.1  Vanilla least-squares regression for Doob’s matching
For any $t\in(0,T)$, the Doob’s $h$-function $h_{t}^{*}\coloneq h^{*}(t,\cdot)$ defined as (3.10) is the unique minimizer of the following implicit Doob’s matching objective:
(4.1) $\mathcal{J}_{t}(h_{t})$ $=\mathbb{E}^{\mathbb{P}}\big[\|h_{t}(\mathbf{X}_{t}^{\leftarrow})-w(\mathbf{X}_{T}^{\leftarrow})\|_{2}^{2}\big]$
$=\mathbb{E}_{\mathbf{X}_{0}\sim p_{0}}\mathbb{E}_{\bm{\epsilon}\sim\mathcal{N}(\mathbf{0},\mathbf{I}_{d})}\big[\|h_{t}(\mu_{T-t}\mathbf{X}_{0}+\sigma_{T-t}\bm{\epsilon})-w(\mathbf{X}_{0})\|_{2}^{2}\big],$
where $w:\mathbb{R}^{d}\to\mathbb{R}_{\geq 0}$ is the known weight function in (3.1). The following proposition justifies the use of $\mathcal{J}_{t}$ as a surrogate for the explicit $L^{2}$-distance.
Proposition 4.1.
For every $t\in(0,T)$, the Doob’s $h$-function $h_{t}^{*}$ in (3.10) minimizes the implicit Doob’s matching objective (4.1). Further,
${\mathcal{J}}_{t}(h_{t})={\mathbb{E}}^{{\mathbb{P}}}\big[\|h_{t}(\mathbf{X}_{t}^{\leftarrow})-h_{t}^{*}(\mathbf{X}_{t}^{\leftarrow})\|_{2}^{2}\big]+V_{t}^{2},$
where $V_{t}^{2}\coloneq{\mathbb{E}}^{{\mathbb{P}}}[\mathrm{Var}(w(\mathbf{X}_{T}^{\leftarrow})|\mathbf{X}_{t}^{\leftarrow})]$ is a constant independent of $h_{t}$.
The proof of Proposition 4.1 is provided in Appendix B.
4.2  Limitations of vanilla regression
Crucially, computing the Doob’s guidance $\nabla\log h_{t}^{*}$ in (3.14) requires estimating not only the function $h_{t}^{*}$ itself but also its gradient $\nabla h_{t}^{*}$, as the guidance is given by:
$\nabla\log h_{t}^{*}(\mathbf{x})=\frac{\nabla h_{t}^{*}(\mathbf{x})}{h_{t}^{*}(\mathbf{x})},\quad\mathbf{x}\in{\mathbb{R}}^{d}.$
According to Proposition 4.1, the objective $\mathcal{J}_{t}$ is only coercive with respect to the $L^{2}(p_{T-t})$-norm. Specifically, for any $h_{t}\in L^{2}(p_{T-t})$,
(4.2) ${\mathcal{J}}_{t}(h_{t})-{\mathcal{J}}_{t}(h_{t}^{*})=\|h_{t}-h_{t}^{*}\|_{L^{2}(p_{T-t})}^{2}.$
However, even if $h_{t}$ is close to $h_{t}^{*}$ in the $L^{2}$ sense, the gradient $\nabla h_{t}$ may remain highly oscillatory. This leads to an unstable plug-in estimator for the Doob’s guidance, a difficulty noted in prior works such as Tang and Xu (2024, Section 3.2.1) and Mou (2025, Section 3.2.2). Similar issues arise in related contexts, including classifier guidance (Dhariwal and Nichol, 2021), Monte Carlo regression (Uehara et al., 2025b, Section 2.2), and Ouyang et al. (2024).
To illustrate this fundamental limitation, consider the sequence of functions $f_{n}:{\mathcal{I}}\to{\mathbb{R}}$ defined by $x\mapsto n^{-1}\sin(nx)$, alongside the zero function $f_{0}\equiv 0$, where ${\mathcal{I}}:=[0,2\pi]$. While $\|f_{n}-f_{0}\|_{L^{2}({\mathcal{I}})}\to 0$ as $n\to\infty$, the distance between their derivatives does not vanish, i.e., $\lim_{n\to\infty}\|f_{n}^{\prime}-f_{0}^{\prime}\|_{L^{2}({\mathcal{I}})}\neq 0$. As a result, the convergence of the plug-in Doob’s guidance estimator derived from vanilla regression (4.1), as utilized in Uehara et al. (2025b); Ouyang et al. (2024), cannot be guaranteed in general.
To mitigate this, Tang and Xu (2024) estimate Doob’s $h$-function and its gradient separately via a martingale approach. In contrast, in the remainder of this work, we propose an approach to simultaneously estimate both the function and its gradient.
4.3  Variationally stable Doob’s matching
To simultaneously estimate the Doob’s $h$-function and its gradient, we adopt a gradient-regularized regression (Drucker and Le Cun, 1991, 1992; Ding et al., 2025b). The population risk is defined by incorporating an additional Sobolev regularization to (4.1):
(4.3) $h_{t}^{\lambda}=\operatorname*{arg\,min}_{h_{t}:{\mathbb{R}}^{d}\to{\mathbb{R}}^{d}}{\mathcal{J}}_{t}^{\lambda}(h_{t})$ $\coloneq{\mathcal{J}}_{t}(h_{t})+\overbrace{\lambda{\mathbb{E}}^{{\mathbb{P}}}\big[\|\nabla h_{t}(\mathbf{X}_{t}^{\leftarrow})\|_{2}^{2}\big]}^{\text{gradient regularization}}$
$={\mathcal{J}}_{t}(h_{t})+\lambda{\mathbb{E}}_{\mathbf{X}_{0}\sim p_{0}}{\mathbb{E}}_{{\bm{\varepsilon}}\sim{\mathcal{N}}({\mathbf{0}},\mathbf{I}_{d})}\big[\|\nabla h_{t}(\mu_{T-t}\mathbf{X}_{0}+\sigma_{T-t}{\bm{\varepsilon}})\|_{2}^{2}\big],$
where $\lambda>0$ is a regularization parameter. The following results characterize the regularization gap and the variational stability of this formulation.
Proposition 4.2 (Regularization gap).
Let $\lambda>0$, $h_{t}^{*}$ be the Doob’s $h$-function defined as (3.10), and $h_{t}^{\lambda}$ be the minimizer of ${\mathcal{J}}_{t}^{\lambda}$ defined as (4.3). Then $h_{t}^{*}\in H^{2}(p_{T-t})$, and
$\|h_{t}^{\lambda}-h_{t}^{*}\|_{L^{2}(p_{T-t})}^{2}$ $\leq\lambda^{2}\|\Delta h_{t}^{*}+\nabla h_{t}^{*}\cdot\nabla\log p_{T-t}\|_{L^{2}(p_{T-t})}^{2},$
$\|\nabla h_{t}^{\lambda}-\nabla h_{t}^{*}\|_{L^{2}(p_{T-t})}^{2}$ $\leq\lambda\|\Delta h_{t}^{*}+\nabla h_{t}^{*}\cdot\nabla\log p_{T-t}\|_{L^{2}(p_{T-t})}^{2}.$
The proof of Proposition 4.2 is provided in Appendix B. This proposition demonstrates that as $\lambda\to 0$, the minimizer $h_{t}^{\lambda}$ of the regularized objective (4.3) converges to $h_{t}^{*}$ in $H^{1}$-norm. More importantly, the objective is variationally stable in the $H^{1}$ sense:
Proposition 4.3 (Variational stability).
Let $\lambda>0$, and $h_{t}^{\lambda}$ be the minimizer of ${\mathcal{J}}_{t}^{\lambda}$ defined as (4.3). Then for any $h_{t}\in H^{1}(p_{T-t})$, we have
$\frac{1}{\max\{\lambda,1\}}\big\{{\mathcal{J}}_{t}^{\lambda}(h_{t})-{\mathcal{J}}_{t}^{\lambda}(h_{t}^{\lambda})\big\}\leq\|h_{t}-h_{t}^{\lambda}\|_{H^{1}(p_{T-t})}^{2}\leq\frac{1}{\min\{\lambda,1\}}\big\{{\mathcal{J}}_{t}^{\lambda}(h_{t})-{\mathcal{J}}_{t}^{\lambda}(h_{t}^{\lambda})\big\}.$
The proof of Proposition 4.3 is provided in Appendix B. We refer to the regularized objective ${\mathcal{J}}_{t}^{\lambda}$ as variationally stable because the convergence in this objective functional necessitates simultaneous convergence in both the function values and their gradients, i.e., stability in the $H^{1}$ sense. Such variational stability ensures that any candidate function $h_{t}$ achieving a low objective value $\mathcal{J}_{t}^{\lambda}(h_{t})$ is guaranteed to be an approximation of the ground-truth Doob’s $h$-function in both value and gradient.
Comparison between vanilla and gradient-regularized Doob’s matching
The distinction between vanilla Doob’s matching (4.1) and the proposed gradient-regularized Doob’s matching (4.3) is fundamental for stable diffusion guidance estimation. While vanilla regression guarantees convergence in the $L^{2}$-norm; see (4.2), it can be unstable in the $H^{1}$ sense. In contrast, the proposed objective ${\mathcal{J}}_{t}^{\lambda}$ is coercive with respect to the $H^{1}$ norm; that is, for a fixed $\lambda>0$,
${\mathcal{J}}_{t}^{\lambda}(h_{t})-{\mathcal{J}}_{t}^{\lambda}(h_{t}^{\lambda})\simeq\|h_{t}-h_{t}^{\lambda}\|_{H^{1}(p_{T-t})}^{2}.$
This property ensures the simultaneous estimation of Doob’s $h$-function and its gradient, thereby yielding a plug-in estimator for Doob’s guidance with mathematical guarantees.
4.4  Doob’s guidance estimation
Since the expectation in the population risk (4.3) is computationally intractable, we approximate it by empirical risk using independent and identically distributed samples:
(4.4) $\widehat{{\mathcal{J}}}_{t}^{\lambda}(h_{t})\coloneq\widehat{{\mathcal{J}}}_{t}(h_{t})+\frac{\lambda}{n}\sum_{i=1}^{n}\|\nabla h_{t}(\mu_{T-t}\mathbf{X}_{0}^{i}+\sigma_{T-t}{\bm{\varepsilon}}^{i})\|_{2}^{2},$
where the empirical least-squares risk is defined as
(4.5) $\widehat{{\mathcal{J}}}_{t}(h_{t})\coloneq\frac{1}{n}\sum_{i=1}^{n}\|h_{t}(\mu_{T-t}\mathbf{X}_{0}^{i}+\sigma_{T-t}{\bm{\varepsilon}}^{i})-w(\mathbf{X}_{0}^{i})\|_{2}^{2}.$
Here $\mathbf{X}_{0}^{1},\ldots,\mathbf{X}_{0}^{n}$ are independent and identically distributed random variables drawn from the reference distribution $p_{0}$, and ${\bm{\varepsilon}}_{1},\ldots,{\bm{\varepsilon}}_{n}$ are independent standard Gaussian random variables. Then one has a gradient-regularized empirical risk minimizer:
(4.6) $\widehat{h}_{t}^{\lambda}\in\operatorname*{arg\,min}_{h_{t}\in{\mathscr{H}}_{t}}\widehat{{\mathcal{J}}}_{t}^{\lambda}(h_{t}),$
where ${\mathscr{H}}_{t}$ is a hypothesis class, which is chosen as a neural network class in this work.
The Doob’s matching with gradient regularization (4.6) yields a valid plug-in estimator of the Doob’s guidance:
(4.7) $\widehat{\mathbf{g}}_{t}^{\lambda}(\mathbf{z})\coloneq\nabla\log\widehat{h}_{t}^{\lambda}(\mathbf{z})=\frac{\nabla\widehat{h}_{t}^{\lambda}(\mathbf{z})}{\widehat{h}_{t}^{\lambda}(\mathbf{z})}\approx\frac{\nabla h_{t}^{*}(\mathbf{z})}{h_{t}^{*}(\mathbf{z})}=\nabla\log h_{t}^{*}(\mathbf{z}),\quad\mathbf{z}\in{\mathbb{R}}^{d}.$
4.5  A summary of computing procedure
By a similar argument as Section 2.4, we have the exponential integrator for the controllable diffusion model:
(4.8) $\,\mathrm{d}\widehat{\mathbf{Z}}_{t}^{\leftarrow}$ $=(\widehat{\mathbf{Z}}_{t}^{\leftarrow}+2\widehat{\mathbf{s}}(kh,\widehat{\mathbf{Z}}_{kh}^{\leftarrow})+2\widehat{\mathbf{g}}^{\lambda}(kh,\widehat{\mathbf{Z}}_{kh}^{\leftarrow}))\,\mathrm{d}t+\sqrt{2}\,\mathrm{d}\mathbf{B}_{t},\quad t\in(kh,(k+1)h),$
$\widehat{\mathbf{Z}}_{0}^{\leftarrow}$ $\sim{\mathcal{N}}({\mathbf{0}},\mathbf{I}_{d}),$
where $0\leq k\leq K-1$, the pre-trained reference score estimator $\widehat{\mathbf{s}}$ is defined as (2.5), and the Doob’s guidance estimator $\widehat{\mathbf{g}}^{\lambda}$ is defined as (4.7).
We apply post-processing to the generated particle $\widehat{\mathbf{Z}}_{T-T_{0}}^{\leftarrow}$ to ensure numerical stability and facilitate the theoretical analysis presented in Theorem 5.6 (Lee et al., 2023a; Chen et al., 2023a). First, we assume the target distribution $q_{0}$ is concentrated on a domain centered at the origin, such as a distribution with compact support (Assumption 1) or with light tails. Consequently, we introduce a truncation operator to the particles obtained from the controllable diffusion model (4.8). Second, because the controllable diffusion process is terminated at an early-stopping time $T_{0}$, there exists a mean shift between the target distribution $q_{0}$ and the early-stopping distribution $q_{T_{0}}\approx\widehat{q}_{T-T_{0}}$, as indicated by (2.2). To mitigate this drift, we employ a scaling operator. Specifically, for $R>0$, we define a truncation operator ${\mathcal{T}}_{R}:\mathbf{z}\mapsto\mathbf{z}{\mathbbm{1}}_{B({\mathbf{0}},R)}(\mathbf{z})$ and a scaling operator ${\mathcal{M}}:\mathbf{z}\mapsto\mu_{T_{0}}^{-1}\mathbf{z}$. The final processed particle is defined as
(4.9) ${\mathcal{M}}\circ{\mathcal{T}}_{R}(\widehat{\mathbf{Z}}_{T-T_{0}}^{\leftarrow})=\mu_{T_{0}}^{-1}\widehat{\mathbf{Z}}_{T-T_{0}}^{\leftarrow}{\mathbbm{1}}_{B({\mathbf{0}},R)}(\widehat{\mathbf{Z}}_{T-T_{0}}^{\leftarrow}),$
and we denote its density by $({\mathcal{M}}\circ{\mathcal{T}}_{R})_{\sharp}\widehat{q}_{T-T_{0}}$.
A complete procedure is summarized in Algorithm 1.
Input: Reference score estimator $\widehat{\mathbf{s}}$, the weight function $w$, the regularization parameter $\lambda$, the step size $h$, and the number of steps $K$.
Output: Particle ${\mathcal{M}}\circ{\mathcal{T}}_{R}(\widehat{\mathbf{Z}}_{T-T_{0}}^{\leftarrow})$ follows the tilted distribution $q_{0}$ approximately.
1 # Doob’s matching
2 Estimate Doob’s $h$-function by $\widehat{h}_{t}^{\lambda}$ via gradient regularized Doob’s matching (4.6).
3 # Controllable generation
4 Generate the initial particle $\widehat{\mathbf{Z}}_{0}^{\leftarrow}\sim{\mathcal{N}}({\mathbf{0}},\mathbf{I}_{d})$.
5 for $k=0,\ldots,K-1$ do
6   Evaluate the reference score: $\widehat{\mathbf{s}}_{k}\leftarrow\widehat{\mathbf{s}}(kh,\widehat{\mathbf{Z}}_{kh}^{\leftarrow})$.
7   Evaluate the Doob’s guidance: $\widehat{\mathbf{g}}_{k}\leftarrow\nabla\log\widehat{h}^{\lambda}(kh,\widehat{\mathbf{Z}}_{kh}^{\leftarrow})$.
8   Exponential integrator: $\widehat{\mathbf{Z}}_{(k+1)h}^{\leftarrow}\sim{\mathcal{N}}(\exp(h)\widehat{\mathbf{Z}}_{kh}^{\leftarrow}+2\phi^{2}(h)(\widehat{\mathbf{s}}_{k}+\widehat{\mathbf{g}}_{k}),\phi^{2}(2h)\mathbf{I}_{d})$.
9 end for
10# Truncation and scaling
11 ${\mathcal{M}}\circ{\mathcal{T}}_{R}(\widehat{\mathbf{Z}}_{T-T_{0}}^{\leftarrow})\leftarrow\mu_{T_{0}}^{-1}\widehat{\mathbf{Z}}_{T-T_{0}}^{\leftarrow}{\mathbbm{1}}_{B({\mathbf{0}},R)}(\widehat{\mathbf{Z}}_{T-T_{0}}^{\leftarrow})$.
return ${\mathcal{M}}\circ{\mathcal{T}}_{R}(\widehat{\mathbf{Z}}_{T-T_{0}}^{\leftarrow})$
Algorithm 1 Inference-time alignment via variationally stable Doob’s matching
5  Convergence Analysis
In this section, we derive a non-asymptotic convergence rate for the variationally stable Doob’s matching (4.6) and the induced controllable diffusion model (4.8). Furthermore, we demonstrate that this convergence rate mitigates the curse of dimensionality under mild assumptions.
5.1  Assumptions
We begin by outlining the essential technical assumptions required for our theoretical results.
Assumption 1 (Bounded support).
The support of the target distribution $q_{0}$ is a compact set contained within the hypercube $\{\mathbf{x}_{0}\in{\mathbb{R}}^{d}:\|\mathbf{x}_{0}\|_{\infty}\leq 1\}$.
Assumption 1 is a standard condition imposed on the data distribution (Lee et al., 2023a; Oko et al., 2023; Chang et al., 2025a; Beyler and Bach, 2025). This constraint is well-motivated by practical applications; for instance, image and video data consist of bounded pixel values, thereby satisfying this requirement.
Assumption 2 (Bounded weight function).
The weight function $w$ defined in (3.1) is bounded from above and bounded away from zero. Specifically, there exist constants $0<\underline{B}<1<\overline{B}<\infty$ such that
$\underline{B}\leq w(\mathbf{x})\leq\bar{B},\quad\text{for all}\penalty 10000\ \mathbf{x}\in\mathrm{supp}(q_{0}).$
Assumption 2 implies that the reference distribution $p_{0}$ and the tilted distribution $q_{0}$ satisfy mutual absolute continuity, ensuring that their supports coincide (i.e., $\mathrm{supp}(p_{0})=\mathrm{supp}(q_{0})$). This condition is crucial for establishing the regularity of Doob’s $h$-function. Furthermore, the ratio $\kappa\coloneq\bar{B}/\underline{B}$ serves as a condition number that characterizes the difficulty of the controllable diffusion task, as discussed in the context of posterior sampling by Purohit et al. (2025); Ding et al. (2024); Chang et al. (2025b).
Under Assumptions 1 and 2, we establish the regularity properties of Doob’s $h$-function defined in (3.10).
Proposition 5.1.
Suppose Assumptions 1 and 2 hold. Then for all $t\in(0,T)$ and $\mathbf{x}\in{\mathbb{R}}^{d}$, the following bounds hold:
  1. (i)
    $\underline{B}\leq h_{t}^{*}(\mathbf{x})\leq\bar{B}$;
  2. (ii)
    $\max_{1\leq k\leq d}|D_{k}h_{t}^{*}(\mathbf{x})|\leq 2\sigma_{T-t}^{-2}\bar{B}$; and
  3. (iii)
    $\max_{1\leq k,\ell\leq d}|D_{k\ell}^{2}h_{t}^{*}(\mathbf{x})|\leq 6\sigma_{T-t}^{-4}\bar{B}$,
where $D_{k}$ and $D_{k\ell}^{2}$ denote the first-order and second-order partial derivatives with respect to the input coordinates, respectively.
The proof of Proposition 5.1 is deferred to Appendix C. It is worth noting that Proposition 5.1 relies solely on the boundedness of the weight function $w$, without requiring the existence or smoothness of its gradients. Nevertheless, we establish that Doob’s $h$-function admits bounded derivatives. This result stems from the definition of Doob’s $h$-function as a posterior expectation under a Gaussian likelihood; the inherent smoothness of the Gaussian kernel endows the posterior expectation with strong regularity properties.
Assumption 3 (Reference score estimation error).
The reference score estimator $\widehat{\mathbf{s}}$ defined in (2.5) satisfies the following error bound:
$\frac{1}{T}\sum_{k=0}^{K-1}h{\mathbb{E}}^{{\mathbb{P}}}\left[\|\widehat{\mathbf{s}}(kh,\mathbf{X}_{kh}^{\leftarrow})-\nabla\log p_{T-kh}(\mathbf{X}_{kh}^{\leftarrow})\|_{2}^{2}\right]\leq\varepsilon_{\mathrm{ref}}^{2}.$
Assumption 3 requires the $L^{2}$-error of the reference score estimator $\widehat{\mathbf{s}}$ to be bounded with respect to the reference path measure $\mathbb{P}$}. In our setting, where numerous samples from the reference distribution $p_{0}$ are available, estimators satisfying this bound can be obtained via implicit score matching (Hyvärinen, 2005), sliced score matching (Song et al., 2020), or denoising score matching (Vincent, 2011). While one can derive explicit bounds of reference score matching as Tang and Yang (2024); Oko et al. (2023); Fu et al. (2024); Ding et al. (2025a); Yakovlev and Puchkin (2025a); Yakovlev et al. (2025) using non-parametric regression theory for deep neural networks (Bauer and Kohler, 2019; Schmidt-Hieber, 2020; Kohler and Langer, 2021; Jiao et al., 2023), we adopt this condition to maintain clarity of presentation, following the convention of Lee et al. (2023a); Chen et al. (2023a); Beyler and Bach (2025); Kremling et al. (2025).
5.2  Error bounds for the Doob’s guidance estimator
We begin by introducing the concept of Vapnik-Chervonenkis (VC) dimension (Vapnik and Chervonenkis, 1971; Anthony et al., 1999; Bartlett et al., 2019), which measures the complexity of a function class.
Definition 1 (VC-dimension).
Let $\mathscr{H}$} be a class of functions mapping from $\mathcal{X}$} to $\mathbb{R}$}. For any num-negative integer $m$, the growth function of $\mathscr{H}$} is defined as
$\Pi_{{\mathscr{H}}}(m)\coloneq\max_{x_{1},\ldots,x_{m}\in{\mathcal{X}}}|\{(\mathrm{sgn}\,h(x_{1}),\ldots,\mathrm{sgn}\,h(x_{m})):h\in{\mathscr{H}}\}|.$
We say $\mathscr{H}$} shatters the set $\{x_{1},\ldots,x_{m}\}$, if
$|\{(\mathrm{sgn}\,h(x_{1}),\ldots,\mathrm{sgn}\,h(x_{m})):h\in{\mathscr{H}}\}|=2^{m}.$
The Vapnik-Chervonenkis dimension of $\mathscr{H}$}, denoted by $\mathrm{VCdim}({\mathscr{H}})$, is the size of the largest shattered set, i.e., the largest $m$ such that $\Pi_{{\mathscr{H}}}(m)=2^{m}$.
To simplify notation, we define the gradient classes and their associated VC-dimensions. For a differentiable hypothesis class $\mathscr{H}$} consisting of functions mapping from ${\mathbb{R}}^{d}$ to $\mathbb{R}$}, the VC-dimension of the gradient hypothesis class is defined as
$\mathrm{VCdim}(\nabla{\mathscr{H}})\coloneq\max_{1\leq k\leq d}\mathrm{VCdim}(D_{k}{\mathscr{H}}),\quad D_{k}{\mathscr{H}}\coloneq\{D_{k}h:h\in{\mathscr{H}}\},$
where $D_{k}$ represents the derivative with respect to the $k$-th entry of the input.
The following lemma provides an oracle inequality for the variationally stable Doob’s matching (4.6).
Lemma 5.2 (Oracle inequality).
Suppose Assumptions 1 and 2 hold. Let $t\in(0,T)$ and let ${\mathscr{H}}_{t}$ be a hypothesis class. Let $\widehat{h}_{t}^{\lambda}$ be the gradient-regularized empirical risk minimizer defined as (4.6), and let $h_{t}^{*}$ be the Doob’s $h$-function defined as (3.10). Then the following inequalities hold:
${\mathbb{E}}\Big[\|\widehat{h}_{t}^{\lambda}-h_{t}^{*}\|_{L^{2}(p_{T-t})}^{2}\Big]$ $\lesssim\underbrace{\inf_{h_{t}\in{\mathscr{H}}_{t}}\Big\{\|h_{t}-h_{t}^{*}\|_{L^{2}(p_{T-t})}^{2}+\lambda\|\nabla h_{t}-\nabla h_{t}^{*}\|_{L^{2}(p_{T-t})}^{2}\Big\}}_{\text{(I)}}$
$\quad+\underbrace{\bar{B}^{2}\Big(\frac{\mathrm{VCdim}({\mathscr{H}}_{t})}{n\log^{-1}n}\Big)^{\frac{1}{2}}+\frac{\lambda d\bar{B}^{2}}{\sigma_{T-t}^{4}}\Big(\frac{\mathrm{VCdim}(\nabla{\mathscr{H}}_{t})}{n\log^{-1}n}\Big)^{\frac{1}{2}}}_{\text{(II)}}+\underbrace{\frac{\lambda^{2}d\bar{B}^{2}}{\sigma_{T-t}^{8}}}_{\text{(III)}},$
${\mathbb{E}}\Big[\|\nabla\widehat{h}_{t}^{\lambda}-\nabla h_{t}^{*}\|_{L^{2}(p_{T-t})}^{2}\Big]$ $\lesssim\underbrace{\inf_{h_{t}\in{\mathscr{H}}_{t}}\Big\{\frac{1}{\lambda}\|h_{t}-h_{t}^{*}\|_{L^{2}(p_{T-t})}^{2}+\|\nabla h_{t}-\nabla h_{t}^{*}\|_{L^{2}(p_{T-t})}^{2}\Big\}}_{\text{(I))}}$
$\quad+\underbrace{\frac{\bar{B}^{2}}{\lambda}\Big(\frac{\mathrm{VCdim}({\mathscr{H}}_{t})}{n\log^{-1}n}\Big)^{\frac{1}{2}}+\frac{d\bar{B}^{2}}{\sigma_{T-t}^{4}}\Big(\frac{\mathrm{VCdim}(\nabla{\mathscr{H}}_{t})}{n\log^{-1}n}\Big)^{\frac{1}{2}}}_{\text{(II)}}+\underbrace{\frac{\lambda d\bar{B}^{2}}{\sigma_{T-t}^{8}}}_{\text{(III)}},$
where the notation $\lesssim$ hides absolute constants.
The proof of Lemma 5.2 is deferred to Appendix D. Both oracle inequalities for Doob’s $h$-function and its gradient decompose the error into three components: approximation error, generalization error, and regularization gap.
  1. (I)
    The approximation error is defined as the minimal $H^{1}$-distance between functions in the hypothesis class ${\mathscr{H}}_{t}$ and the ground-truth Doob’s $h$-function $h_{t}^{*}$, measuring the approximation capability of ${\mathscr{H}}_{t}$.
  2. (II)
    The generalization error captures the error arising from finite-sample approximation, which vanishes as the number of samples approaches infinity.
  3. (III)
    The regularization gap is introduced by the gradient regularization in the objective functional, which causes the minimizer of the variationally stable objective (4.3) to deviate from the ground-truth Doob’s $h$-function $h_{t}^{*}$ (3.10). This gap has been analyzed in Proposition 4.2.
Comparison with oracle inequality of vanilla regression
Lemma 5.2 is analogous to the oracle inequality found in regression problems. Let $\widehat{h}_{t}$ be the vanilla estimator estimated by minimizing (4.5) over the hypothesis class ${\mathscr{H}}_{t}$. Informally, the following oracle inequality holds:
(5.1) ${\mathbb{E}}\Big[\|\widehat{h}_{t}-h_{t}^{*}\|_{L^{2}(p_{T-t})}^{2}\Big]\lesssim\underbrace{\inf_{h_{t}\in{\mathscr{H}}_{t}}\|h_{t}-h_{t}^{*}\|_{L^{2}(p_{T-t})}^{2}}_{\text{approximation}}+\underbrace{\bar{B}^{2}\Big(\frac{\mathrm{VCdim}({\mathscr{H}}_{t})}{n\log^{-1}n}\Big)^{\frac{1}{2}}}_{\text{generalization}}.$
Comparing (5.1) with Lemma 5.2 reveals several crucial differences:
  1. (i)
    The approximation error in Lemma 5.2 is measured in the $H^{1}$-norm, whereas in (5.1), it is measured in the $L^{2}$-norm. This distinction is natural because we require the estimator to converge in the $H^{1}$-norm; thus, simultaneous approximation of the function and its derivatives is essential. Simultaneous approximation using neural networks has been investigated in various contexts (Li et al., 2019; Gühring et al., 2020; Gühring and Raslan, 2021; Duan et al., 2022a, b; Lu et al., 2022b; Shen et al., 2022, 2024; Belomestny et al., 2023; Yakovlev and Puchkin, 2025b).
  2. (ii)
    In vanilla regression, the generalization error in (5.1) depends only on the complexity of the hypothesis class. In contrast, the generalization error bounds in Lemma 5.2 also depend on the complexity of the derivative classes $\nabla{\mathscr{H}}_{t}$. This occurs because the objective functional of the gradient-regularized Doob’s matching (4.3) includes the gradient norm term. Consequently, the error from finite-sample approximation is influenced not only by the complexity of the hypothesis class but also by that of the derivative class.
  3. (iii)
    The most significant difference lies in the regularization error. If we focus solely on the oracle inequality for $\widehat{h}_{t}^{\lambda}$, letting $\lambda$ go to zero reduces the expression to the vanilla regression oracle inequality (5.1). However, the bound for the gradient $\nabla\widehat{h}_{t}^{\lambda}$ diverges as the regularization parameter $\lambda$ approaches zero. This highlights the key advantage of our gradient-regularized method: the gradient-regularized is essential for guaranteeing simultaneous convergence of both the estimator value and its gradient. Additionally, there exists a trade-off with respect to $\lambda$ in the oracle inequality for $\widehat{h}_{t}^{\lambda}$: a larger $\lambda$ leads to larger regularization error, while reduces the approximation and generalization errors.
Given the oracle inequality for a general hypothesis class ${\mathscr{H}}_{t}$, we consider the specifical case that ${\mathscr{H}}_{t}$ is chosen as a neural network class, with the aim of deriving non-asymptotic convergence rates. We begin by formally defining the neural network class.
Definition 2 (Neural network class).
A function implemented by a neural network $h:{\mathbb{R}}^{N_{0}}\to{\mathbb{R}}^{N_{L+1}}$ is defined by
$h(\mathbf{x})=T_{L}(\varrho(T_{L-1}(\cdots\varrho(T_{0}(\mathbf{x}))\cdots))),$
where the activation function $\varrho$ is applied component-wise and $T_{\ell}(\mathbf{x}):=\mathbf{A}_{\ell}\mathbf{x}+\mathbf{b}_{\ell}$ is an affine transformation with $\mathbf{A}_{\ell}\in{\mathbb{R}}^{N_{\ell+1}\times N_{\ell}}$ and $\mathbf{b}_{\ell}\in{\mathbb{R}}^{N_{\ell}}$ for $\ell=0,\ldots,L$. In this paper, we consider the case where $N_{0}=d+1$ and $N_{L+1}=1$. The number $L$ is called the depth of neural networks. Additionally, $S:=\sum_{\ell=0}^{L}(\|\mathbf{A}_{\ell}\|_{0}+\|\mathbf{b}_{\ell}\|_{0})$ represents the total number of non-zero weights within the neural network. We denote by $N(L,S)$ the set of neural networks with depth at most $L$ and the number of non-zero weights at most $S$.
The following theorem establishes the convergence rate of the estimated Doob’s guidance given in (4.7).
Theorem 5.3 (Convergence rate of Doob’s guidance).
Suppose Assumptions 1 and 2 hold. Let $t\in(0,T)$. Set the hypothesis class ${\mathscr{H}}_{t}$ as
(5.2) ${\mathscr{H}}_{t}\coloneq\left\{h_{t}\in N(L,S):\begin{aligned} &\sup_{\mathbf{x}\in{\mathbb{R}}^{d}}h_{t}(\mathbf{x})\leq\bar{B},\penalty 10000\ \inf_{\mathbf{x}\in{\mathbb{R}}^{d}}h_{t}(\mathbf{x})\geq\underline{B},\\ &\max_{1\leq k\leq d}\sup_{\mathbf{x}\in{\mathbb{R}}^{d}}|D_{k}h_{t}(\mathbf{x})|\leq 2\sigma_{T-t}^{-2}\bar{B}\end{aligned}\right\},$
where $L={\mathcal{O}}(\log n)$ and $S={\mathcal{O}}(n^{\frac{d}{d+8}})$. Let $\widehat{h}_{t}^{\lambda}$ be the gradient-regularized empirical risk minimizer defined as (4.6), and let $h_{t}^{*}$ be the Doob’s $h$-function defined as (3.10). Then the following inequality holds:
${\mathbb{E}}\Big[\|\nabla\log\widehat{h}_{t}^{\lambda}-\nabla\log h_{t}^{*}\|_{L^{2}(p_{T-t})}^{2}\Big]\leq C\sigma_{T-t}^{-8}n^{-\frac{2}{d+8}}\log^{4}n,$
provided that the regularization parameter $\lambda$ is set as $\lambda={\mathcal{O}}(n^{-\frac{2}{d+8}})$, where $C$ is a constant depending only on $d$, $\bar{B}$, and $\underline{B}$.
The proof of Theorem 5.3 is deferred to Appendix D. This theorem demonstrates that the $L^{2}$-error of the Doob’s guidance estimator (4.7) converges to the exact Doob’s guidance in (3.14) as the sample size increases, provided that the size of the neural network is appropriately chosen. However, since the prefactor $\sigma_{T-t}^{-8}$ diverges as $t$ approaches the terminal time $T$, early stopping in controllable diffusion models (4.8) is required to ensure the validity of the Doob’s guidance estimator.
Remark 5.4 (Comparisons with previous work).
Simultaneous estimation of a function and its gradient using deep neural network has been investigated by Ding et al. (2025b). The most important distinction in our work lies in the elimination of the convexity assumption on the hypothesis class. Specifically, Ding et al. (2025b, Lemma 7) propose an oracle inequality under the assumption that the hypothesis class is convex. Furthermore, Ding et al. (2025b, Theorem 3) requires the estimator to be a minimizer of the gradient-regularized empirical risk over the convex hull of a neural network class, which is intractable in practice. In contrast, Lemma 5.2 eliminates the requirement of convexity for the hypothesis class, and Theorem 5.3 removes the need for the convex hull of the neural network class. This aligns the theoretical analysis more closely with practical computing.
5.3  Error bounds for the controllable diffusion models
In this subsection, we establish a non-asymptotic convergence rate for the controllable diffusion models (4.8). We begin by presenting an error decomposition for the KL-divergence between the early-stopping distribution $q_{T_{0}}$ and the distribution of $\widehat{\mathbf{Z}}_{T-T_{0}}^{\leftarrow}$.
Lemma 5.5 (Error decomposition).
Suppose Assumptions 1, 2, and 3 hold. Let $\widehat{q}_{T-T_{0}}$ be the marginal density of $\widehat{\mathbf{Z}}_{T-T_{0}}^{\leftarrow}$ defined in (4.8). Then it follows that
$\mathrm{KL}(q_{T_{0}}\|\widehat{q}_{T-T_{0}})$ $\lesssim\underbrace{\frac{\bar{B}}{\underline{B}}\sum_{k=0}^{K-1}h{\mathbb{E}}^{{\mathbb{P}}}\left[\|\nabla\log\widehat{h}_{kh}(\mathbf{X}_{kh}^{\leftarrow})-\nabla\log h_{kh}^{*}(\mathbf{X}_{kh}^{\leftarrow})\|_{2}^{2}\right]}_{\text{(I)}}$
$\quad+\underbrace{\frac{\bar{B}}{\underline{B}}T\varepsilon_{\mathrm{ref}}^{2}}_{\text{(II)}}+\underbrace{d\exp(-T)}_{\text{(III)}}+\underbrace{\frac{d^{2}T^{2}}{\sigma_{T_{0}}^{4}K}}_{\text{(IV)}},$
where the notation $\lesssim$ hides absolute constants.
The proof of Lemma 5.5 is deferred to Appendix E. Lemma 5.5 decomposes the KL-divergence between the early-stopping distribution $q_{T_{0}}$ and the distribution of $\widehat{\mathbf{Z}}_{T-T_{0}}^{\leftarrow}$ into four components: (I) Doob’s guidance error, (II) the reference score error, (III) the initialization error, and (IV) the discretization error. Specifically, Doob’s guidance error represents the average error of Doob’s $h$-guidance estimator at each time point, which has been investigated in Theorem 5.3; the reference score error is the average error of the reference score estimator at each time point, which is discussed in Assumption 3; the initialization error arises from replacing the initial distribution $q_{T}=p_{T}$ with a Gaussian distribution in (4.8); and the discretization error is induced by the exponential integrator.
While Lemma 5.5 characterizes the error between the estimated distribution $\widehat{q}_{T-T_{0}}$ and the early-stopping distribution $q_{T_{0}}$, our primary interest lies in the discrepancy between $\widehat{q}_{T-T_{0}}$ and the target tilted distribution $q_{0}$ defined in (3.1). Since the KL-divergence does not satisfy the triangular inequality, we instead propose an error bound in 2-Wasserstein distance. The following theorem establishes the 2-Wasserstein distance between the scaled and truncated distribution $({\mathcal{M}}\circ{\mathcal{T}}_{R})_{\sharp}\widehat{q}_{T-T_{0}}$ defined in (4.9) and the target tilted distribution $q_{0}$.
Theorem 5.6 (Convergence rate of controllable diffusion models).
Suppose Assumptions 1, 2, and 3 hold. Let $\varepsilon\in(0,1)$. Set the hypothesis classes $\{{\mathscr{H}}_{T-kh}\}_{k=0}^{K=1}$ as (5.2) with the same depth $L$ and number of non-zero parameters $S$ as Theorem 5.3. Let $\widehat{q}_{T-T_{0}}$ be the marginal density of $\widehat{\mathbf{Z}}_{T-T_{0}}^{\leftarrow}$ defined in (4.8), and let $({\mathcal{M}}\circ{\mathcal{T}}_{R})_{\sharp}\widehat{q}_{T-T_{0}}$ defined as (4.9). Then it follows that
(5.3) ${\mathbb{E}}\Big[{\mathcal{W}}_{2}^{2}(q_{0},({\mathcal{M}}\circ{\mathcal{T}}_{R})_{\sharp}\widehat{q}_{T-T_{0}})\Big]\leq C\varepsilon\log^{3}\Big(\frac{1}{\varepsilon}\Big),$
provided that the truncation radius $R$, the terminal time $T$, the step size $h$, the number of steps $K$, the error of reference score $\varepsilon_{\mathrm{ref}}$, the number of samples $n$ for Doob’s matching, and the early-stopping time $T_{0}$ are set, respectively, as
$R\asymp\log^{\frac{1}{2}}\Big(\frac{1}{\varepsilon}\Big),\quad T\asymp\log\Big(\frac{1}{\varepsilon^{2}}\Big),\quad K\gtrsim\frac{1}{\varepsilon^{4}}\log^{2}\Big(\frac{1}{\varepsilon^{2}}\Big),\quad h\lesssim\varepsilon^{4}\log^{-1}\Big(\frac{1}{\varepsilon^{2}}\Big)$
$\varepsilon_{\mathrm{ref}}^{2}\lesssim\varepsilon^{2}\log^{-1}\Big(\frac{1}{\varepsilon^{2}}\Big),\quad n\gtrsim\frac{1}{\varepsilon^{3(d+8)}}\log^{\frac{d+8}{2}}\Big(\frac{1}{\varepsilon^{2}}\Big).$
Here $C$ is a constant depending only on $d$, $\bar{B}$, and $\underline{B}$.
The proof of Theorem 5.6 is deferred to Appendix E. This theorem establishes a non-asymptotic convergence rate for the controllable diffusion model (4.8) using the variationally stable Doob’s matching (4.6). Crucially, it provides theoretical guidance for selecting hyper-parameters, including the truncation radius $R$ in (4.9), the step size $h$, the number of steps $K$ in (4.8), the early stopping time $T_{0}$, the terminal time $T$, the reference score error $\varepsilon_{\mathrm{ref}}$ (Assumption 3), and the sample size $n$ for Doob’s matching in (4.4).
However, this rate suffers from the curse of dimensionality (CoD), implying that the required number of samples $n$ grows exponentially as the error tolerance $\varepsilon$ decays. We address this challenge in the remainder of this section under a low-dimensional subspace assumption.
5.4  Adaptivity to low-dimensionality
In this subsection, we demonstrate that the convergence rate mitigates the curse of dimensionality under a low-dimensional subspace assumption, a setting previously explored in Chen et al. (2023b, Section 3) and Oko et al. (2023, Section 6).
Assumption 4 (Low-dimensional subspace).
Let $d^{*}\ll d$ be an integer, and $\mathbf{P}\in{\mathbb{R}}^{d\times d^{*}}$ be a column orthogonal matrix, i.e., $\mathbf{P}^{\top}\mathbf{P}=\mathbf{I}_{d^{*}}$. Let $\bar{p}_{0}$ be a probability density with a compact support contained within a hypercube $\{\bar{\mathbf{x}}_{0}\in{\mathbb{R}}^{d^{*}}:\|\bar{\mathbf{x}}_{0}\|_{\infty}\leq 1\}$. The reference density $p_{0}$ is a push-forward of $\bar{p}_{0}$ by the linear map $\mathbf{P}$, i.e., $p_{0}\coloneq\mathbf{P}_{\sharp}\bar{p}_{0}$.
Consequently, the reference density $p_{0}$ is supported on a linear subspace $\{\mathbf{P}\bar{\mathbf{x}}_{0}\in{\mathbb{R}}^{d}:\bar{\mathbf{x}}_{0}\in{\mathbb{R}}^{d^{*}}\}$ with an ambient dimension $d$, and a much smaller intrinsic dimension $d^{*}\ll d$.
Before proceeding, we define the forward and time-reversal process in the low-dimensional latent space. Analogously to (2.1), the forward process reads
(5.4) $\mathrm{d}\bar{\mathbf{X}}_{t}=-\bar{\mathbf{X}}_{t}\,\mathrm{d}t+\sqrt{2}\,\mathrm{d}\bar{\mathbf{B}}_{t},\quad t\in(0,T),\penalty 10000\ \bar{\mathbf{X}}_{0}\sim\bar{p}_{0},$
where $\bar{\mathbf{B}}_{t}$ is a $d^{*}$-dimensional standard Brownian motion, and $T>0$ is the terminal time. The transition distribution of this forward process is given by:
(5.5) $(\bar{\mathbf{X}}_{t}|\bar{\mathbf{X}}_{0}=\bar{\mathbf{x}}_{0})\sim\mathcal{N}(\mu_{t}\bar{\mathbf{x}}_{0},\sigma_{t}^{2}\mathbf{I}_{d^{*}}).$
Let $\bar{p}_{t}$ denote the marginal density of $\bar{\mathbf{X}}_{t}$ for $t\in(0,T)$. The corresponding time-reversal process (Anderson, 1982) is defined as
(5.6) $\,\mathrm{d}\bar{\mathbf{X}}_{t}^{\leftarrow}$ $=(\bar{\mathbf{X}}_{t}^{\leftarrow}+2\nabla\log\bar{p}_{T-t}(\bar{\mathbf{X}}_{t}^{\leftarrow}))\,\mathrm{d}t+\sqrt{2}\,\mathrm{d}\bar{\mathbf{B}}_{t},\quad t\in(0,T),$
$\bar{\mathbf{X}}_{0}^{\leftarrow}$ $\sim\bar{p}_{T}.$
As shown by Anderson (1982), the path measure of the time-reversal process $(\bar{\mathbf{X}}_{t}^{\leftarrow})_{0\leq t\leq T}$ corresponds exactly to the reverse of the forward process $(\bar{\mathbf{X}}_{t})_{0\leq t\leq T}$.
The following result establishes a relationship between the ground-truth Doob’s $h$-function (3.10) and its analogue $\bar{h}_{t}^{*}:{\mathbb{R}}^{d^{*}}\to{\mathbb{R}}$ in the low-dimensional latent space. In other words, it provides a low-dimensional representation of the ground-truth Doob’s $h$-function.
Proposition 5.7 (Low-dimensional representation).
Suppose Assumptions 4 and 2 hold. Then for any $t\in(0,T)$ and $\mathbf{x}\in{\mathbb{R}}^{d}$, we have
(5.7) $h_{t}^{*}(\mathbf{x})=\bar{h}_{t}^{*}(\mathbf{P}^{\top}\mathbf{x})\coloneq{\mathbb{E}}[w(\mathbf{P}\bar{\mathbf{X}}_{T}^{\leftarrow})\mid\bar{\mathbf{X}}_{t}^{\leftarrow}=\mathbf{P}^{\top}\mathbf{x}].$
The proof of Proposition 5.7 is provided in Appendix F. Proposition 5.7 implies that estimating the ground-truth Doob’s $h$-function reduces to estimating its low-dimensional counterpart $\bar{h}_{t}^{*}:{\mathbb{R}}^{d^{*}}\rightarrow{\mathbb{R}}$, thereby enabling the Doob’s guidance estimator to adapt to low-dimensional structures.
Analogously to Proposition 5.1, we can establish the regularity properties of the low-dimensional representation of Doob’s $h$-function $\bar{h}_{t}^{*}:{\mathbb{R}}^{d^{*}}\to{\mathbb{R}}$ in (5.7).
Proposition 5.8.
Suppose Assumptions 4 and 2 hold. Then for all $t\in(0,T)$ and $\bar{\mathbf{x}}\in{\mathbb{R}}^{d^{*}}$, the following bounds hold:
  1. (i)
    $\underline{B}\leq\bar{h}_{t}^{*}(\bar{\mathbf{x}})\leq\bar{B}$;
  2. (ii)
    $\max_{1\leq k\leq d}|D_{k}\bar{h}_{t}^{*}(\bar{\mathbf{x}})|\leq 2\sigma_{T-t}^{-2}\bar{B}$; and
  3. (iii)
    $\max_{1\leq k,\ell\leq d}|D_{k\ell}^{2}\bar{h}_{t}^{*}(\bar{\mathbf{x}})|\leq 6\sigma_{T-t}^{-4}\bar{B}$,
where $D_{k}$ and $D_{k\ell}^{2}$ denote the first-order and second-order partial derivatives with respect to the input coordinates, respectively.
The proof of Proposition 5.7 is provided in Appendix F. Based on these results, we derive the convergence rates for the variationally stable Doob’s matching under the assumption of low-dimensional subspace.
Theorem 5.9 (Adaptivity to intrinsic dimension).
Suppose Assumptions 4 and 2 hold. Let $t\in(0,T)$. Set the hypothesis class ${\mathscr{H}}_{t}$ as
(5.8) ${\mathscr{H}}_{t}\coloneq\left\{h_{t}\in N(L,S):\begin{aligned} &\sup_{\mathbf{x}\in{\mathbb{R}}^{d}}h_{t}(\mathbf{x})\leq\bar{B},\penalty 10000\ \inf_{\mathbf{x}\in{\mathbb{R}}^{d}}h_{t}(\mathbf{x})\geq\underline{B},\\ &\max_{1\leq k\leq d}\sup_{\mathbf{x}\in{\mathbb{R}}^{d}}|D_{k}h_{t}(\mathbf{x})|\leq 2\sigma_{T-t}^{-2}\bar{B}\end{aligned}\right\},$
where $L={\mathcal{O}}(\log n)$ and $S={\mathcal{O}}(n^{\frac{d^{*}}{d^{*}+8}})$. Let $\widehat{h}_{t}^{\lambda}$ be the gradient-regularized empirical risk minimizer defined as (4.6), and let $h_{t}^{*}$ be the Doob’s $h$-function defined as (3.10). Then the following inequality holds:
${\mathbb{E}}\Big[\|\nabla\log\widehat{h}_{t}^{\lambda}-\nabla\log h_{t}^{*}\|_{L^{2}(p_{T-t})}^{2}\Big]\leq C\sigma_{T-t}^{-8}n^{-\frac{2}{d^{*}+8}}\log^{4}n,$
provided that the regularization parameter $\lambda$ is set as $\lambda={\mathcal{O}}(n^{-\frac{2}{d^{*}+8}})$, where $C$ is a constant depending only on $d^{*}$, $\bar{B}$, and $\underline{B}$.
The proof of Theorem 5.9 is provided in Appendix F. This result confirms that the convergence rate eliminates the exponential dependence on the ambient dimension $d$, scaling exponentially solely with the intrinsic dimension $d^{*}\ll d$. This effectively mitigates the curse of dimensionality in Theorem 5.3.
The following corollary is a direct consequence of Theorem 5.9, derived using arguments similar to those in Theorem 5.6.
Corollary 5.10.
Suppose Assumptions 4, 2, and 3 hold. Let $\varepsilon\in(0,1)$. Set the hypothesis classes $\{{\mathscr{H}}_{T-kh}\}_{k=0}^{K=1}$ as (5.2) with the same depth $L$ and number of non-zero parameters $S$ as Theorem 5.3. Let $\widehat{q}_{T-T_{0}}$ be the marginal density of $\widehat{\mathbf{Z}}_{T-T_{0}}^{\leftarrow}$ defined in (4.8), and let $({\mathcal{M}}\circ{\mathcal{T}}_{R})_{\sharp}\widehat{q}_{T-T_{0}}$ defined as (4.9). Then it follows that
(5.9) ${\mathbb{E}}\Big[{\mathcal{W}}_{2}^{2}(q_{0},({\mathcal{M}}\circ{\mathcal{T}}_{R})_{\sharp}\widehat{q}_{T-T_{0}})\Big]\leq C\varepsilon\log^{3}\Big(\frac{1}{\varepsilon}\Big).$
provided that the truncation radius $R$, the terminal time $T$, the step size $h$, the number of steps $K$, the error of reference score $\varepsilon_{\mathrm{ref}}$, the number of samples $n$ for Doob’s matching, and the early-stopping time $T_{0}$ are set, respectively, as
$R\asymp\log^{\frac{1}{2}}\Big(\frac{1}{\varepsilon}\Big),\quad T\asymp\log\Big(\frac{1}{\varepsilon^{2}}\Big),\quad K\gtrsim\frac{1}{\varepsilon^{4}}\log^{2}\Big(\frac{1}{\varepsilon^{2}}\Big),\quad h\lesssim\varepsilon^{4}\log^{-1}\Big(\frac{1}{\varepsilon^{2}}\Big)$
$\varepsilon_{\mathrm{ref}}^{2}\lesssim\varepsilon^{2}\log^{-1}\Big(\frac{1}{\varepsilon^{2}}\Big),\quad n\gtrsim\frac{1}{\varepsilon^{3(d^{*}+8)}}\log^{\frac{d+8}{2}}\Big(\frac{1}{\varepsilon^{2}}\Big).$
Here $C$ is a constant depending only on $^{*}, dd$, $\bar{B}$, and $\underline{B}$.
Crucially, the convergence rates in Corollary 5.10 depend only polynomially on the ambient dimension $d$, while the sample complexity depends exponentially solely on the intrinsic dimension $d^{*}\ll d$. This result significantly mitigates the curse of dimensionality.
Remark 5.11.
Adaptivity to low dimensionality plays a pivotal role in the analysis of diffusion and flow-based models. One line of work studies the adaptivity of score or velocity estimator to low dimensionality under Assumption 4 or its variants; see e.g., Chen et al. (2023b); Oko et al. (2023); Yakovlev and Puchkin (2025a); Ding et al. (2025a). A second line of work focuses on the adaptivity of the sampling procedure; see e.g., Li and Yan (2024); Huang et al. (2024); Potaptchik et al. (2025). These works provide valuable insights for extending Corollary 5.10 to achieve provable adaptivity across reference score estimation, sampling, and guidance estimation, thereby completely eliminating dependence on the ambient dimension. We leave this unified analysis, which is outside the scope of the current work, for future research. Importantly, even within the current framework, the dependence on the ambient dimension $d$ remains only polynomial.
6  Conclusions
In this work, we proposed variationally stable Doob’s matching, a principled inference-time alignment framework for diffusion models grounded in the theory of Doob’s $h$-transform. Our approach reformulates guidance as the gradient of the logarithm of an underlying Doob’s $h$-function, providing a mathematically consistent mechanism for tilting a pre-trained diffusion model toward a target distribution without retraining the reference score network. By leveraging gradient-regularized regression, Doob’s matching simultaneously estimates both the $h$-function and its gradient, thereby providing a consistent estimator for Doob’s guidance.
From a theoretical perspective, we established non-asymptotic convergence rates for the proposed guidance estimator, showing that the estimated Doob’s guidance converges to the true guidance under suitable choices of the hypothesis class and regularization parameter. Building on this result, we further derived non-asymptotic convergence guarantees for the induced controllable diffusion process, demonstrating that the generated distribution converges to the target distribution in the $2$-Wasserstein distance. These results provide an end-to-end theoretical guarantees for inference-time aligned diffusion models that explicitly account for guidance estimation error, reference score estimation error, initialization bias, and discretization error. Furthermore, we show that our convergence rates depend solely on the intrinsic dimension of the linear subspace rather than the ambient dimension. This highlights the estimator’s adaptivity to low-dimensional structures, effectively mitigating the curse of dimensionality.
Appendix A Derivations in Section 3
Lemma A.1.
The Doob’s $h$-transform $h^{*}(t,\mathbf{X}_{t}^{\leftarrow})$ defined as (3.10) is a martingale, and satisfies the following SDE:
$\mathrm{d}h^{*}(t,\mathbf{X}_{t}^{\leftarrow})=\nabla h^{*}(t,\mathbf{X}_{t}^{\leftarrow})^{\top}\sqrt{2}\,\mathrm{d}\mathbf{B}_{t}.$
Proof of Lemma A.1.
This proof is divided into two parts.
Part 1. Martingale. Due to the Markov property of the diffusion process $\mathbf{X}_{t}^{\leftarrow}$ [Øksendal, 2003, Theorem 7.1.2], using (3.10) implies
$M_{t}\coloneq h^{*}(t,\mathbf{X}_{t}^{\leftarrow})={\mathbb{E}}^{{\mathbb{P}}}[w(\mathbf{X}_{T}^{\leftarrow})|{\mathcal{F}}_{t}].$
It is apparent that $M_{t}$ is ${\mathcal{F}}_{t}$-measurable for each $t\in(0,T)$, thus $M_{t}$ is adapted to $\mathbb{F}$}. Then we show that $M_{t}$ is integrable under $\mathbb{P}$}. Indeed,
${\mathbb{E}}^{{\mathbb{P}}}[|M_{t}|]\leq{\mathbb{E}}^{{\mathbb{P}}}[{\mathbb{E}}^{{\mathbb{P}}}[|w(\mathbf{X}_{T}^{\leftarrow})||{\mathcal{F}}_{t}]]={\mathbb{E}}^{{\mathbb{P}}}[w(\mathbf{X}_{T}^{\leftarrow})]=Z<\infty,$
where the first inequality holds from Jensen’s inequality, and the first equality used the law of total expectation and the fact that $w(\mathbf{X}_{T}^{\leftarrow})$. We next show the martingale property. For each $0\leq s\leq t\leq T$,
${\mathbb{E}}^{{\mathbb{P}}}[M_{t}|{\mathcal{F}}_{s}]={\mathbb{E}}^{{\mathbb{P}}}[{\mathbb{E}}^{{\mathbb{P}}}[w(\mathbf{X}_{T}^{\leftarrow})|{\mathcal{F}}_{t}]|{\mathcal{F}}_{s}]={\mathbb{E}}^{{\mathbb{P}}}[w(\mathbf{X}_{T}^{\leftarrow})|{\mathcal{F}}_{s}]=M_{s},$
where the second equality involves the tower property of conditional expectation, and the fact that ${\mathcal{F}}_{s}\subseteq{\mathcal{F}}_{t}$. Therefore, $M_{t}$ is a martingale.
Part 2. Stochastic dynamics. Applying Itô’s formula to $h^{*}(t,\mathbf{X}_{t}^{\leftarrow})$ yields
$\,\mathrm{d}h^{*}(t,\mathbf{X}_{t}^{\leftarrow})$ $=\partial_{t}h^{*}(t,\mathbf{X}_{t}^{\leftarrow})\,\mathrm{d}t+\nabla h^{*}(t,\mathbf{X}_{t}^{\leftarrow})^{\top}\,\mathrm{d}\mathbf{X}_{t}^{\leftarrow}+\frac{1}{2}(\,\mathrm{d}\mathbf{X}_{t}^{\leftarrow})^{\top}\nabla^{2}h^{*}(t,\mathbf{X}_{t}^{\leftarrow})\,\mathrm{d}\mathbf{X}_{t}^{\leftarrow}$
$=\nabla h^{*}(t,\mathbf{X}_{t}^{\leftarrow})^{\top}\sqrt{2}\,\mathrm{d}\mathbf{B}_{t},$
where the last equality holds from the fact that martingale has zero drift. This completes the proof. ∎
Proposition 3.1.
The Radon-Nikodym derivative process $L_{t}$, as defined in (3.8), admits the following representation:
$L_{t}={\mathbb{E}}^{{\mathbb{P}}}\Big[\frac{\,\mathrm{d}{\mathbb{Q}}}{\,\mathrm{d}{\mathbb{P}}}\Big|{\mathcal{F}}_{t}\Big]=\frac{h^{*}(t,\mathbf{X}_{t}^{\leftarrow})}{{\mathbb{E}}^{{\mathbb{P}}}[w(\mathbf{X}_{T}^{\leftarrow})]},$
where $h^{*}:[0,T]\times{\mathbb{R}}^{d}\rightarrow{\mathbb{R}}$, referred to as Doob’s $h$-function, is defined as the conditional expectation of the terminal weight:
(A.1) $h^{*}(t,\mathbf{x})\coloneq{\mathbb{E}}^{{\mathbb{P}}}[w(\mathbf{X}_{T}^{\leftarrow})\mid\mathbf{X}_{t}^{\leftarrow}=\mathbf{x}].$
Furthermore, the log-likelihood ratio satisfies the following SDE:
$\mathrm{d}(\log L_{t})=\nabla\log h^{*}(t,\mathbf{X}_{t}^{\leftarrow})^{\top}\sqrt{2}\,\mathrm{d}\mathbf{B}_{t}-\|\nabla\log h^{*}(t,\mathbf{X}_{t}^{\leftarrow})\|_{2}^{2}\,\mathrm{d}t.$
Proof of Proposition 3.1.
The proof is divided into three parts.
Part 1. The equivalent definition of $L_{t}$. In this part, we aim to show the conditional expectation in Proposition 3.1 is identical to the likelihood $L_{t}$ defined as (3.8). Define
$\bar{L}_{t}={\mathbb{E}}^{{\mathbb{P}}}\Big[\frac{\,\mathrm{d}{\mathbb{Q}}}{\,\mathrm{d}{\mathbb{P}}}\Big|{\mathcal{F}}_{t}\Big].$
For each event $A\in{\mathcal{F}}_{t}\subseteq{\mathcal{F}}_{T}$, we have
${\mathbb{Q}}(A)=\int_{A}\frac{\,\mathrm{d}{\mathbb{Q}}}{\,\mathrm{d}{\mathbb{P}}}\,\mathrm{d}{\mathbb{P}}=\int_{A}{\mathbb{E}}\Big[\frac{\,\mathrm{d}{\mathbb{Q}}}{\,\mathrm{d}{\mathbb{P}}}\Big|{\mathcal{F}}_{t}\Big]\,\mathrm{d}{\mathbb{P}}=\int_{A}\bar{L}_{t}\,\mathrm{d}{\mathbb{P}},$
where the second equality is due to $A\in{\mathcal{F}}_{t}$. This means $\bar{L}_{t}$ acts as the density for the measure $\mathbb{Q}$} with respect to $\mathbb{P}$} when restricting to the $\sigma$-algebra ${\mathcal{F}}_{t}$. Thus $L_{t}\equiv\bar{L}_{t}$ for each $t\in(0,T)$.
Part 2. The expression of Doob’s $h$-transform. It is straightforward that
$L_{t}={\mathbb{E}}^{{\mathbb{P}}}\Big[\frac{\,\mathrm{d}{\mathbb{Q}}}{\,\mathrm{d}{\mathbb{P}}}\Big|{\mathcal{F}}_{t}\Big]=\frac{{\mathbb{E}}^{{\mathbb{P}}}[w(\mathbf{X}_{T}^{\leftarrow})|{\mathcal{F}}_{t}]}{{\mathbb{E}}^{{\mathbb{P}}}[w(\mathbf{X}_{T}^{\leftarrow})]}=\frac{{\mathbb{E}}^{{\mathbb{P}}}[w(\mathbf{X}_{T}^{\leftarrow})|\mathbf{X}_{t}^{\leftarrow}]}{{\mathbb{E}}^{{\mathbb{P}}}[w(\mathbf{X}_{T}^{\leftarrow})]}=\frac{h^{*}(t,\mathbf{X}_{t}^{\leftarrow})}{{\mathbb{E}}^{{\mathbb{P}}}[w(\mathbf{X}_{T}^{\leftarrow})]},$
where the second equality follows from (3.9), the third equality involves the Markovity of the diffusion process $\mathbf{X}_{t}^{\leftarrow}$, and the last equality is due to the definition of the $h$-function (3.10).
Part 3. The stochastic dynamics of log-likelihood. We can now derive the dynamics of the likelihood $L_{t}$. Letting $Z={\mathbb{E}}^{{\mathbb{P}}}[w(\mathbf{X}_{T}^{\leftarrow})]$, we have:
$\,\mathrm{d}L_{t}$ $=\frac{1}{Z}\,\mathrm{d}h^{*}(t,\mathbf{X}_{t}^{\leftarrow})$
$=\frac{1}{Z}(\nabla h^{*}(t,\mathbf{X}_{t}^{\leftarrow}))^{\top}\sqrt{2}\,\mathrm{d}\mathbf{B}_{t}$
$=\frac{h^{*}(t,\mathbf{X}_{t}^{\leftarrow})}{Z}\frac{\nabla h^{*}(t,\mathbf{X}_{t}^{\leftarrow})^{\top}}{h^{*}(t,\mathbf{X}_{t}^{\leftarrow})}\sqrt{2}\,\mathrm{d}\mathbf{B}_{t}$
(A.2) $=L_{t}(\nabla\log h^{*}(t,\mathbf{X}_{t}^{\leftarrow}))^{\top}\sqrt{2}\,\mathrm{d}\mathbf{B}_{t},$
where the second equality is due to Lemma A.1. Using Itô’s formula for log-likelihood yields
$\,\mathrm{d}(\log L_{t})$ $=\frac{\,\mathrm{d}L_{t}}{L_{t}}-\frac{1}{2}\Big\langle\frac{\,\mathrm{d}L_{t}}{L_{t}},\frac{\,\mathrm{d}L_{t}}{L_{t}}\Big\rangle$
$=\sqrt{2}\nabla\log h^{*}(t,\mathbf{X}_{t}^{\leftarrow})^{\top}\,\mathrm{d}\mathbf{B}_{t}-\frac{1}{2}\|\sqrt{2}\nabla\log h^{*}(t,\mathbf{X}_{t}^{\leftarrow})\|_{2}^{2}\,\mathrm{d}t$
$=\sqrt{2}\nabla\log h^{*}(t,\mathbf{X}_{t}^{\leftarrow})^{\top}\,\mathrm{d}\mathbf{B}_{t}-\|\nabla\log h^{*}(t,\mathbf{X}_{t}^{\leftarrow})\|_{2}^{2}\,\mathrm{d}t,$
where the second equality is due to (A.2). This completes the proof. ∎
Proposition 3.2.
Let the reference process $\mathbf{X}_{t}^{\leftarrow}$ satisfy the SDE (2.4) under the path measure $\mathbb{P}$}, and let $\mathbb{Q}$} be the target path measure defined by (3.9). Assume that the Novikov condition holds:
(A.3) ${\mathbb{E}}^{{\mathbb{P}}}\Big[\exp\Big(\int_{0}^{T}\|\nabla\log h^{*}(s,\mathbf{X}_{s}^{\leftarrow})\|_{2}^{2}\,\mathrm{d}s\Big)\Big]<\infty.$
Define a process $(\widetilde{\mathbf{B}}_{t})_{1\leq t\leq T}$ by
(A.4) $\mathrm{d}\widetilde{\mathbf{B}}_{t}=\,\mathrm{d}\mathbf{B}_{t}-\sqrt{2}\nabla\log h^{*}(t,\mathbf{X}_{t}^{\leftarrow})\,\mathrm{d}t,$
where $\mathbf{B}_{t}$ is a Brownian motion under $\mathbb{P}$}. Then $\widetilde{\mathbf{B}}_{t}$ is a standard Brownian motion under $\mathbb{Q}$}. Further, under the path measure $\mathbb{Q}$}, the reference process $\mathbf{X}_{t}^{\leftarrow}$ in (2.4) evolves according to:
(A.5) $\mathrm{d}\mathbf{X}_{t}^{\leftarrow}=(\mathbf{X}_{t}^{\leftarrow}+2\nabla\log p_{T-t}(\mathbf{X}_{t}^{\leftarrow})+2\nabla\log h^{*}(t,\mathbf{X}_{t}^{\leftarrow}))\,\mathrm{d}t+\sqrt{2}\,\mathrm{d}\widetilde{\mathbf{B}}_{t}.$
Proof of Proposition 3.2.
The derivation proceeds in two steps.
Step 1. The Brownian motion under the target path measure $\mathbb{Q}$}. Using Proposition 3.1 and noting that $L_{0}\equiv 1$ as (3.9), we have
$L_{t}=\exp\Big(\int_{0}^{t}\nabla\log h^{*}(s,\mathbf{X}_{s}^{\leftarrow})^{\top}\sqrt{2}\,\mathrm{d}\mathbf{B}_{s}-\|\nabla\log h^{*}(s,\mathbf{X}_{s}^{\leftarrow})\|_{2}^{2}\,\mathrm{d}s\Big).$
According to Karatzas and Shreve [1998, Corollary 5.13], the Novikov condition (3.11) implies that $L_{t}$ is a martingale. Then applying Girsanov’s theorem [Øksendal, 2003, Theorem 8.6.6] yields that $\widetilde{\mathbf{B}}_{t}$ is a Brownian motion under $\mathbb{Q}$}.
Step 2. Dynamics under the target path measure. Recall that under the reference path measure $\mathbb{P}$}, the process evolves as
$\mathrm{d}\mathbf{X}_{t}^{\leftarrow}=(\mathbf{X}_{t}^{\leftarrow}+2\nabla\log p_{T-t}(\mathbf{X}_{t}^{\leftarrow}))\,\mathrm{d}t+\sqrt{2}\,\mathrm{d}\mathbf{B}_{t},$
where $\mathbf{B}_{t}$ is a $\mathbb{P}$}-Brownian motion. Substituting the relationship (3.12) into this SDE implies
$\,\mathrm{d}\mathbf{X}_{t}^{\leftarrow}$ $=(\mathbf{X}_{t}^{\leftarrow}+2\nabla\log p_{T-t}(\mathbf{X}_{t}^{\leftarrow}))\,\mathrm{d}t+\sqrt{2}(\,\mathrm{d}\widetilde{\mathbf{B}}_{t}+\sqrt{2}\nabla\log h^{*}(t,\mathbf{X}_{t}^{\leftarrow})\,\mathrm{d}t)$
$=(\mathbf{X}_{t}^{\leftarrow}+2\nabla\log p_{T-t}(\mathbf{X}_{t}^{\leftarrow})+2\nabla\log h^{*}(t,\mathbf{X}_{t}^{\leftarrow}))\,\mathrm{d}t+\sqrt{2}\,\mathrm{d}\widetilde{\mathbf{B}}_{t},$
which is the dynamics of $\mathbf{X}_{t}^{\leftarrow}$ under the target path measure $\mathbb{Q}$}. This completes the proof. ∎
Appendix B Derivations in Section 4
Proposition 4.1.
For every $t\in(0,T)$, the Doob’s $h$-function $h_{t}^{*}$ in (3.10) minimizes the implicit Doob’s matching loss (4.1). Further,
${\mathcal{J}}_{t}(h_{t})={\mathbb{E}}^{{\mathbb{P}}}\big[\|h_{t}(\mathbf{X}_{t}^{\leftarrow})-h_{t}^{*}(\mathbf{X}_{t}^{\leftarrow})\|_{2}^{2}\big]+V_{t}^{2},$
where $V_{t}^{2}\coloneq{\mathbb{E}}^{{\mathbb{P}}}[\mathrm{Var}(w(\mathbf{X}_{T}^{\leftarrow})|\mathbf{X}_{t}^{\leftarrow})]$ is a constant independent of $h_{t}$.
Proof of Proposition 4.1.
By a direct calculation, we have
(B.1) ${\mathcal{J}}_{t}(h_{t})$ $={\mathbb{E}}^{{\mathbb{P}}}\big[\|h_{t}(\mathbf{X}_{t}^{\leftarrow})-w(\mathbf{X}_{T}^{\leftarrow})\|_{2}^{2}\big]$
$={\mathbb{E}}^{{\mathbb{P}}}\big[\|h_{t}(\mathbf{X}_{t}^{\leftarrow})-h_{t}^{*}(\mathbf{X}_{t}^{\leftarrow})+h_{t}^{*}(\mathbf{X}_{t}^{\leftarrow})-w(\mathbf{X}_{T}^{\leftarrow})\|_{2}^{2}\big]$
$={\mathbb{E}}^{{\mathbb{P}}}\big[\|h_{t}(\mathbf{X}_{t}^{\leftarrow})-h_{t}^{*}(\mathbf{X}_{t}^{\leftarrow})\|_{2}^{2}]+{\mathbb{E}}^{{\mathbb{P}}}\big[\|h_{t}^{*}(\mathbf{X}_{t}^{\leftarrow})-w(\mathbf{X}_{T}^{\leftarrow})\|_{2}^{2}\big]$
$\quad+2{\mathbb{E}}^{{\mathbb{P}}}\big[\langle h_{t}(\mathbf{X}_{t}^{\leftarrow})-h_{t}^{*}(\mathbf{X}_{t}^{\leftarrow}),h_{t}^{*}(\mathbf{X}_{t}^{\leftarrow})-w(\mathbf{X}_{T}^{\leftarrow})\rangle\big],$
where $h_{t}^{*}$ is defined as (3.10). For the second term in (B.1), we have
${\mathbb{E}}^{{\mathbb{P}}}\big[\|h_{t}^{*}(\mathbf{X}_{t}^{\leftarrow})-w(\mathbf{X}_{T}^{\leftarrow})\|_{2}^{2}\big]$
$={\mathbb{E}}^{{\mathbb{P}}}\big[\|{\mathbb{E}}^{{\mathbb{P}}}[w(\mathbf{X}_{T}^{\leftarrow})|\mathbf{X}_{t}^{\leftarrow}]-w(\mathbf{X}_{T}^{\leftarrow})\|_{2}^{2}\big]$
$={\mathbb{E}}^{{\mathbb{P}}}{\mathbb{E}}^{{\mathbb{P}}}\big[\|{\mathbb{E}}^{{\mathbb{P}}}[w(\mathbf{X}_{T}^{\leftarrow})|\mathbf{X}_{t}^{\leftarrow}]-w(\mathbf{X}_{T}^{\leftarrow})\|_{2}^{2}|\mathbf{X}_{t}^{\leftarrow}\big]$
(B.2) $={\mathbb{E}}^{{\mathbb{P}}}[\mathrm{Var}(w(\mathbf{X}_{T}^{\leftarrow})|\mathbf{X}_{t}^{\leftarrow})],$
where the first equality holds from the definition of $h_{t}^{*}$ (3.10), and the second equality is due to the law of the total expectation. For the third term in (B.1), we find
${\mathbb{E}}^{{\mathbb{P}}}\big[\langle h_{t}(\mathbf{X}_{t}^{\leftarrow})-h_{t}^{*}(\mathbf{X}_{t}^{\leftarrow}),h_{t}^{*}(\mathbf{X}_{t}^{\leftarrow})-w(\mathbf{X}_{T}^{\leftarrow})\rangle\big]$
$={\mathbb{E}}^{{\mathbb{P}}}\big[\langle h_{t}(\mathbf{X}_{t}^{\leftarrow})-{\mathbb{E}}^{{\mathbb{P}}}[w(\mathbf{X}_{T}^{\leftarrow})|\mathbf{X}_{t}^{\leftarrow}],{\mathbb{E}}^{{\mathbb{P}}}[w(\mathbf{X}_{T}^{\leftarrow})|\mathbf{X}_{t}^{\leftarrow}]-w(\mathbf{X}_{T}^{\leftarrow})\rangle\big]$
$={\mathbb{E}}^{{\mathbb{P}}}\big[\langle h_{t}(\mathbf{X}_{t}^{\leftarrow})-{\mathbb{E}}^{{\mathbb{P}}}[w(\mathbf{X}_{T}^{\leftarrow})|\mathbf{X}_{t}^{\leftarrow}],{\mathbb{E}}^{{\mathbb{P}}}[w(\mathbf{X}_{T}^{\leftarrow})|\mathbf{X}_{t}^{\leftarrow}]-{\mathbb{E}}^{{\mathbb{P}}}[w(\mathbf{X}_{T}^{\leftarrow})|\mathbf{X}_{t}^{\leftarrow}]\rangle\big]$
(B.3) $=0,$
where the first equality holds from the definition of $h_{t}^{*}$ (3.10), and the second equality is due to the law of the total expectation. Substituting (B.2) and (B.3) into (B.1) completes the proof. ∎
Lemma B.1.
Suppose Assumptions 1 and 2 hold. Assume that $v_{t}\in H^{1}(p_{T-t})$. Then
$(\nabla h_{t}^{*},\nabla v_{t})_{L^{2}(p_{T-t})}=(\Delta h_{t}^{*}+\nabla h_{t}^{*}\cdot\nabla\log p_{T-t},v_{t})_{L^{2}(p_{T-t})}.$
Proof of Lemma B.1.
We first construct a sequence of cut-off functions $\{\psi_{k}\}_{k=1}^{\infty}\subseteq C_{0}^{\infty}({\mathbb{R}}^{d})$, satisfying
  1. (i)
    $\psi_{k}(\mathbf{x})=1$ for $\|\mathbf{x}\|_{2}\leq k$,
  2. (ii)
    $\psi_{k}(\mathbf{x})=0$ for $\|\mathbf{x}\|_{2}\geq 2k$,
  3. (iii)
    $\psi_{k}(\mathbf{x})\in(0,1)$ for $\mathbf{x}\in{\mathbb{R}}^{d}$, and
  4. (iv)
    $\|\nabla\psi_{k}(\mathbf{x})\|_{2}\leq Ck^{-1}$ for some constant $C$ independent of $\mathbf{x}$ and $k$.
See Brezis [2011, Theorem 8.7] for a detailed construction of such cut-off functions. Then we focus on the compactly supported approximations $\{\psi_{k}v_{t}\}_{k=1}^{\infty}$:
$-\langle\nabla h_{t}^{*},\nabla(\psi_{k}v_{t})\rangle_{L^{2}(p_{T-t})}$
$=-\int_{{\mathbb{R}}^{d}}\langle\nabla h_{t}^{*}(\mathbf{x}),\nabla(\psi_{k}v_{t})(\mathbf{x})\rangle p_{T-t}(\mathbf{x})\,\mathrm{d}\mathbf{x}$
(B.4) $=-\int_{{\mathbb{R}}^{d}}\nabla\cdot(\nabla h_{t}^{*}(\mathbf{x})p_{T-t}(\mathbf{x})(\psi_{k}v_{t})(\mathbf{x}))\,\mathrm{d}\mathbf{x}+\int_{{\mathbb{R}}^{d}}\nabla\cdot(\nabla h_{t}^{*}(\mathbf{x})p_{T-t}(\mathbf{x}))(\psi_{k}v_{t})(\mathbf{x})\,\mathrm{d}\mathbf{x}$
(B.5) $=\int_{{\mathbb{R}}^{d}}\nabla\cdot(\nabla h_{t}^{*}(\mathbf{x})p_{T-t}(\mathbf{x}))(\psi_{k}v_{t})(\mathbf{x})\,\mathrm{d}\mathbf{x},$
where we used the Gauss’s divergence theorem and the fact that $\psi_{k}v_{t}\in H_{0}^{1}(B({\mathbf{0}},k))$ with $B({\mathbf{0}},k)\coloneq\{\mathbf{x}:\|\mathbf{x}\|_{2}\leq k\}$. For the left-hand side of (B.5), we have
$\int_{{\mathbb{R}}^{d}}\langle\nabla h_{t}^{*}(\mathbf{x}),\nabla(\psi_{k}v_{t})(\mathbf{x})\rangle p_{T-t}(\mathbf{x})\,\mathrm{d}\mathbf{x}$
(B.6) $=\int_{{\mathbb{R}}^{d}}\psi_{k}(\mathbf{x})\langle\nabla h_{t}^{*}(\mathbf{x}),\nabla v_{t}(\mathbf{x})\rangle p_{T-t}(\mathbf{x})\,\mathrm{d}\mathbf{x}+\int_{{\mathbb{R}}^{d}}\langle\nabla h_{t}^{*}(\mathbf{x}),\nabla\psi_{k}(\mathbf{x})\rangle v_{t}(\mathbf{x})p_{T-t}(\mathbf{x})\,\mathrm{d}\mathbf{x}.$
For the second term in (B.6), we have
$\Big|\int_{{\mathbb{R}}^{d}}\langle\nabla h_{t}^{*}(\mathbf{x}),\nabla\psi_{k}(\mathbf{x})\rangle v_{t}(\mathbf{x})p_{T-t}(\mathbf{x})\,\mathrm{d}\mathbf{x}\Big|$
$\leq\int_{{\mathbb{R}}^{d}}\|\nabla h_{t}^{*}(\mathbf{x})\|_{2}\|\nabla\psi_{k}(\mathbf{x})\|_{2}|v_{t}(\mathbf{x})|p_{T-t}(\mathbf{x})\,\mathrm{d}\mathbf{x}$
$\leq\frac{C}{k}\int_{{\mathbb{R}}^{d}}\|\nabla h_{t}^{*}(\mathbf{x})\|_{2}|v_{t}(\mathbf{x})|p_{T-t}(\mathbf{x})\,\mathrm{d}\mathbf{x}$
$\leq\frac{C}{k}\|\nabla h_{t}^{*}\|_{L^{2}(p_{T-t})}\|v_{t}\|_{L^{2}(p_{T-t})},$
where the second inequality holds from the definition of the cut-off function, and the last inequality is due to Cauchy-Schwarz inequality and the fact that $h_{t}^{*}\in H^{1}(p_{T-t})$, which is a direct conclusion of Lemmas C.1 and C.4. Taking limitation with respect to $k\to\infty$ and using Lebesgue’s dominated convergence theorem yields
(B.7) $\lim_{k\to\infty}\int_{{\mathbb{R}}^{d}}\langle\nabla h_{t}^{*}(\mathbf{x}),\nabla\psi_{k}(\mathbf{x})\rangle v_{t}(\mathbf{x})p_{T-t}(\mathbf{x})\,\mathrm{d}\mathbf{x}=0.$
Combining (B.5), (B.6), and (B.7) and taking limitation with respect to $k\to\infty$ completes the proof. ∎
Proposition 4.2.
Let $\lambda>0$, $h_{t}^{*}$ be the Doob’s $h$-function defined as (3.10), and $h_{t}^{\lambda}$ be the minimizer of ${\mathcal{J}}_{t}^{\lambda}$ defined as (4.3). Then $h_{t}^{*}\in H^{2}(p_{T-t})$, and
$\|h_{t}^{\lambda}-h_{t}^{*}\|_{L^{2}(p_{T-t})}^{2}$ $\leq\lambda^{2}\|\Delta h_{t}^{*}+\nabla h_{t}^{*}\cdot\nabla\log p_{T-t}\|_{L^{2}(p_{T-t})}^{2},$
$\|\nabla h_{t}^{\lambda}-\nabla h_{t}^{*}\|_{L^{2}(p_{T-t})}^{2}$ $\leq\lambda\|\Delta h_{t}^{*}+\nabla h_{t}^{*}\cdot\nabla\log p_{T-t}\|_{L^{2}(p_{T-t})}^{2}.$
Proof of Proposition 4.2.
First, $h_{t}^{*}\in H^{2}(p_{T-t})$ is a direct conclusion of Lemmas C.1, C.4, and C.5. It remains to prove two inequalities. Using Proposition 4.1 and (4.3), we have
${\mathcal{J}}_{t}^{\lambda}(h_{t})=\|h_{t}-h_{t}^{*}\|_{L^{2}(p_{T-t})}^{2}+V_{t}^{2}+\lambda\|\nabla h_{t}\|_{L^{2}(p_{T-t})}^{2}$
Since $h_{t}^{\lambda}$ is the minimizer of ${\mathcal{J}}_{t}^{\lambda}$, the methods of variation imply that for any $v_{t}\in H^{1}(p_{T-t})$,
$\delta{\mathcal{J}}_{t}^{\lambda}(h_{t}^{\lambda},v_{t})=\langle h_{t}^{\lambda}-h_{t}^{*},v_{t}\rangle_{L^{2}(p_{T-t})}+\lambda\langle\nabla h_{t}^{\lambda},\nabla v_{t}\rangle_{L^{2}(p_{T-t})}=0,$
which implies
$\langle h_{t}^{\lambda}-h_{t}^{*},v_{t}\rangle_{L^{2}(p_{T-t})}+\lambda\langle\nabla h_{t}^{\lambda}-\nabla h_{t}^{*},\nabla v_{t}\rangle_{L^{2}(p_{T-t})}$
$=-\lambda\langle\nabla h_{t}^{*},\nabla v_{t}\rangle_{L^{2}(p_{T-t})}$
$=\lambda\langle\Delta h_{t}^{*}+\nabla h_{t}^{*}\cdot\nabla\log p_{T-t},v_{t}\rangle_{L^{2}(p_{T-t})},$
where the last equality invokes Lemma B.1. Substituting $v_{t}\coloneq h_{t}^{\lambda}-h_{t}^{*}$ yields
$\|h_{t}^{\lambda}-h_{t}^{*}\|_{L^{2}(p_{T-t})}^{2}+\lambda\|\nabla h_{t}^{\lambda}-\nabla h_{t}^{*}\|_{L^{2}(p_{T-t})}^{2}$
$=\lambda\langle\Delta h_{t}^{*}+\nabla h_{t}^{*}\cdot\nabla\log p_{T-t},h_{t}^{\lambda}-h_{t}^{*}\rangle_{L^{2}(p_{T-t})}$
(B.8) $\leq\lambda\|\Delta h_{t}^{*}+\nabla h_{t}^{*}\cdot\nabla\log p_{T-t}\|_{L^{2}(p_{T-t})}\|h_{t}^{\lambda}-h_{t}^{*}\|_{L^{2}(p_{T-t})},$
where the last inequality is due to Cauchy-Schwarz inequality. A direct conclusion is
$\|h_{t}^{\lambda}-h_{t}^{*}\|_{L^{2}(p_{T-t})}^{2}\leq\lambda^{2}\|\Delta h_{t}^{*}+\nabla h_{t}^{*}\cdot\nabla\log p_{T-t}\|_{L^{2}(p_{T-t})}^{2}.$
Then plugging this equality into (B.8) yields
$\|\nabla h_{t}^{\lambda}-\nabla h_{t}^{*}\|_{L^{2}(p_{T-t})}^{2}\leq\lambda\|\Delta h_{t}^{*}+\nabla h_{t}^{*}\cdot\nabla\log p_{T-t}\|_{L^{2}(p_{T-t})}^{2},$
which completes the proof. ∎
Proposition 4.3.
Let $\lambda>0$, and $h_{t}^{\lambda}$ be the minimizer of ${\mathcal{J}}_{t}^{\lambda}$ defined as (4.3). Then for any $h_{t}\in H^{1}(p_{T-t})$, we have
$\frac{1}{\max\{\lambda,1\}}\big\{{\mathcal{J}}_{t}^{\lambda}(h_{t})-{\mathcal{J}}_{t}^{\lambda}(h_{t}^{\lambda})\big\}\leq\|h_{t}-h_{t}^{\lambda}\|_{H^{1}(p_{T-t})}^{2}\leq\frac{1}{\min\{\lambda,1\}}\big\{{\mathcal{J}}_{t}^{\lambda}(h_{t})-{\mathcal{J}}_{t}^{\lambda}(h_{t}^{\lambda})\big\}.$
Proof of Proposition 4.3.
Using Proposition 4.1 and (4.3), we have
${\mathcal{J}}_{t}^{\lambda}(h_{t})=\|h_{t}-h_{t}^{*}\|_{L^{2}(p_{T-t})}^{2}+V_{t}^{2}+\lambda\|\nabla h_{t}\|_{L^{2}(p_{T-t})}^{2}.$
Since $h_{t}^{\lambda}$ is the minimizer of ${\mathcal{J}}_{t}^{\lambda}$, the methods of variation imply
(B.9) $\delta{\mathcal{J}}_{t}^{\lambda}(h_{t}^{\lambda},v_{t})=\langle h_{t}^{\lambda}-h_{t}^{*},v_{t}\rangle_{L^{2}(p_{T-t})}+\lambda\langle\nabla h_{t}^{\lambda},\nabla v_{t}\rangle_{L^{2}(p_{T-t})}=0,$
for any $v_{t}\in H^{1}(p_{T-t})$. A direct calculation yields
${\mathcal{J}}_{t}^{\lambda}(h_{t})={\mathcal{J}}_{t}^{\lambda}(h_{t}-h_{t}^{\lambda}+h_{t}^{\lambda})$
$=\|h_{t}-h_{t}^{\lambda}+h_{t}^{\lambda}-h_{t}^{*}\|_{L^{2}(p_{T-t})}^{2}+V_{t}^{2}+\lambda\|\nabla h_{t}-\nabla h_{t}^{\lambda}+\nabla h_{t}^{\lambda}\|_{L^{2}(p_{T-t})}^{2}$
$={\mathcal{J}}_{t}^{\lambda}(h_{t}^{\lambda})+\|h_{t}-h_{t}^{\lambda}\|_{L^{2}(p_{T-t})}^{2}+\lambda\|\nabla h_{t}-\nabla h_{t}^{\lambda}\|_{L^{2}(p_{T-t})}^{2}$
$\quad+2\langle h_{t}-h_{t}^{\lambda},h_{t}^{\lambda}-h_{t}^{*}\rangle_{L^{2}(p_{T-t})}+2\lambda\langle\nabla h_{t}-\nabla h_{t}^{\lambda},\nabla h_{t}^{\lambda}\rangle_{L^{2}(p_{T-t})}$
$={\mathcal{J}}_{t}^{\lambda}(h_{t}^{\lambda})+\|h_{t}-h_{t}^{\lambda}\|_{L^{2}(p_{T-t})}^{2}+\lambda\|\nabla h_{t}-\nabla h_{t}^{\lambda}\|_{L^{2}(p_{T-t})}^{2},$
where the last equality holds from (B.9). This completes the proof. ∎
Appendix C Derivations in Section 5.1
Lemma C.1.
Suppose Assumption 2 holds. Then for all $t\in(0,T)$,
$\underline{B}\leq h_{t}^{*}(\mathbf{x})\leq\bar{B},\quad\mathbf{x}\in{\mathbb{R}}^{d}.$
Proof of Lemma C.1.
A direct conclusion of Assumption 2 and the definition of Doob’s $h$-function $h_{t}^{*}$ (3.10). ∎
Lemma C.2 (Tweedie’s formula).
Let $t\in(0,T)$, and let $\mathbf{X}_{t}$ be defined as (2.1). Then
$\nabla\log p_{t}(\mathbf{x})+\frac{\mathbf{x}}{\sigma_{t}^{2}}=\frac{\mu_{t}}{\sigma_{t}^{2}}{\mathbb{E}}[\mathbf{X}_{0}|\mathbf{X}_{t}=\mathbf{x}],\quad\mathbf{x}\in{\mathbb{R}}^{d}.$
Proof of Lemma C.2.
It is straightforward that
$\nabla\log p_{t}(\mathbf{x})$ $=\frac{\nabla p_{t}(\mathbf{x})}{p_{t}(\mathbf{x})}$
$=\frac{1}{p_{t}(\mathbf{x})}\int\nabla_{\mathbf{x}}\varphi_{d}(\mathbf{x};\mu_{t}\mathbf{x}_{0},\sigma_{t}^{2}\mathbf{I}_{d})p_{0}(\mathbf{x}_{0})\,\mathrm{d}\mathbf{x}_{0}$
$=-\frac{1}{p_{t}(\mathbf{x})}\int\Big(\frac{\mathbf{x}-\mu_{t}\mathbf{x}_{0}}{\sigma_{t}^{2}}\Big)\varphi_{d}(\mathbf{x};\mu_{t}\mathbf{x}_{0},\sigma_{t}^{2}\mathbf{I}_{d})p_{0}(\mathbf{x}_{0})\,\mathrm{d}\mathbf{x}_{0}$
$=-\frac{1}{p_{t}(\mathbf{x})}\frac{\mathbf{x}}{\sigma_{t}^{2}}\int\varphi_{d}(\mathbf{x};\mu_{t}\mathbf{x}_{0},\sigma_{t}^{2}\mathbf{I}_{d})p_{0}(\mathbf{x}_{0})\,\mathrm{d}\mathbf{x}_{0}$
$\quad+\frac{\mu_{t}}{\sigma_{t}^{2}}\int\mathbf{x}_{0}\varphi_{d}(\mathbf{x};\mu_{t}\mathbf{x}_{0},\sigma_{t}^{2}\mathbf{I}_{d})\frac{p_{0}(\mathbf{x}_{0})}{p_{t}(\mathbf{x})}\,\mathrm{d}\mathbf{x}_{0}$
$=-\frac{\mathbf{x}}{\sigma_{t}^{2}}+\frac{\mu_{t}}{\sigma_{t}^{2}}{\mathbb{E}}[\mathbf{X}_{0}|\mathbf{X}_{t}=\mathbf{x}],$
where the second equality is due to (2.2), and last equality invokes the Bayes’ rule. This completes the proof. ∎
Lemma C.3.
Let $g:{\mathbb{R}}^{d}\rightarrow{\mathbb{R}}^{d}$ be an integrable function. Let $t\in(0,T)$, and let $\mathbf{X}_{t}$ be defined as (2.1). Then for each $\mathbf{x}\in{\mathbb{R}}^{d}$,
$\nabla_{\mathbf{x}}{\mathbb{E}}[g(\mathbf{X}_{0})|\mathbf{X}_{t}=\mathbf{x}]=\frac{\mu_{t}}{\sigma_{t}^{2}}\mathrm{Cov}(\mathbf{X}_{0},g(\mathbf{X}_{0})|\mathbf{X}_{t}=\mathbf{x}),$
where the $k$-th entry of $\mathrm{Cov}(\mathbf{X}_{0},g(\mathbf{X}_{0})|\mathbf{X}_{t}=\mathbf{x})$ is defined as $\mathrm{Cov}(X_{0,k},g(\mathbf{X}_{0})|\mathbf{X}_{t}=\mathbf{x})$ with $\mathbf{X}_{0}=(X_{0,1},\ldots,X_{0,d})$.
Proof of Lemma C.3.
According to Bayes’ rule, we have
${\mathbb{E}}[g(\mathbf{X}_{0})|\mathbf{X}_{t}=\mathbf{x}]=\frac{1}{p_{t}(\mathbf{x})}\int g(\mathbf{x}_{0})\varphi_{d}(\mathbf{x};\mu_{t}\mathbf{x}_{0},\sigma_{t}^{2}\mathbf{I}_{d})p_{0}(\mathbf{x}_{0})\,\mathrm{d}\mathbf{x}_{0}.$
Taking gradient with respect to $\mathbf{x}$ on both sides of the equality yields
$\nabla_{\mathbf{x}}{\mathbb{E}}[g(\mathbf{X}_{0})|\mathbf{X}_{t}=\mathbf{x}]$ $=\frac{1}{p_{t}(\mathbf{x})}\int g(\mathbf{x}_{0})\nabla_{\mathbf{x}}\varphi_{d}(\mathbf{x};\mu_{t}\mathbf{x}_{0},\sigma_{t}^{2}\mathbf{I}_{d})p_{0}(\mathbf{x}_{0})\,\mathrm{d}\mathbf{x}_{0}$
$\quad-\frac{\nabla p_{t}(\mathbf{x})}{p_{t}^{2}(\mathbf{x})}\int g(\mathbf{x}_{0})\varphi_{d}(\mathbf{x};\mu_{t}\mathbf{x}_{0},\sigma_{t}^{2}\mathbf{I}_{d})p_{0}(\mathbf{x}_{0})\,\mathrm{d}\mathbf{x}_{0}$
$=-\frac{1}{p_{t}(\mathbf{x})}\int g(\mathbf{x}_{0})\Big(\frac{\mathbf{x}-\mu_{t}\mathbf{x}_{0}}{\sigma_{t}^{2}}\Big)\varphi_{d}(\mathbf{x};\mu_{t}\mathbf{x}_{0},\sigma_{t}^{2}\mathbf{I}_{d})p_{0}(\mathbf{x}_{0})\,\mathrm{d}\mathbf{x}_{0}$
$\quad-\frac{\nabla\log p_{t}(\mathbf{x})}{p_{t}(\mathbf{x})}\int g(\mathbf{x}_{0})\varphi_{d}(\mathbf{x};\mu_{t}\mathbf{x}_{0},\sigma_{t}^{2}\mathbf{I}_{d})p_{0}(\mathbf{x}_{0})\,\mathrm{d}\mathbf{x}_{0}$
$=-\Big(\nabla\log p_{t}(\mathbf{x})+\frac{\mathbf{x}}{\sigma_{t}^{2}}\Big)\int g(\mathbf{x}_{0})\varphi_{d}(\mathbf{x};\mu_{t}\mathbf{x}_{0},\sigma_{t}^{2}\mathbf{I}_{d})\frac{p_{0}(\mathbf{x}_{0})}{p_{t}(\mathbf{x})}\,\mathrm{d}\mathbf{x}_{0}$
$\quad+\frac{\mu_{t}}{\sigma_{t}^{2}}\int\mathbf{x}_{0}g(\mathbf{x}_{0})\varphi_{d}(\mathbf{x};\mu_{t}\mathbf{x}_{0},\sigma_{t}^{2}\mathbf{I}_{d})\frac{p_{0}(\mathbf{x}_{0})}{p_{t}(\mathbf{x})}\,\mathrm{d}\mathbf{x}_{0}$
$=\frac{\mu_{t}}{\sigma_{t}^{2}}(-{\mathbb{E}}[\mathbf{X}_{0}|\mathbf{X}_{t}=\mathbf{x}]{\mathbb{E}}[g(\mathbf{X}_{0})|\mathbf{X}_{t}=\mathbf{x}]+{\mathbb{E}}[\mathbf{X}_{0}g(\mathbf{X}_{0})|\mathbf{X}_{t}=\mathbf{x}])$
$=\frac{\mu_{t}}{\sigma_{t}^{2}}\mathrm{Cov}(\mathbf{X}_{0},g(\mathbf{X}_{0})|\mathbf{X}_{t}=\mathbf{x}),$
where the fourth equality follows from Lemma C.2 and the Bayes’ rule. Here the $k$-th entry of the conditional covariance $\mathrm{Cov}(\mathbf{X}_{0},g(\mathbf{X}_{0})|\mathbf{X}_{t}=\mathbf{x})$ is defined as $\mathrm{Cov}(X_{0,k},g(\mathbf{X}_{0})|\mathbf{X}_{t}=\mathbf{x})$, where $\mathbf{X}_{0}=(X_{0,1},\ldots,X_{0,d})$. This completes the proof. ∎
Lemma C.4.
Suppose Assumptions 1 and 2 hold. Then for each $\mathbf{x}\in{\mathbb{R}}^{d}$ and $t>0$,
$\max_{1\leq k\leq d}|D_{k}h_{t}^{*}(\mathbf{x})|\leq\frac{2\bar{B}}{\sigma_{T-t}^{2}},$
where $D_{k}$ denote the differential operator with respect to the $k$-th entry of $\mathbf{x}$.
Proof of Lemma C.4.
According to the definition of Doob’s $h$-function $h_{t}^{*}$ (3.10) and the property of the time-reversal process (2.4), we have
$h^{*}(t,\mathbf{x})\coloneq{\mathbb{E}}^{{\mathbb{P}}}[w(\mathbf{X}_{T}^{\leftarrow})|\mathbf{X}_{t}^{\leftarrow}=\mathbf{x}]={\mathbb{E}}[w(\mathbf{X}_{0})|\mathbf{X}_{T-t}=\mathbf{x}],$
where the second expectation is with respect to the path measure of the forward process (2.1). Then it follows from Lemma C.3 that
(C.1) $\nabla h_{t}^{*}(\mathbf{x})=\frac{\mu_{T-t}}{\sigma_{T-t}^{2}}\mathrm{Cov}(\mathbf{X}_{0},w(\mathbf{X}_{0})|\mathbf{X}_{T-t}=\mathbf{x}).$
Then it follows from Assumptions 1 and 2 that for each $\mathbf{x}\in{\mathbb{R}}^{d}$,
(C.2) $\|\mathrm{Cov}(\mathbf{X}_{0},w(\mathbf{X}_{0})|\mathbf{X}_{T-t}=\mathbf{x})\|_{\infty}=\max_{1\leq k\leq d}\mathrm{Cov}(X_{0,k},w(\mathbf{X}_{0})|\mathbf{X}_{T-t}=\mathbf{x})\leq 2\bar{B}.$
Substituting (C.2) into (C.1) yields
$\|\nabla h_{t}^{*}(\mathbf{x})\|_{\infty}=\frac{\mu_{t}}{\sigma_{t}^{2}}\|\mathrm{Cov}(\mathbf{X}_{0},w(\mathbf{X}_{0})|\mathbf{X}_{T-t}=\mathbf{x})\|_{\infty}\leq\frac{2\bar{B}}{\sigma_{T-t}^{2}},$
where we used the fact that $\mu_{t}=\exp(-t)<1$. This completes the proof. ∎
Lemma C.5.
Suppose Assumptions 1 and 2 hold. Then for each $\mathbf{x}\in{\mathbb{R}}^{d}$ and $t>0$,
$|D_{k\ell}^{2}h_{t}^{*}(\mathbf{x})|\leq\frac{6\bar{B}}{\sigma_{T-t}^{4}},$
where $D_{k\ell}^{2}$ denote the second-order differential operator with respect to $k$-th and $\ell$-th entry.
Proof of Lemma C.5.
Taking derivative with respect to the $\ell$-th entry of $\mathbf{x}$ on both sides of (C.1) implies
(C.3) $D_{k\ell}^{2}h_{t}^{*}(\mathbf{x})=\frac{\mu_{T-t}}{\sigma_{T-t}^{2}}D_{\ell}\mathrm{Cov}(X_{0,k},w(\mathbf{X}_{0})|\mathbf{X}_{T-t}=\mathbf{x}).$
It remains to estimate the derivative of the conditional covariance. Indeed,
$D_{\ell}\mathrm{Cov}(X_{0,k},w(\mathbf{X}_{0})|\mathbf{X}_{T-t}=\mathbf{x})$
$=D_{\ell}{\mathbb{E}}[X_{0,k}|\mathbf{X}_{T-t}=\mathbf{x}]{\mathbb{E}}[w(\mathbf{X}_{0})|\mathbf{X}_{T-t}=\mathbf{x}]$
$\quad+{\mathbb{E}}[X_{0,k}|\mathbf{X}_{T-t}=\mathbf{x}]D_{\ell}{\mathbb{E}}[w(\mathbf{X}_{0})|\mathbf{X}_{T-t}=\mathbf{x}]$
$\quad-D_{\ell}{\mathbb{E}}[X_{0,k}w(\mathbf{X}_{0})|\mathbf{X}_{T-t}=\mathbf{x}]$
$=\frac{\mu_{T-t}}{\sigma_{T-t}^{2}}\mathrm{Cov}(X_{0,\ell},X_{0,k}|\mathbf{X}_{t}=\mathbf{x}){\mathbb{E}}[w(\mathbf{X}_{0})|\mathbf{X}_{T-t}=\mathbf{x}]$
$\quad+{\mathbb{E}}[X_{0,k}|\mathbf{X}_{T-t}=\mathbf{x}]\frac{\mu_{T-t}}{\sigma_{T-t}^{2}}\mathrm{Cov}(w(\mathbf{X}_{0}),X_{0,\ell}|\mathbf{X}_{T-t}=\mathbf{x})$
$\quad-\frac{\mu_{T-t}}{\sigma_{T-t}^{2}}\mathrm{Cov}(X_{0,k}w(\mathbf{X}_{0}),X_{0,\ell}|\mathbf{X}_{T-t}=\mathbf{x}),$
where the last equality holds from Lemma C.3. Concequently, for each $\mathbf{x}\in{\mathbb{R}}^{d}$ and $t>0$,
(C.4) $|D_{\ell}\mathrm{Cov}(X_{0,k},w(\mathbf{X}_{0})|\mathbf{X}_{T-t}=\mathbf{x})|\leq\frac{6\mu_{T-t}\bar{B}}{\sigma_{T-t}^{2}},$
where we used Assumptions 1 and 2. Substituting (C.4) into (C.3) completes the proof. ∎
Proposition 5.1.
Suppose Assumptions 1 and 2 hold. Then for all $t\in(0,T)$ and $\mathbf{x}\in{\mathbb{R}}^{d}$, the following bounds hold:
  1. (i)
    $\underline{B}\leq h_{t}^{*}(\mathbf{x})\leq\bar{B}$;
  2. (ii)
    $\max_{1\leq k\leq d}|D_{k}h_{t}^{*}(\mathbf{x})|\leq 2\sigma_{T-t}^{-2}\bar{B}$; and
  3. (iii)
    $\max_{1\leq k,\ell\leq d}|D_{k\ell}^{2}h_{t}^{*}(\mathbf{x})|\leq 6\sigma_{T-t}^{-4}\bar{B}$,
where $D_{k}$ and $D_{k\ell}^{2}$ denote the first-order and second-order partial derivatives with respect to the input coordinates, respectively.
Proof of Proposition 5.1.
A direct conclusion of Lemmas C.1, C.4, and C.5. ∎
Appendix D Derivations in Section 5.2
D.1  Oracle inequality of variationally stable Doob’s matching
Lemma 5.2.
Suppose Assumptions 1 and 2 hold. Let $t\in(0,T)$ and let ${\mathscr{H}}_{t}$ be a hypothesis class. Let $\widehat{h}_{t}^{\lambda}$ be the gradient-regularized empirical risk minimizer defined as (4.6), and let $h_{t}^{*}$ be the Doob’s $h$-function defined as (3.10). Then the following inequalities hold:
${\mathbb{E}}\Big[\|\widehat{h}_{t}^{\lambda}-h_{t}^{*}\|_{L^{2}(p_{T-t})}^{2}\Big]$ $\lesssim\inf_{h_{t}\in{\mathscr{H}}_{t}}\Big\{\|h_{t}-h_{t}^{*}\|_{L^{2}(p_{T-t})}^{2}+\lambda\|\nabla h_{t}-\nabla h_{t}^{*}\|_{L^{2}(p_{T-t})}^{2}\Big\}$
$\quad+\bar{B}^{2}\Big(\frac{\mathrm{VCdim}({\mathscr{H}}_{t})}{n\log^{-1}n}\Big)^{\frac{1}{2}}+\frac{\lambda d\bar{B}^{2}}{\sigma_{T-t}^{4}}\Big(\frac{\mathrm{VCdim}(\nabla{\mathscr{H}}_{t})}{n\log^{-1}n}\Big)^{\frac{1}{2}}+\frac{\lambda^{2}d\bar{B}^{2}}{\sigma_{T-t}^{8}},$
${\mathbb{E}}\Big[\|\nabla\widehat{h}_{t}^{\lambda}-\nabla h_{t}^{*}\|_{L^{2}(p_{T-t})}^{2}\Big]$ $\lesssim\inf_{h_{t}\in{\mathscr{H}}_{t}}\Big\{\frac{1}{\lambda}\|h_{t}-h_{t}^{*}\|_{L^{2}(p_{T-t})}^{2}+\|\nabla h_{t}-\nabla h_{t}^{*}\|_{L^{2}(p_{T-t})}^{2}\Big\}$
$\quad+\frac{\bar{B}^{2}}{\lambda}\Big(\frac{\mathrm{VCdim}({\mathscr{H}}_{t})}{n\log^{-1}n}\Big)^{\frac{1}{2}}+\frac{d\bar{B}^{2}}{\sigma_{T-t}^{4}}\Big(\frac{\mathrm{VCdim}(\nabla{\mathscr{H}}_{t})}{n\log^{-1}n}\Big)^{\frac{1}{2}}+\frac{\lambda d\bar{B}^{2}}{\sigma_{T-t}^{8}},$
where the notation $\lesssim$ hides absolute constants.
Proof of Lemma 5.2.
It follows from Proposition 4.3 and Lemma D.1 that
${\mathbb{E}}\Big[\|\widehat{h}_{t}^{\lambda}-h_{t}^{\lambda}\|_{L^{2}(p_{T-t})}^{2}\Big]+\lambda{\mathbb{E}}\Big[\|\nabla\widehat{h}_{t}^{\lambda}-\nabla h_{t}^{\lambda}\|_{L^{2}(p_{T-t})}^{2}\Big]$
$={\mathbb{E}}\Big[{\mathcal{J}}_{t}^{\lambda}(\widehat{h}_{t}^{\lambda})-{\mathcal{J}}_{t}^{\lambda}(h_{t}^{\lambda})\Big]$
$\leq\inf_{h_{t}\in{\mathscr{H}}_{t}}\Big\{\|h_{t}-h_{t}^{\lambda}\|_{L^{2}(p_{T-t})}^{2}+\lambda\|\nabla h_{t}-\nabla h_{t}^{\lambda}\|_{L^{2}(p_{T-t})}^{2}\Big\}$
$\quad+80\bar{B}^{2}\Big(\frac{\mathrm{VCdim}({\mathscr{H}}_{t})}{n\log^{-1}n}\Big)^{\frac{1}{2}}+8\frac{\lambda d\bar{B}^{2}}{\sigma_{T-t}^{4}}\Big(\frac{\mathrm{VCdim}(\nabla{\mathscr{H}}_{t})}{n\log^{-1}n}\Big)^{\frac{1}{2}}$
$\leq\inf_{h_{t}\in{\mathscr{H}}_{t}}\Big\{2\|h_{t}-h_{t}^{*}\|_{L^{2}(p_{T-t})}^{2}+2\lambda\|\nabla h_{t}-\nabla h_{t}^{*}\|_{L^{2}(p_{T-t})}^{2}\Big\}$
$\quad+80\bar{B}^{2}\Big(\frac{\mathrm{VCdim}({\mathscr{H}}_{t})}{n\log^{-1}n}\Big)^{\frac{1}{2}}+8\frac{\lambda d\bar{B}^{2}}{\sigma_{T-t}^{4}}\Big(\frac{\mathrm{VCdim}(\nabla{\mathscr{H}}_{t})}{n\log^{-1}n}\Big)^{\frac{1}{2}}$
$\quad+2\|h_{t}^{*}-h_{t}^{\lambda}\|_{L^{2}(p_{T-t})}^{2}+2\lambda\|\nabla h_{t}^{*}-\nabla h_{t}^{\lambda}\|_{L^{2}(p_{T-t})}^{2},$
where the last inequality holds from the triangular inequality. Using the triangular inequality again, we have
(D.1) ${\mathbb{E}}\Big[\|\widehat{h}_{t}^{\lambda}-h_{t}^{*}\|_{L^{2}(p_{T-t})}^{2}\Big]+\lambda{\mathbb{E}}\Big[\|\nabla\widehat{h}_{t}^{\lambda}-\nabla h_{t}^{*}\|_{L^{2}(p_{T-t})}^{2}\Big]$
$\leq 2{\mathbb{E}}\Big[\|\widehat{h}_{t}^{\lambda}-h_{t}^{\lambda}\|_{L^{2}(p_{T-t})}^{2}\Big]+2\lambda{\mathbb{E}}\Big[\|\nabla\widehat{h}_{t}^{\lambda}-\nabla h_{t}^{\lambda}\|_{L^{2}(p_{T-t})}^{2}\Big]$
$\quad+2\|h_{t}^{*}-h_{t}^{\lambda}\|_{L^{2}(p_{T-t})}^{2}+2\lambda\|\nabla h_{t}^{*}-\nabla h_{t}^{\lambda}\|_{L^{2}(p_{T-t})}^{2}$
$\leq\underbrace{\inf_{h_{t}\in{\mathscr{H}}_{t}}\Big\{4\|h_{t}-h_{t}^{*}\|_{L^{2}(p_{T-t})}^{2}+4\lambda\|\nabla h_{t}-\nabla h_{t}^{*}\|_{L^{2}(p_{T-t})}^{2}\Big\}}_{\text{approximation error}}$
$\quad+\underbrace{160\bar{B}^{2}\Big(\frac{\mathrm{VCdim}({\mathscr{H}}_{t})}{n\log^{-1}n}\Big)^{\frac{1}{2}}+16\frac{\lambda d\bar{B}^{2}}{\sigma_{T-t}^{4}}\Big(\frac{\mathrm{VCdim}(\nabla{\mathscr{H}}_{t})}{n\log^{-1}n}\Big)^{\frac{1}{2}}}_{\text{generalization error}}$
$\quad+\underbrace{6\|h_{t}^{*}-h_{t}^{\lambda}\|_{L^{2}(p_{T-t})}^{2}+6\lambda\|\nabla h_{t}^{*}-\nabla h_{t}^{\lambda}\|_{L^{2}(p_{T-t})}^{2}}_{\text{regularization gap}}.$
Combining Proposition 4.2 and Lemma D.6 yields
[…]
Substituting (D.2) into (D.1) completes the proof. ∎
Lemma D.1.
Suppose Assumptions 1 and 2 hold. Let $t\in(0,T)$ and let ${\mathscr{H}}_{t}$ be a hypothesis class. Let $\widehat{h}_{t}^{\lambda}$ be the gradient-regularized empirical risk minimizer defined as (4.6). Then we have
${\mathbb{E}}\big[{\mathcal{J}}_{t}^{\lambda}(\widehat{h}_{t}^{\lambda})-{\mathcal{J}}_{t}^{\lambda}(h_{t}^{\lambda})\big]$ $\leq\inf_{h_{t}\in{\mathscr{H}}_{t}}\Big\{\|h_{t}-h_{t}^{\lambda}\|_{L^{2}(p_{T-t})}^{2}+\lambda\|\nabla h_{t}-\nabla h_{t}^{\lambda}\|_{L^{2}(p_{T-t})}^{2}\Big\}$
$\quad+80\bar{B}^{2}\Big(\frac{\mathrm{VCdim}({\mathscr{H}}_{t})}{n\log^{-1}n}\Big)^{\frac{1}{2}}+8\frac{\lambda d\bar{B}^{2}}{\sigma_{T-t}^{4}}\Big(\frac{\mathrm{VCdim}(\nabla{\mathscr{H}}_{t})}{n\log^{-1}n}\Big)^{\frac{1}{2}},$
where the notation $\lesssim$ hides absolute constants.
Proof of Lemma D.1.
For any $h_{t}\in{\mathscr{H}}_{t}$, we have
${\mathcal{J}}_{t}^{\lambda}(\widehat{h}_{t}^{\lambda})-{\mathcal{J}}_{t}^{\lambda}(h_{t}^{\lambda})$
$={\mathcal{J}}_{t}^{\lambda}(\widehat{h}_{t}^{\lambda})-\widehat{{\mathcal{J}}}_{t}^{\lambda}(\widehat{h}_{t}^{\lambda})+\widehat{{\mathcal{J}}}_{t}^{\lambda}(\widehat{h}_{t}^{\lambda})-\widehat{{\mathcal{J}}}_{t}^{\lambda}(h_{t})+\widehat{{\mathcal{J}}}_{t}^{\lambda}(h_{t})-{\mathcal{J}}_{t}^{\lambda}(h_{t})+{\mathcal{J}}_{t}^{\lambda}(h_{t})-{\mathcal{J}}_{t}^{\lambda}(h_{t}^{\lambda})$
$\leq{\mathcal{J}}_{t}^{\lambda}(\widehat{h}_{t}^{\lambda})-\widehat{{\mathcal{J}}}_{t}^{\lambda}(\widehat{h}_{t}^{\lambda})+\widehat{{\mathcal{J}}}_{t}^{\lambda}(h_{t})-{\mathcal{J}}_{t}^{\lambda}(h_{t})+{\mathcal{J}}_{t}^{\lambda}(h_{t})-{\mathcal{J}}_{t}^{\lambda}(h_{t}^{\lambda}),$
where the inequality holds from that fact that $\widehat{h}_{t}^{\lambda}$ is the minimizer of $\widehat{{\mathcal{J}}}_{t}^{\lambda}$ over the hypothesis class ${\mathscr{H}}_{t}$. Taking expectation on both sides of the inequality yields
${\mathbb{E}}\big[{\mathcal{J}}_{t}^{\lambda}(\widehat{h}_{t}^{\lambda})-{\mathcal{J}}_{t}^{\lambda}(h_{t}^{\lambda})\big]$ $={\mathbb{E}}\big[{\mathcal{J}}_{t}^{\lambda}(\widehat{h}_{t}^{\lambda})-\widehat{{\mathcal{J}}}_{t}^{\lambda}(\widehat{h}_{t}^{\lambda})\big]+{\mathcal{J}}_{t}^{\lambda}(h_{t})-{\mathcal{J}}_{t}^{\lambda}(h_{t}^{\lambda})$
$\leq{\mathbb{E}}\Bigg[\sup_{h_{t}\in{\mathscr{H}}_{t}}{\mathcal{J}}_{t}^{\lambda}(h_{t})-\widehat{{\mathcal{J}}}_{t}^{\lambda}(h_{t})\Bigg]+{\mathcal{J}}_{t}^{\lambda}(h_{t})-{\mathcal{J}}_{t}^{\lambda}(h_{t}^{\lambda}),$
where the equality holds from ${\mathbb{E}}[\widehat{{\mathcal{J}}}_{t}(h_{t})]={\mathcal{J}}_{t}(h_{t})$ for each fixed $h_{t}$, and the inequality is due to $h_{t}\in{\mathscr{H}}_{t}$. By taking infimum on both sides of the inequality with respect to $h_{t}\in{\mathscr{H}}_{t}$, we have
(D.3) ${\mathbb{E}}\big[{\mathcal{J}}_{t}^{\lambda}(\widehat{h}_{t}^{\lambda})-{\mathcal{J}}_{t}^{\lambda}(h_{t}^{\lambda})\big]\leq\underbrace{{\mathbb{E}}\Bigg[\sup_{h_{t}\in{\mathscr{H}}_{t}}{\mathcal{J}}_{t}^{\lambda}(h_{t})-\widehat{{\mathcal{J}}}_{t}^{\lambda}(h_{t})\Bigg]}_{\text{generalization error}}+\underbrace{\inf_{h_{t}\in{\mathscr{H}}_{t}}\Big\{{\mathcal{J}}_{t}^{\lambda}(h_{t})-{\mathcal{J}}_{t}^{\lambda}(h_{t}^{\lambda})\Big\}}_{\text{approximation error}}.$
The rest of the proof is divided into three steps.
Step 1. Generalization error in (D.3).
For the generalization error in (D.3), we have the following decomposition:
(D.4) ${\mathbb{E}}\Bigg[\sup_{h_{t}\in{\mathscr{H}}_{t}}{\mathcal{J}}_{t}^{\lambda}(h_{t})-\widehat{{\mathcal{J}}}_{t}^{\lambda}(h_{t})\Bigg]$
$=\underbrace{{\mathbb{E}}\Bigg[\sup_{h_{t}\in{\mathscr{H}}_{t}}{\mathbb{E}}\big[(h_{t}(\mathbf{X}_{T-t})-w(\mathbf{X}_{0}))^{2}\big]-\frac{1}{n}\sum_{i=1}^{n}(h_{t}(\mathbf{X}_{T-t}^{i})-w(\mathbf{X}_{0}^{i}))^{2}\Bigg]}_{\text{(G1)}}$
$\quad+\lambda\underbrace{{\mathbb{E}}\Bigg[\sup_{h_{t}\in{\mathscr{H}}_{t}}{\mathbb{E}}\big[\|\nabla h_{t}(\mathbf{X}_{T-t})\|_{2}^{2}\big]-\frac{1}{n}\sum_{i=1}^{n}\|\nabla h_{t}(\mathbf{X}_{T-t}^{i})\|_{2}^{2}\Bigg]}_{\text{(G2)}}.$
We start from the term (G1) in (D.4). First, recall Proposition 4.1:
(D.5) ${\mathbb{E}}\big[(h_{t}(\mathbf{X}_{T-t})-w(\mathbf{X}_{0}))^{2}\big]={\mathbb{E}}\big[(h_{t}(\mathbf{X}_{T-t})-h_{t}^{*}(\mathbf{X}_{T-t}))^{2}\big]+V_{t}.$
For the empirical counterpart, it is straightforward that
$-\frac{1}{n}\sum_{i=1}^{n}(h_{t}(\mathbf{X}_{T-t}^{i})-w(\mathbf{X}_{0}^{i}))^{2}$
$=-\frac{1}{n}\sum_{i=1}^{n}(h_{t}(\mathbf{X}_{T-t}^{i})-h_{t}^{*}(\mathbf{X}_{T-t}^{i})+{\mathbb{E}}[w(\mathbf{X}_{0}^{i})|\mathbf{X}_{T-t}^{i}]-w(\mathbf{X}_{0}^{i}))^{2}$
$=-\frac{1}{n}\sum_{i=1}^{n}(h_{t}(\mathbf{X}_{T-t}^{i})-h_{t}^{*}(\mathbf{X}_{T-t}^{i}))^{2}-\frac{1}{n}\sum_{i=1}^{n}({\mathbb{E}}[w(\mathbf{X}_{0}^{i})|\mathbf{X}_{T-t}^{i}]-w(\mathbf{X}_{0}^{i}))^{2}$
$\quad-\frac{2}{n}\sum_{i=1}^{n}(h_{t}(\mathbf{X}_{T-t}^{i})-h_{t}^{*}(\mathbf{X}_{T-t}^{i}))({\mathbb{E}}[w(\mathbf{X}_{0}^{i})|\mathbf{X}_{T-t}^{i}]-w(\mathbf{X}_{0}^{i})),$
where the first equality invokes the definition of the Doob’s $h$-function $h_{t}^{*}$ in (3.10). Taking expectation with respect to $\{(\mathbf{X}_{0}^{i},\mathbf{X}_{T-t}^{i})\}_{i=1}^{n}$ yields
${\mathbb{E}}\Bigg[\sup_{h_{t}\in{\mathscr{H}}_{t}}{\mathbb{E}}\big[(h_{t}(\mathbf{X}_{T-t})-w(\mathbf{X}_{0}))^{2}\big]-\frac{1}{n}\sum_{i=1}^{n}(h_{t}(\mathbf{X}_{T-t}^{i})-w(\mathbf{X}_{0}^{i}))^{2}\Bigg]$
$={\mathbb{E}}\Bigg[\sup_{h_{t}\in{\mathscr{H}}_{t}}{\mathbb{E}}\big[(h_{t}(\mathbf{X}_{T-t})-h_{t}^{*}(\mathbf{X}_{T-t}))^{2}\big]-\frac{1}{n}\sum_{i=1}^{n}(h_{t}(\mathbf{X}_{T-t}^{i})-h_{t}^{*}(\mathbf{X}_{T-t}^{i}))^{2}\Bigg]$
$\quad+2{\mathbb{E}}\Bigg[\frac{1}{n}\sum_{i=1}^{n}h_{t}(\mathbf{X}_{T-t}^{i})({\mathbb{E}}[w(\mathbf{X}_{0}^{i})|\mathbf{X}_{T-t}^{i}]-w(\mathbf{X}_{0}^{i}))\Bigg]$
(D.6) $\leq 64\bar{B}^{2}\Big(\frac{\mathrm{VCdim}({\mathscr{H}}_{t})}{n\log^{-1}n}\Big)^{\frac{1}{2}}+16\bar{B}^{2}\Big(\frac{\mathrm{VCdim}({\mathscr{H}}_{t})}{n\log^{-1}n}\Big)^{\frac{1}{2}}=80\bar{B}^{2}\Big(\frac{\mathrm{VCdim}({\mathscr{H}}_{t})}{n\log^{-1}n}\Big)^{\frac{1}{2}},$
where the equality invokes (D.5), and the inequality holds from Lemmas D.2 and D.5. For the term (G2) in (D.4), using Lemma D.3 implies
(D.7) ${\mathbb{E}}\Bigg[\sup_{h_{t}\in{\mathscr{H}}_{t}}{\mathbb{E}}\big[\|\nabla h_{t}(\mathbf{X}_{T-t})\|_{2}^{2}\big]-\frac{1}{n}\sum_{i=1}^{n}\|\nabla h_{t}(\mathbf{X}_{T-t}^{i})\|_{2}^{2}\Bigg]\leq\frac{8d\bar{B}^{2}}{\sigma_{T-t}^{4}}\Big(\frac{\mathrm{VCdim}(\nabla{\mathscr{H}}_{t})}{n\log^{-1}n}\Big)^{\frac{1}{2}}.$
Substituting (D.6) and (D.7) into (D.4) yields a generalization error bound:
(D.8) ${\mathbb{E}}\Bigg[\sup_{h_{t}\in{\mathscr{H}}_{t}}{\mathcal{J}}_{t}^{\lambda}(h_{t})-\widehat{{\mathcal{J}}}_{t}^{\lambda}(h_{t})\Bigg]\leq 80\bar{B}^{2}\Big(\frac{\mathrm{VCdim}({\mathscr{H}}_{t})}{n\log^{-1}n}\Big)^{\frac{1}{2}}+\frac{8d\bar{B}^{2}}{\sigma_{T-t}^{4}}\Big(\frac{\mathrm{VCdim}(\nabla{\mathscr{H}}_{t})}{n\log^{-1}n}\Big)^{\frac{1}{2}}.$
Step 2. Approximation error in (D.3). According to the proof of Proposition 4.3, we have
(D.9) ${\mathcal{J}}_{t}^{\lambda}(h_{t})-{\mathcal{J}}_{t}^{\lambda}(h_{t}^{\lambda})=\|h_{t}-h_{t}^{\lambda}\|_{L^{2}(p_{T-t})}^{2}+\lambda\|\nabla h_{t}-\nabla h_{t}^{\lambda}\|_{L^{2}(p_{T-t})}^{2}.$
Step 3. Conclusion. Substituting (D.8) and (D.9) into (D.3) completes the proof. ∎
D.2  Auxilary lemmas for the oracle inequality
According to the standard techniques of symmetrization [Mohri et al., 2018, Theorem 3.3], we have the following generalization bounds. We introduce the concept of Rademacher complexity [Bartlett and Mendelson, 2002, Mohri et al., 2018], which is crucial for analyzing the generalization error.
Definition 3 (Rademacher complexity).
Let $\mathscr{H}$} be a function class, and let $\mathbf{X}^{1:n}\coloneq(\mathbf{X}^{1},\ldots,\mathbf{X}^{n})$ be a set of samples. The empirical Rademacher complexity of $\mathscr{H}$} with respect to $\mathbf{X}_{1:n}$ is defined as
${\mathfrak{R}}({\mathscr{H}}|\mathbf{X}^{1:n})\coloneq{\mathbb{E}}\Big[\sup_{h\in{\mathscr{H}}}\frac{1}{n}\sum_{i=1}^{n}\varepsilon_{i}h(\mathbf{X}^{i})\Big|\mathbf{X}^{1:n}\Big],$
where $\varepsilon_{1},\ldots,\varepsilon_{n}$ are i.i.d. Rademacher random variables. The Rademacher complexity of $\mathscr{H}$} is the expectation of empirical Rademacher complexity with respect to the distribution of $\mathbf{X}^{1:n}$ defined as
${\mathfrak{R}}_{n}({\mathscr{H}})\coloneq{\mathbb{E}}[\widehat{{\mathfrak{R}}}({\mathscr{H}}|\mathbf{X}^{1:n})]={\mathbb{E}}\Big[\sup_{h\in{\mathscr{H}}}\frac{1}{n}\sum_{i=1}^{n}\varepsilon_{i}h(\mathbf{X}^{i})\Big].$
Lemma D.2.
Suppose Assumptions 1 and 2 hold. Then
${\mathbb{E}}\Bigg[\sup_{h_{t}\in{\mathscr{H}}_{t}}\|h_{t}-h_{t}^{*}\|_{L^{2}(p_{T-t})}^{2}-\frac{1}{n}\sum_{i=1}^{n}(h_{t}(\mathbf{X}_{T-t}^{i})-h_{t}^{*}(\mathbf{X}_{T-t}^{i}))^{2}\Bigg]\leq 64\bar{B}^{2}\Big(\frac{\mathrm{VCdim}({\mathscr{H}}_{t})}{n\log^{-1}n}\Big)^{\frac{1}{2}},$
where the expectation is taken with respect to $\mathbf{X}_{T-t}^{1},\ldots,\mathbf{X}_{T-t}^{n}\sim^{\mathrm{i.i.d.}}p_{T-t}$.
Proof of Lemma D.2.
Let $\mathbf{X}_{T-t}^{1,\prime},\ldots,\mathbf{X}_{T-t}^{n,\prime}$ be independent copies of $\mathbf{X}_{T-t}^{1},\ldots,\mathbf{X}_{T-t}^{n}$. Let $\varepsilon_{1},\ldots,\varepsilon_{n}$ be a set of i.i.d. Rademacher variables, which are independent of $\mathbf{X}_{T-t}^{1,\prime},\ldots,\mathbf{X}_{T-t}^{n,\prime}$ and $\mathbf{X}_{T-t}^{1},\ldots,\mathbf{X}_{T-t}^{n}$. It follows that
${\mathbb{E}}\Bigg[\sup_{h_{t}\in{\mathscr{H}}_{t}}\|h_{t}-h_{t}^{*}\|_{L^{2}(p_{T-t})}^{2}-\frac{1}{n}\sum_{i=1}^{n}(h_{t}(\mathbf{X}_{T-t}^{i})-h_{t}^{*}(\mathbf{X}_{T-t}^{i}))^{2}\Bigg]$
$\leq{\mathbb{E}}\Bigg[\sup_{h_{t}\in{\mathscr{H}}_{t}}{\mathbb{E}}\big[(h_{t}(\mathbf{X}_{T-t})-h_{t}^{*}(\mathbf{X}_{T-t}))^{2}\big]-\frac{1}{n}\sum_{i=1}^{n}(h_{t}(\mathbf{X}_{T-t}^{i})-h_{t}^{*}(\mathbf{X}_{T-t}^{i}))^{2}\Bigg]$
$={\mathbb{E}}\Bigg[\sup_{h_{t}\in{\mathscr{H}}_{t}}{\mathbb{E}}\Bigg[\frac{1}{n}\sum_{i=1}^{n}(h_{t}(\mathbf{X}_{T-t}^{i,\prime})-h_{t}^{*}(\mathbf{X}_{T-t}^{i,\prime}))^{2}\Bigg]-\frac{1}{n}\sum_{i=1}^{n}(h_{t}(\mathbf{X}_{T-t}^{i})-h_{t}^{*}(\mathbf{X}_{T-t}^{i}))^{2}\Bigg]$
$\leq{\mathbb{E}}\Bigg[\sup_{h_{t}\in{\mathscr{H}}_{t}}\frac{1}{n}\sum_{i=1}^{n}\Big\{(h_{t}(\mathbf{X}_{T-t}^{i,\prime})-h_{t}^{*}(\mathbf{X}_{T-t}^{i,\prime}))^{2}-(h_{t}(\mathbf{X}_{t}^{i})-h_{t}^{*}(\mathbf{X}_{t}^{i}))^{2}\Big\}\Bigg]$
$={\mathbb{E}}\Bigg[\sup_{h_{t}\in{\mathscr{H}}_{t}}\frac{1}{n}\sum_{i=1}^{n}\varepsilon_{i}\Big\{(h_{t}(\mathbf{X}_{T-t}^{i,\prime})-h_{t}^{*}(\mathbf{X}_{T-t}^{i,\prime}))^{2}-(h_{t}(\mathbf{X}_{T-t}^{i})-h_{t}^{*}(\mathbf{X}_{T-t}^{i}))^{2}\Big\}\Bigg]$
(D.10) $=2{\mathbb{E}}\Bigg[\sup_{h_{t}\in{\mathscr{H}}_{t}}\frac{1}{n}\sum_{i=1}^{n}\varepsilon_{i}(h_{t}(\mathbf{X}_{T-t}^{i})-h_{t}^{*}(\mathbf{X}_{T-t}^{i}))^{2}\Bigg]\leq 8\bar{B}{\mathfrak{R}}_{n}({\mathscr{H}}_{t}),$
where the second inequality holds from Jensen’s inequality, and last inequality is due to Ledoux-Talagrand contraction inequality [Mohri et al., 2018, Lemma 5.7] and Lemma C.1.
It remains to bound the Rademacher complexity ${\mathfrak{R}}_{n}({\mathscr{H}}_{t})$ in (D.10). Let $\delta>0$ and ${\mathscr{H}}_{t}^{\delta}$ be an $L^{\infty}(\mathbf{X}_{T-t}^{1:n})$ $\delta$-cover of ${\mathscr{H}}_{t}$ satisfying $|{\mathscr{H}}_{t}^{\delta}|=N(\delta,{\mathscr{H}}_{t},L^{\infty}(\mathbf{X}_{T-t}^{1:n}))$. Then for any $h_{t}\in{\mathscr{H}}_{t}$, there exists $h_{t}^{\delta}\in{\mathscr{H}}_{t}^{\delta}$ such that
$\frac{1}{n}\sum_{i=1}^{n}\varepsilon_{i}h_{t}(\mathbf{X}_{t}^{i})-\frac{1}{n}\sum_{i=1}^{n}\varepsilon_{i}h_{t}^{\delta}(\mathbf{X}_{t}^{i})\leq\delta.$
As a consequence,
${\mathfrak{R}}({\mathscr{H}}_{t}\mid\mathbf{X}_{T-t}^{1:n})$ $={\mathbb{E}}\Bigg[\sup_{h_{t}\in{\mathscr{H}}_{t}}\frac{1}{n}\sum_{i=1}^{n}\varepsilon_{i}h_{t}(\mathbf{X}_{t}^{i})\mid\mathbf{X}_{T-t}^{1:n}\Bigg]$
$\leq{\mathbb{E}}\Bigg[\sup_{h_{t}^{\delta}\in{\mathscr{H}}_{t}^{\delta}}\frac{1}{n}\sum_{i=1}^{n}\varepsilon_{i}h_{t}^{\delta}(\mathbf{X}_{t}^{i})\mid\mathbf{X}_{T-t}^{1:n}\Bigg]+\delta$
$\leq\bar{B}\Big(\frac{2\log|{\mathscr{H}}_{t}^{\delta}|}{n}\Big)^{\frac{1}{2}}+\delta$
$=\bar{B}\Big(\frac{2\log N(\delta,{\mathscr{H}}_{t},L^{\infty}(\mathbf{X}_{T-t}^{1:n}))}{n}\Big)^{\frac{1}{2}}+\delta,$
where $\varepsilon_{1},\ldots,\varepsilon_{n}$ are a sequence of i.i.d. Rademacher variables, the second inequality follows from Massart’s lemma [Mohri et al., 2018, Theorem 3.7], and the equality is due to the definition of ${\mathscr{H}}_{t}^{\delta}$. Then setting $\delta=\bar{B}/\sqrt{n}$ yields
(D.11) ${\mathfrak{R}}({\mathscr{H}}_{t}\mid\mathbf{X}_{T-t}^{1:n})\leq\bar{B}\Big(\frac{2\log N(B/\sqrt{n},{\mathscr{H}}_{t},L^{\infty}(\mathbf{X}_{T-t}^{1:n}))}{n}\Big)^{\frac{1}{2}}\leq 8\bar{B}\Big(\frac{\mathrm{VCdim}({\mathscr{H}}_{t})}{n\log^{-1}n}\Big)^{\frac{1}{2}},$
where the last inequality holds from Anthony et al. [1999, Theorem 12.2]. Substituting (D.11) into (D.10) completes the proof. ∎
By a similar argument as Lemma D.2, we have the following generalization bounds for the gradient term.
Lemma D.3.
Suppose Assumptions 1 and 2 hold. Then
${\mathbb{E}}\Bigg[\sup_{h_{t}\in{\mathscr{H}}_{t}}\|\nabla h_{t}\|_{L^{2}(p_{T-t})}^{2}-\frac{1}{n}\sum_{i=1}^{n}\|\nabla h_{t}(\mathbf{X}_{T-t}^{i})\|_{2}^{2}\Bigg]\leq\frac{8d\bar{B}^{2}}{\sigma_{T-t}^{4}}\Big(\frac{\mathrm{VCdim}(\nabla{\mathscr{H}}_{t})}{n\log^{-1}n}\Big)^{\frac{1}{2}},$
where the expectation is taken with respect to $\mathbf{X}_{T-t}^{1},\ldots,\mathbf{X}_{T-t}^{i}\sim^{\mathrm{i.i.d.}}p_{T-t}$.
Proof of Lemma D.3.
It is straightforward that
${\mathbb{E}}\Bigg[\sup_{h_{t}\in{\mathscr{H}}_{t}}\|\nabla h_{t}\|_{L^{2}(p_{T-t})}^{2}-\frac{1}{n}\sum_{i=1}^{n}\|\nabla h_{t}(\mathbf{X}_{T-t}^{i})\|_{2}^{2}\Bigg]$
$\leq\sum_{k=1}^{d}{\mathbb{E}}\Bigg[\sup_{h_{t}\in{\mathscr{H}}_{t}}\|D_{k}h_{t}\|_{L^{2}(p_{T-t})}^{2}-\frac{1}{n}\sum_{i=1}^{n}(D_{k}h_{t}(\mathbf{X}_{T-t}^{i}))^{2}\Bigg]$
$\leq\sum_{k=1}^{d}\frac{8\bar{B}^{2}}{\sigma_{T-t}^{4}}\Big(\frac{\mathrm{VCdim}(D_{k}{\mathscr{H}}_{t})}{n\log^{-1}n}\Big)^{\frac{1}{2}}\leq\frac{8d\bar{B}^{2}}{\sigma_{T-t}^{4}}\Big(\frac{\mathrm{VCdim}(\nabla{\mathscr{H}}_{t})}{n\log^{-1}n}\Big)^{\frac{1}{2}},$
where the first inequality holds from the convexity of supremum and Jensen’s inequality, the second inequality invokes a similar argument as Lemma D.2, and the last inequality holds from the definition of $\mathrm{VCdim}(\nabla{\mathscr{H}}_{t})$. This completes the proof. ∎
The following lemma is an extension of Bartlett and Mendelson [2002, Lemma 4].
Lemma D.4.
Let $\mathbf{z}=(z_{1},\ldots,z_{n})\in{\mathcal{Z}}\subseteq{\mathbb{R}}^{n}$. Let $\xi_{1},\ldots,\xi_{n}$ be a sequence of i.i.d. random variables with $|\xi_{i}|<K$ and ${\mathbb{E}}[\xi_{i}]=0$ for each $1\leq i\leq n$. Then it follows that
${\mathbb{E}}\Bigg[\sup_{\mathbf{z}\in{\mathcal{Z}}}\frac{1}{n}\sum_{i=1}^{n}\xi_{i}z_{i}\Bigg]\leq 2K{\mathbb{E}}\Bigg[\sup_{\mathbf{z}\in{\mathcal{Z}}}\frac{1}{n}\sum_{i=1}^{n}\varepsilon_{i}z_{i}\Bigg],$
where $\varepsilon_{1},\ldots,\varepsilon_{n}$ is a sequence of i.i.d. Rademacher variables.
Proof of Lemma D.4.
The proof relies on the symmetrization technique. Let $\xi_{1}^{\prime},\ldots,\xi_{n}^{\prime}$ be independent copies of $\xi_{1},\ldots,\xi_{n}$. It follows that
${\mathbb{E}}\Bigg[\sup_{\mathbf{z}\in{\mathcal{Z}}}\frac{1}{n}\sum_{i=1}^{n}\xi_{i}z_{i}\Bigg]$ $={\mathbb{E}}\Bigg[\sup_{\mathbf{z}\in{\mathcal{Z}}}{\mathbb{E}}\Bigg[\frac{1}{n}\sum_{i=1}^{n}(\xi_{i}-\xi_{i}^{\prime})z_{i}\Bigg|\xi_{1},\ldots,\xi_{n}\Bigg]\Bigg]$
$\leq{\mathbb{E}}\Bigg[\sup_{\mathbf{z}\in{\mathcal{Z}}}\frac{1}{n}\sum_{i=1}^{n}(\xi_{i}-\xi_{i}^{\prime})z_{i}\Bigg]={\mathbb{E}}\Bigg[\sup_{\mathbf{z}\in{\mathcal{Z}}}\frac{1}{n}\sum_{i=1}^{n}\varepsilon_{i}(\xi_{i}-\xi_{i}^{\prime})z_{i}\Bigg]$
$\leq 2{\mathbb{E}}\Bigg[\sup_{\mathbf{z}\in{\mathcal{Z}}}\frac{1}{n}\sum_{i=1}^{n}\varepsilon_{i}\xi_{i}z_{i}\Bigg]\leq 2K{\mathbb{E}}\Bigg[\sup_{\mathbf{z}\in{\mathcal{Z}}}\frac{1}{n}\sum_{i=1}^{n}\varepsilon_{i}z_{i}\Bigg],$
where the first equality due to ${\mathbb{E}}[\xi_{i}^{\prime}]=0$, the first inequality holds from Jensen’s inequality, and the second equality follows from the fact that distribution of $(\xi_{i}-\xi_{i}^{\prime})$ is symmetric around zero, so it has the same distribution as $\varepsilon_{i}(\xi_{i}-\xi_{i}^{\prime})$. The second inequality comes from the triangular inequality for the supremum, and we used the fact that $\xi_{i}$ and $\xi_{i}^{\prime}$ are identically distributed. The last inequality invokes Ledoux-Talagrand contraction inequality [Mohri et al., 2018, Lemma 5.7] and $\max_{1\leq i\leq n}|\xi_{i}|\leq K$. This completes the proof. ∎
Lemma D.5.
Suppose Assumptions 1 and 2 hold. Then
${\mathbb{E}}\Bigg[\sup_{h_{t}\in{\mathscr{H}}_{t}}\frac{1}{n}\sum_{i=1}^{n}h_{t}(\mathbf{X}_{T-t}^{i})(w(\mathbf{X}_{0}^{i})-{\mathbb{E}}[w(\mathbf{X}_{0}^{i})\mid\mathbf{X}_{T-t}^{i}])\mid\mathbf{X}_{T-t}^{1:n}\Bigg]\leq 16\bar{B}^{2}\Big(\frac{\mathrm{VCdim}({\mathscr{H}}_{t})}{n\log^{-1}n}\Big)^{\frac{1}{2}},$
where the expectation is taken with respect to $\mathbf{X}_{T-t}^{1},\ldots,\mathbf{X}_{T-t}^{i}\sim^{\mathrm{i.i.d.}}p_{T-t}$.
Proof of Lemma D.5.
Define a sequence of auxilary random variables
$\xi_{i}\coloneq w(\mathbf{X}_{0}^{i})-{\mathbb{E}}[w(\mathbf{X}_{0}^{i})\mid\mathbf{X}_{T-t}^{i}].$
It is apparent that ${\mathbb{E}}[\xi_{i}\mid\mathbf{X}_{T-t}^{i}]=0$, and $|\xi_{i}|\leq\bar{B}$. Using Lemma D.4 yields
${\mathbb{E}}\Bigg[\sup_{h_{t}\in{\mathscr{H}}_{t}}\frac{1}{n}\sum_{i=1}^{n}\xi_{i}h_{t}(\mathbf{X}_{T-t}^{i})\mid\mathbf{X}_{T-t}^{1:n}\Bigg]$ $\leq 2\bar{B}{\mathbb{E}}\Bigg[\sup_{h_{t}\in{\mathscr{H}}_{t}}\frac{1}{n}\sum_{i=1}^{n}\varepsilon_{i}h_{t}(\mathbf{X}_{T-t}^{i})\mid\mathbf{X}_{T-t}^{1:n}\Bigg]$
$=2\bar{B}{\mathfrak{R}}({\mathscr{H}}_{t}\mid\mathbf{X}_{T-t}^{1:n}),$
where $\varepsilon_{1},\ldots,\varepsilon_{n}$ is a sequence of i.i.d. Rademacher variables. Here the first inequality follows from Lemma D.4, and the second inequality is due to the fact that $\widehat{h}_{t}^{\lambda}\in{\mathscr{H}}_{t}$, the second inequality holds from Lemma D.4. Finally, using (D.11) completes the proof. ∎
Lemma D.6.
Suppose Assumptions 1 and 2 hold. Then
(D.12) $\|\Delta h_{t}^{*}+\nabla h_{t}^{*}\cdot\nabla\log p_{T-t}\|_{L^{2}(p_{T-t})}\leq\frac{12\sqrt{d}\bar{B}}{\sigma_{T-t}^{4}}.$
Proof of Lemma D.6.
By applying the triangular inequality, we have
$\|\Delta h_{t}^{*}+\nabla h_{t}^{*}\cdot\nabla\log p_{T-t}\|_{L^{2}(p_{T-t})}$
$\leq\|\Delta h_{t}^{*}\|_{L^{2}(p_{T-t})}+\|\nabla h_{t}^{*}\cdot\nabla\log p_{T-t}\|_{L^{2}(p_{T-t})}$
(D.13) $\leq\frac{6\sqrt{d}\bar{B}}{\sigma_{T-t}^{4}}+\frac{2\bar{B}}{\sigma_{T-t}^{2}}\|\nabla\log p_{T-t}\|_{L^{2}(p_{T-t})},$
where the last inequality holds from Lemmas C.4 and C.5. It remains to estimate the $L^{2}(p_{T-t})$-norm of the score $\nabla\log p_{T-t}$ in (D.13). Indeed,
$\|\nabla\log p_{T-t}\|_{L^{2}(p_{T-t})}^{2}$
$=\int\|\frac{\mathbf{x}}{\sigma_{T-t}^{2}}-\frac{\mu_{T-t}}{\sigma_{T-t}^{2}}{\mathbb{E}}[\mathbf{X}_{0}|\mathbf{X}_{T-t}=\mathbf{x}]\|_{2}^{2}p_{T-t}(\mathbf{x})\,\mathrm{d}\mathbf{x}$
$\leq\frac{2}{\sigma_{T-t}^{4}}{\mathbb{E}}\big[\|\mathbf{X}_{T-t}\|_{2}^{2}\big]+\frac{2d\mu_{T-t}^{2}}{\sigma_{T-t}^{4}}$
$=\frac{2}{\sigma_{T-t}^{4}}\Big\{\mu_{T-t}^{2}{\mathbb{E}}\big[\|\mathbf{X}_{0}\|_{2}^{2}\big]+2\mu_{T-t}\sigma_{T-t}{\mathbb{E}}\big[\langle\mathbf{X}_{0},{\bm{\varepsilon}}\rangle\big]+\sigma_{T-t}^{2}{\mathbb{E}}\big[\|{\bm{\varepsilon}}\|_{2}^{2}\big]\Big\}+\frac{2d\mu_{T-t}^{2}}{\sigma_{T-t}^{4}}$
(D.14) $=\frac{2}{\sigma_{T-t}^{4}}\Big\{\mu_{T-t}^{2}{\mathbb{E}}\big[\|\mathbf{X}_{0}\|_{2}^{2}\big]+\sigma_{T-t}^{2}{\mathbb{E}}\big[\|{\bm{\varepsilon}}\|_{2}^{2}\big]\Big\}+\frac{2d\mu_{T-t}^{2}}{\sigma_{T-t}^{4}}\leq\frac{6d}{\sigma_{T-t}^{4}},$
where the first equality is owing to Lemma C.2, the first inequality used Assumption 1, and the second and third equalities hold from $\mathbf{X}_{T-t}\stackrel{{\scriptstyle\,\mathrm{d}}}{{=}}\mu_{T-t}\mathbf{X}_{0}+\sigma_{T-t}{\bm{\varepsilon}}$ where $\mathbf{X}_{0}$ is independent ${\bm{\varepsilon}}\sim{\mathcal{N}}({\mathbf{0}},\mathbf{I}_{d})$. The last inequality also uses Assumption 1. Substituting (D.14) into (D.13) completes the proof. ∎
D.3  Convergence rate of the Doob’s guidance estimation
Lemma D.7.
Suppose Assumptions 1 and 2 hold. Let $R\geq 1$, and let the hypothesis class ${\mathscr{H}}_{t}$ be defined as (5.2) with $L\leq C\log N$ and $S\leq N^{d}$, then there exists $h_{t}\in{\mathscr{H}}_{t}$ such that
$\|h_{t}-h_{t}^{*}\|_{L^{\infty}(B({\mathbf{0}},R))}$ $\leq\frac{C\bar{B}R^{2}}{\sigma_{T-t}^{4}N^{2}},$
$\|\nabla h_{t}-\nabla h_{t}^{*}\|_{L^{\infty}(B({\mathbf{0}},R))}$ $\leq\frac{C\bar{B}R}{\sigma_{T-t}^{4}N},$
where $C$ is a constant only depending on $d$.
Proof of Lemma D.7.
We first rescale the target function $h_{t}^{*}$ to $B({\mathbf{0}},1)$ by $g_{t}^{*}(\mathbf{z})\coloneq h_{t}^{*}(R\mathbf{z})$. According to Ding et al. [2025b, Lemma 6], there exists $g_{t}\in N(L,S)$ such that
$\|g_{t}-g_{t}^{*}\|_{L^{\infty}(B({\mathbf{0}},1))}$ $\leq\frac{C^{\prime}}{N^{2}}\|g_{t}^{*}\|_{C^{2}({\mathbb{R}}^{d})},$
$\|\nabla g_{t}-\nabla g_{t}^{*}\|_{L^{\infty}(B({\mathbf{0}},1))}$ $\leq\frac{C^{\prime}}{N}\|g_{t}^{*}\|_{C^{2}({\mathbb{R}}^{d})},$
where $C^{\prime}$ is a constant only depending on $d$. Note that $D_{k}g_{t}^{*}(\mathbf{z})=RD_{k}h_{t}^{*}(R\mathbf{z})$ for each $1\leq k\leq d$, and $D_{k\ell}^{2}g_{t}^{*}(\mathbf{z})=R^{2}D_{k\ell}^{2}h_{t}^{*}(R\mathbf{z})$ for each $1\leq k,\ell\leq d$. Thus
$\|g_{t}(R^{-1}\cdot)-h_{t}^{*}\|_{L^{\infty}(B({\mathbf{0}},R))}$ $=\|g_{t}(R^{-1}\cdot)-g_{t}^{*}(R^{-1}\cdot)\|_{L^{\infty}(B({\mathbf{0}},1))}\leq\frac{C^{\prime}R^{2}}{N^{2}}\|h_{t}^{*}\|_{C^{2}({\mathbb{R}}^{d})},$
$\|\nabla g_{t}(R^{-1}\cdot)-\nabla h_{t}^{*}\|_{L^{\infty}(B({\mathbf{0}},R))}$ $=\frac{1}{R}\|\nabla g_{t}(R^{-1}\cdot)-\nabla g_{t}^{*}(R^{-1}\cdot)\|_{L^{\infty}(B({\mathbf{0}},1))}\leq\frac{C^{\prime}R}{N}\|h_{t}^{*}\|_{C^{2}({\mathbb{R}}^{d})}.$
Setting $h_{t}\coloneq g_{t}(R^{-1}\cdot)$, and using Lemmas C.1, C.4, and C.5 complete the proof. ∎
Lemma D.8 (Approximation error).
Suppose Assumptions 1 and 2 hold. Let $R\geq 1$, and let the hypothesis class ${\mathscr{H}}_{t}$ be defined as (5.2) with $L\leq C\log N$ and $S\leq N^{d}$, then
$\|h_{t}-h_{t}^{*}\|_{L^{2}(p_{T-t})}^{2}$ $\leq C\frac{\bar{B}^{2}\log^{4}N}{\sigma_{T-t}^{8}N^{4}},$
$\|\nabla h_{t}-\nabla h_{t}^{*}\|_{L^{2}(p_{T-t})}^{2}$ $\leq C\frac{\bar{B}^{2}\log^{2}N}{\sigma_{T-t}^{8}N^{2}}.$
provided that $R^{2}=(4d\mu_{t}^{2}+8\sigma_{t}^{2})\log N^{4}$, where $C$ is a constant only depending on $d$.
Proof of Lemma D.8.
It is straightforward that for each $R\geq 1$,
(D.15) $\|h_{t}-h_{t}^{*}\|_{L^{2}(p_{T-t})}^{2}$ $=\underbrace{\int(h_{t}(\mathbf{x})-h_{t}^{*}(\mathbf{x}))^{2}{\mathbbm{1}}\{\|\mathbf{x}\|_{2}\leq R\}p_{T-t}(\mathbf{x})\,\mathrm{d}\mathbf{x}}_{\text{(i)}}$
$\quad+\underbrace{\int(h_{t}(\mathbf{x})-h_{t}^{*}(\mathbf{x}))^{2}{\mathbbm{1}}\{\|\mathbf{x}\|_{2}>R\}p_{T-t}(\mathbf{x})\,\mathrm{d}\mathbf{x}}_{\text{(ii)}}.$
For term (i) in (D.15), we have
$\int(h_{t}(\mathbf{x})-h_{t}^{*}(\mathbf{x}))^{2}{\mathbbm{1}}\{\|\mathbf{x}\|_{2}\leq R\}p_{T-t}(\mathbf{x})\,\mathrm{d}\mathbf{x}$
(D.16) $\leq\sup_{\|\mathbf{x}\|_{2}\leq R}(h_{t}(\mathbf{x})-h_{t}^{*}(\mathbf{x}))^{2}\leq\frac{C^{2}\bar{B}^{2}R^{4}}{\sigma_{T-t}^{8}N^{4}},$
where the second inequality holds from Lemma D.7. For term (ii) in (D.15), we have
$\int\|h_{t}(\mathbf{x})-h_{t}^{*}(\mathbf{x})\|_{2}^{2}{\mathbbm{1}}\{\|\mathbf{x}\|_{2}>R\}p_{T-t}(\mathbf{x})\,\mathrm{d}\mathbf{x}$
(D.17) $\leq 4\bar{B}^{2}\operatorname{Pr}\{\|\mathbf{X}_{T-t}\|_{2}>R\}\leq 2^{d+3}\bar{B}^{2}\exp\Big(-\frac{R^{2}}{4d\mu_{t}^{2}+8\sigma_{t}^{2}}\Big),$
where the first inequality holds from Lemma C.1, and the second inequality is due to Lemma G.1. Substituting (D.16) and (D.17) into (D.15) yields
(D.18) $\|h_{t}-h_{t}^{*}\|_{L^{2}(p_{T-t})}^{2}\leq\frac{C^{2}\bar{B}^{2}R^{4}}{\sigma_{T-t}^{8}N^{4}}+2^{d+3}\bar{B}^{2}\exp\Big(-\frac{R^{2}}{4d\mu_{t}^{2}+8\sigma_{t}^{2}}\Big).$
Similarly, for the gradient term, we have
(D.19) $\|\nabla h_{t}-\nabla h_{t}^{*}\|_{L^{2}(p_{T-t})}^{2}$ $=\underbrace{\int\|\nabla h_{t}(\mathbf{x})-\nabla h_{t}^{*}(\mathbf{x})\|_{2}^{2}{\mathbbm{1}}\{\|\mathbf{x}\|_{2}\leq R\}p_{T-t}(\mathbf{x})\,\mathrm{d}\mathbf{x}}_{\text{(i)}}$
$\quad+\underbrace{\int\|\nabla h_{t}(\mathbf{x})-\nabla h_{t}^{*}(\mathbf{x})\|_{2}^{2}{\mathbbm{1}}\{\|\mathbf{x}\|_{2}>R\}p_{T-t}(\mathbf{x})\,\mathrm{d}\mathbf{x}}_{\text{(ii)}}.$
For term (i) in (D.19), we have
$\int\|\nabla h_{t}(\mathbf{x})-\nabla h_{t}^{*}(\mathbf{x})\|_{2}^{2}{\mathbbm{1}}\{\|\mathbf{x}\|_{2}\leq R\}p_{T-t}(\mathbf{x})\,\mathrm{d}\mathbf{x}$
(D.20) $\leq\sup_{\|\mathbf{x}\|_{2}\leq R}\|\nabla h_{t}(\mathbf{x})-\nabla h_{t}^{*}(\mathbf{x})\|_{2}^{2}\leq\frac{C^{2}\bar{B}^{2}R^{2}}{\sigma_{T-t}^{8}N^{2}},$
where the second inequality holds from Lemma D.7. For term (ii) in (D.19), we have
$\int\|\nabla h_{t}(\mathbf{x})-\nabla h_{t}^{*}(\mathbf{x})\|_{2}^{2}{\mathbbm{1}}\{\|\mathbf{x}\|_{2}>R\}p_{T-t}(\mathbf{x})\,\mathrm{d}\mathbf{x}$
(D.21) $\leq\frac{16\bar{B}^{2}}{\sigma_{T-t}^{4}}\operatorname{Pr}\{\|\mathbf{X}_{T-t}\|_{2}>R\}\leq 2^{d+5}\frac{\bar{B}^{2}}{\sigma_{T-t}^{4}}\exp\Big(-\frac{R^{2}}{4d\mu_{t}^{2}+8\sigma_{t}^{2}}\Big),$
where the first inequality holds from Lemma C.4, and the second inequality is due to Lemma G.1. Substituting (D.20) and (D.21) into (D.19) yields
(D.22) $\|\nabla h_{t}-\nabla h_{t}^{*}\|_{L^{2}(p_{T-t})}^{2}\leq\frac{C^{2}\bar{B}^{2}R^{2}}{\sigma_{T-t}^{8}N^{2}}+2^{d+5}\frac{\bar{B}^{2}}{\sigma_{T-t}^{4}}\exp\Big(-\frac{R^{2}}{4d\mu_{t}^{2}+8\sigma_{t}^{2}}\Big).$
Setting $R^{2}=(4d\mu_{t}^{2}+8\sigma_{t}^{2})\log N^{4}$ in (D.18) and (D.22) completes the proof. ∎
Lemma D.9 (Generalization error).
Suppose Assumptions 1 and 2 hold. Let the hypothesis class ${\mathscr{H}}_{t}$ be defined as (5.2), then
$\mathrm{VCdim}({\mathscr{H}}_{t})$ $\leq cLS\log(S),$
$\mathrm{VCdim}(\nabla{\mathscr{H}}_{t})$ $\leq cL^{2}S\log(LS),$
where $c$ is an absolute constant.
Proof of Lemma D.9.
Since ${\mathscr{H}}_{t}\subseteq N(L,S)$, using Bartlett et al. [2019, Theorem 7] implies
(D.23) $\mathrm{VCdim}({\mathscr{H}}_{t})\leq\mathrm{VCdim}(N(L,S))\leq c_{1}LS\log(S),$
where $c_{1}$ is an absolute constant. According to Ding et al. [2025b, Lemma 13], we have $\nabla{\mathscr{H}}_{t}\subseteq N(c_{2}L,c_{3}LS)$, where $c_{2}$ and $c_{3}$ are absolute constants. Using Bartlett et al. [2019, Theorem 7] again implies
(D.24) $\mathrm{VCdim}(\nabla{\mathscr{H}}_{t})\leq c_{4}L^{2}S\log(LS),$
where $c_{4}$ is an absolute constant. This completes the proof. ∎
Theorem 5.3.
Suppose Assumptions 1 and 2 hold. Let $t\in(0,T)$. Set the hypothesis class ${\mathscr{H}}_{t}$ as
${\mathscr{H}}_{t}\coloneq\left\{h_{t}\in N(L,S):\begin{aligned} &\sup_{\mathbf{x}\in{\mathbb{R}}^{d}}h_{t}(\mathbf{x})\leq\bar{B},\penalty 10000\ \inf_{\mathbf{x}\in{\mathbb{R}}^{d}}h_{t}(\mathbf{x})\geq\underline{B},\\ &\max_{1\leq k\leq d}\sup_{\mathbf{x}\in{\mathbb{R}}^{d}}|D_{k}h_{t}(\mathbf{x})|\leq 2\sigma_{T-t}^{-2}\bar{B}\end{aligned}\right\},$
where $L={\mathcal{O}}(\log n)$ and $S={\mathcal{O}}(n^{\frac{d}{d+8}})$. Let $\widehat{h}_{t}^{\lambda}$ be the gradient-regularized empirical risk minimizer defined as (4.6), and let $h_{t}^{*}$ be the Doob’s $h$-function defined as (3.10). Then the following inequality holds:
${\mathbb{E}}\Big[\|\nabla\log\widehat{h}_{t}^{\lambda}-\nabla\log h_{t}^{*}\|_{L^{2}(p_{T-t})}^{2}\Big]\leq C\sigma_{T-t}^{-8}n^{-\frac{2}{d+8}},$
provided that the regularization parameter $\lambda$ is set as $\lambda={\mathcal{O}}(n^{-\frac{2}{d+8}})$, where $C$ is a constant depending only on $d$, $\bar{B}$, and $\underline{B}$.
Proof of Theorem 5.3.
Substituting Lemmas D.8, and D.9 into Lemma 5.2 yields
${\mathbb{E}}\Big[\|\widehat{h}_{t}^{\lambda}-h_{t}^{*}\|_{L^{2}(p_{T-t})}^{2}\Big]$
$\leq C\frac{\log^{4}N}{\sigma_{T-t}^{8}N^{4}}+C\lambda\frac{\log^{2}N}{\sigma_{T-t}^{8}N^{2}}+C\Big(\frac{N^{d}\log^{2}N}{n\log^{-1}n}\Big)^{\frac{1}{2}}+C\frac{\lambda}{\sigma_{T-t}^{4}}\Big(\frac{N^{d}\log^{4}N}{n\log^{-1}n}\Big)^{\frac{1}{2}}+C\frac{\lambda^{2}}{\sigma_{T-t}^{8}},$
where $C$ is a constant only depending on $d$ and $\bar{B}$, and we used the fact $L\leq C^{\prime}\log N$ and $S\leq N^{d}$ in Lemma D.8. Similarly,
${\mathbb{E}}\big[\|\nabla\widehat{h}_{t}^{\lambda}-\nabla h_{t}^{*}\|_{L^{2}(p_{T-t})}^{2}\big]$
$\leq\frac{C}{\lambda}\frac{\log^{4}N}{\sigma_{T-t}^{8}N^{4}}+C\frac{\log^{2}N}{\sigma_{T-t}^{8}N^{2}}+\frac{C}{\lambda}\Big(\frac{N^{d}\log^{2}N}{n\log^{-1}n}\Big)^{\frac{1}{2}}+\frac{C}{\sigma_{T-t}^{4}}\Big(\frac{N^{d}\log^{4}N}{n\log^{-1}n}\Big)^{\frac{1}{2}}+\frac{C\lambda}{\sigma_{T-t}^{8}}.$
By setting $N={\mathcal{O}}(n^{\frac{1}{d+8}})$ and $\lambda={\mathcal{O}}(n^{-\frac{2}{d+8}})$, we have
(D.25) ${\mathbb{E}}\big[\|\widehat{h}_{t}^{\lambda}-h_{t}^{*}\|_{L^{2}(p_{T-t})}^{2}\big]$ $\lesssim\frac{1}{\sigma_{T-t}^{8}}n^{-\frac{4}{d+8}}\log^{4}n,$
${\mathbb{E}}\big[\|\nabla\widehat{h}_{t}^{\lambda}-\nabla h_{t}^{*}\|_{L^{2}(p_{T-t})}^{2}\big]$ $\lesssim\frac{1}{\sigma_{T-t}^{8}}n^{-\frac{2}{d+8}}\log^{4}n.$
Consequently,
$\|\nabla\log\widehat{h}_{t}^{\lambda}-\nabla\log h_{t}^{*}\|_{L^{2}(p_{T-t})}^{2}$
$=\Big\|\frac{\nabla\widehat{h}_{t}^{\lambda}}{\widehat{h}_{t}^{\lambda}}-\frac{\nabla h_{t}^{*}}{\widehat{h}_{t}^{\lambda}}+\frac{\nabla h_{t}^{*}}{\widehat{h}_{t}^{\lambda}}-\frac{\nabla h_{t}^{*}}{h_{t}^{*}}\Big\|_{L^{2}(p_{T-t})}^{2}$
$\leq 2\Big\|\frac{\nabla\widehat{h}_{t}^{\lambda}}{\widehat{h}_{t}^{\lambda}}-\frac{\nabla h_{t}^{*}}{\widehat{h}_{t}^{\lambda}}\Big\|_{L^{2}(p_{T-t})}^{2}+2\Big\|\frac{\nabla h_{t}^{*}}{\widehat{h}_{t}^{\lambda}}-\frac{\nabla h_{t}^{*}}{h_{t}^{*}}\Big\|_{L^{2}(p_{T-t})}^{2}$
$\leq\frac{2}{\underline{B}^{2}}\|\nabla\widehat{h}_{t}^{\lambda}-\nabla h_{t}^{*}\|_{L^{2}(p_{T-t})}^{2}+2\frac{\bar{B}^{2}}{\underline{B}^{4}}\|\widehat{h}_{t}^{\lambda}-h_{t}^{*}\|_{L^{2}(p_{T-t})}^{2}$
$\leq C^{\prime}\frac{1}{\sigma_{T-t}^{8}}n^{-\frac{2}{d+8}}\log^{4}n,$
where the second inequality is owing to Lemmas C.1 and C.4, and the last inequality holds from (D.25). This completes the proof. ∎
Appendix E Derivations in Section 5.3
E.1  Error decomposition of the controllable diffusion models
Lemma 5.5.
Suppose Assumptions 1, 2, and 3 hold. Let $\widehat{q}_{T-T_{0}}$ be the marginal density of $\widehat{\mathbf{Z}}_{T-T_{0}}^{\leftarrow}$ defined in (4.8). Then it follows that
$\mathrm{KL}(q_{T_{0}}\|\widehat{q}_{T-T_{0}})$ $\lesssim\frac{\bar{B}}{\underline{B}}\sum_{k=0}^{K-1}h{\mathbb{E}}^{{\mathbb{P}}}\left[\|\nabla\log\widehat{h}_{kh}(\mathbf{X}_{kh}^{\leftarrow})-\nabla\log h_{kh}^{*}(\mathbf{X}_{kh}^{\leftarrow})\|_{2}^{2}\right]$
$\quad+\frac{\bar{B}}{\underline{B}}T\varepsilon_{\mathrm{ref}}^{2}+d\exp(-T)+\frac{d^{2}T^{2}}{\sigma_{T_{0}}^{4}K},$
where the notation $\lesssim$ hides absolute constants.
Proof of Lemma 5.5.
According to Chen et al. [2023a, Proposition C.3], we have
(E.1) $\mathrm{KL}(q_{T_{0}}\|\widehat{q}_{T-T_{0}})$ $\lesssim\underbrace{\mathrm{KL}(q_{T}\|\gamma_{d})}_{\text{initial error}}+\underbrace{\sum_{k=0}^{K-1}h{\mathbb{E}}^{{\mathbb{Q}}}\left[\|\widehat{\mathbf{s}}(kh,\mathbf{Z}_{kh}^{\leftarrow})-\nabla\log p_{T-kh}(\mathbf{Z}_{kh}^{\leftarrow})\|_{2}^{2}\right]}_{\text{base score error}}$
$\quad+\underbrace{\sum_{k=0}^{K-1}h{\mathbb{E}}^{{\mathbb{Q}}}\left[\|\nabla\log\widehat{h}_{kh}(\mathbf{Z}_{kh}^{\leftarrow})-\nabla\log h_{kh}^{*}(\mathbf{Z}_{kh}^{\leftarrow})\|_{2}^{2}\right]}_{\text{Doob's guidance error}}$
$\quad+\underbrace{\sum_{k=0}^{K-1}\int_{kh}^{(k+1)h}{\mathbb{E}}^{{\mathbb{Q}}}\left[\|\nabla\log q_{T-t}(\mathbf{Z}_{t}^{\leftarrow})-\nabla\log q_{T-kh}(\mathbf{Z}_{kh}^{\leftarrow})\|_{2}^{2}\right]\,\mathrm{d}t}_{\text{discretization error}},$
where $\gamma_{d}$ is the density of a standard Gaussian distribution ${\mathcal{N}}({\mathbf{0}},\mathbf{I}_{d})$.
Step 1. Initial error in (E.1). Using Chen et al. [2023a, Lemma C.4], we have
(E.2) $\mathrm{KL}(q_{T}\|\gamma_{d})\lesssim d\exp(-T).$
Step 2. Reference score error and Doob’s guidance error in (E.1). Under Assumption 2, it is apparent that ${\mathbb{E}}^{{\mathbb{P}}}[w(\mathbf{X}_{T}^{\leftarrow})]\geq\underline{B}$ and $\sup_{\mathbf{x}\in{\mathbb{R}}^{d}}h^{*}(t,\mathbf{x})\leq\bar{B}$ for each $t\in(T_{0},T)$. Hence, the density ratio is uniformly bounded
(E.3) $\sup_{\mathbf{x}\in{\mathbb{R}}^{d}}\frac{q_{T-t}(\mathbf{x})}{p_{T-t}(\mathbf{x})}=\frac{h^{*}(t,\mathbf{x})}{{\mathbb{E}}^{{\mathbb{P}}}[w(\mathbf{X}_{T}^{\leftarrow})]}\leq\frac{\bar{B}}{\underline{B}}.$
For the reference score error term in (E.1), it follows for each $0\leq k\leq K-1$ that
${\mathbb{E}}^{{\mathbb{Q}}}\left[\|\widehat{\mathbf{s}}(kh,\mathbf{Z}_{kh}^{\leftarrow})-\nabla\log p_{T-kh}(\mathbf{Z}_{kh}^{\leftarrow})\|_{2}^{2}\right]$
$=\int\|\widehat{\mathbf{s}}(kh,\mathbf{z})-\nabla\log p_{T-kh}(\mathbf{z})\|_{2}^{2}q_{T-kh}(\mathbf{z})\,\mathrm{d}\mathbf{z}$
$=\int\|\widehat{\mathbf{s}}(kh,\mathbf{x})-\nabla\log p_{T-kh}(\mathbf{x})\|_{2}^{2}\frac{q_{T-kh}(\mathbf{x})}{p_{T-kh}(\mathbf{x})}p_{T-kh}(\mathbf{x})\,\mathrm{d}\mathbf{x}$
$\leq\frac{\bar{B}}{\underline{B}}\int\|\widehat{\mathbf{s}}(kh,\mathbf{x})-\nabla\log p_{T-kh}(\mathbf{x})\|_{2}^{2}p_{T-kh}(\mathbf{x})\,\mathrm{d}\mathbf{x}$
$=\frac{\bar{B}}{\underline{B}}{\mathbb{E}}^{{\mathbb{P}}}\left[\|\widehat{\mathbf{s}}(kh,\mathbf{X}_{kh}^{\leftarrow})-\nabla\log p_{T-kh}(\mathbf{X}_{kh}^{\leftarrow})\|_{2}^{2}\right],$
where the inequality holds from (E.3) and Hölder’s inequality. Consequently,
$\sum_{k=0}^{K-1}h{\mathbb{E}}^{{\mathbb{Q}}}\left[\|\widehat{\mathbf{s}}(kh,\mathbf{Z}_{kh}^{\leftarrow})-\nabla\log p_{T-kh}(\mathbf{Z}_{kh}^{\leftarrow})\|_{2}^{2}\right]$
(E.4) $\leq\frac{\bar{B}}{\underline{B}}\sum_{k=0}^{K-1}h{\mathbb{E}}^{{\mathbb{P}}}\left[\|\widehat{\mathbf{s}}(kh,\mathbf{X}_{kh}^{\leftarrow})-\nabla\log p_{T-kh}(\mathbf{X}_{kh}^{\leftarrow})\|_{2}^{2}\right]\leq T\frac{\bar{B}}{\underline{B}}\varepsilon_{\mathrm{ref}}^{2},$
where the last inequality is owing to Assumption 3. By a similar argument, we have
$\sum_{k=0}^{K-1}h{\mathbb{E}}^{{\mathbb{Q}}}\left[\|\nabla\log\widehat{h}_{kh}(\mathbf{Z}_{kh}^{\leftarrow})-\nabla\log h_{kh}^{*}(\mathbf{Z}_{kh}^{\leftarrow})\|_{2}^{2}\right]$
(E.5) $\leq\frac{\bar{B}}{\underline{B}}\sum_{k=0}^{K-1}h{\mathbb{E}}^{{\mathbb{P}}}\left[\|\nabla\log\widehat{h}_{kh}(\mathbf{X}_{kh}^{\leftarrow})-\nabla\log h_{kh}^{*}(\mathbf{X}_{kh}^{\leftarrow})\|_{2}^{2}\right].$
Step 3. Discretization error in (E.1). According to Chen et al. [2023a, Lemma D.1], we have
(E.6) ${\mathbb{E}}^{{\mathbb{Q}}}\big[\|\nabla\log q_{T-t}(\mathbf{Z}_{t}^{\leftarrow})-\nabla\log q_{T-kh}(\mathbf{Z}_{kh}^{\leftarrow})\|_{2}^{2}\big]\lesssim\frac{dG_{k}T}{K},$
for any $t\in(kh,(k+1)h)$, provided that $\nabla\log q_{T-t}$ is $G$-Lipschitz for any $t\in(kh,(k+1)h)$. Then it remains to estimate the Lipschitz constant $G_{k}$. Using Lemmas C.2 and C.3 yields
$\nabla^{2}\log q_{T-t}(\mathbf{z})$ $=-\frac{1}{\sigma_{T-t}^{2}}\mathbf{I}_{d}+\frac{\mu_{T-t}}{\sigma_{T-t}^{2}}\nabla{\mathbb{E}}[\mathbf{Z}_{0}|\mathbf{Z}_{t}=\mathbf{z}]$
$=-\frac{1}{\sigma_{T-t}^{2}}\mathbf{I}_{d}+\frac{\mu_{T-t}^{2}}{\sigma_{T-t}^{4}}\mathrm{Cov}(\mathbf{Z}_{0}|\mathbf{Z}_{t}=\mathbf{z}).$
As a consequence, for each $0\leq k\leq K-1$,
$G_{k}$ $\leq\sup_{t\in(T_{0},T)}\sup_{\mathbf{z}\in{\mathbb{R}}^{d}}\|\nabla^{2}\log q_{T-t}(\mathbf{z})\|_{\mathrm{op}}$
(E.7) $\leq\sup_{t\in(T_{0},T)}\frac{1}{\sigma_{T-t}^{2}}+\frac{\mu_{T-t}^{2}}{\sigma_{T-t}^{4}}\|\mathrm{Cov}(\mathbf{Z}_{0}|\mathbf{Z}_{t}=\mathbf{z})\|_{\mathrm{op}}\lesssim\frac{d}{\sigma_{T_{0}}^{4}},$
where the first inequality holds from the triangular inequality, and the second inequality is due to the boundedness of $\mathbf{Z}_{0}$ under Assumptions 1 and 2. Combining (E.6) and (E.7) implies
(E.8) $\sum_{k=0}^{K-1}\int_{kh}^{(k+1)h}{\mathbb{E}}^{{\mathbb{Q}}}\left[\|\nabla\log q_{T-t}(\mathbf{Z}_{t}^{\leftarrow})-\nabla\log q_{T-kh}(\mathbf{Z}_{kh}^{\leftarrow})\|_{2}^{2}\right]\,\mathrm{d}t\lesssim\frac{d^{2}T^{2}}{\sigma_{T_{0}}^{4}K}.$
Step 4. Conclusions. Substituting (E.2), (E.4), (E.5), (E.8) into (E.1) completes the proof. ∎
Corollary E.1.
Suppose Assumptions 1, 2, and 3 hold. Let $\delta\in(0,1)$. Set the hypothesis classes $\{{\mathscr{H}}_{T-kh}\}_{k=0}^{K=1}$ as (5.2) with the same depth $L$ and number of non-zero parameters $S$. Let $\widehat{q}_{T-T_{0}}$ be the marginal density of $\widehat{\mathbf{Z}}_{T-T_{0}}^{\leftarrow}$ defined in (4.8). Then it follows that
$\|q_{T_{0}}-\widehat{q}_{T-T_{0}}\|_{\mathrm{TV}}^{2}\leq\frac{C\delta^{2}}{\sigma_{T_{0}}^{8}}\log\Big(\frac{1}{\delta}\log\Big(\frac{\sigma_{T_{0}}^{8}}{\delta^{2}}\Big)\Big),$
where $C$ is a constant depending only on $d$, $\bar{B}$, and $\underline{B}$, and
$T\asymp\log\Big(\frac{\sigma_{T_{0}}^{8}}{\delta^{2}}\Big),\quad K\gtrsim\frac{\sigma_{T_{0}}^{4}}{\delta^{2}}\log^{2}\Big(\frac{\sigma_{T_{0}}^{8}}{\delta^{2}}\Big),\quad h\lesssim\frac{\delta^{2}}{\sigma_{T_{0}}^{4}}\log^{-1}\Big(\frac{\sigma_{T_{0}}^{8}}{\delta^{2}}\Big)$
$\varepsilon_{\mathrm{ref}}^{2}\lesssim\frac{\delta^{2}}{\sigma_{T_{0}}^{8}}\log^{-1}\Big(\frac{\sigma_{T_{0}}^{8}}{\delta^{2}}\Big),\quad n\gtrsim\frac{1}{\delta^{d+8}}\log^{\frac{d+8}{2}}\Big(\frac{\sigma_{T_{0}}^{8}}{\delta^{2}}\Big).$
Proof of Corollary E.1.
Combining Theorem 5.3 and Lemma 5.5 yields
$\mathrm{KL}(q_{T_{0}}\|\widehat{q}_{T-T_{0}})$ $\leq\frac{C}{\sigma_{T_{0}}^{8}}\Big\{\underbrace{Tn^{-\frac{2}{d+8}}\log^{4}n}_{\text{(i)}}+\underbrace{T\sigma_{T_{0}}^{8}\varepsilon_{\mathrm{ref}}^{2}}_{\text{(ii)}}+\underbrace{\sigma_{T_{0}}^{8}\exp(-T)}_{\text{(iii)}}+\underbrace{T^{2}\sigma_{T_{0}}^{4}\frac{1}{K}}_{\text{(iv)}}\Big\},$
where $C$ is a constant depending only on $d$, $\bar{B}$, and $\underline{B}$. By setting
$T\asymp\log\Big(\frac{\sigma_{T_{0}}^{8}}{\delta^{2}}\Big),\quad K\gtrsim\frac{\sigma_{T_{0}}^{4}}{\delta^{2}}\log^{2}\Big(\frac{\sigma_{T_{0}}^{8}}{\delta^{2}}\Big),$
$\varepsilon_{\mathrm{ref}}^{2}\lesssim\frac{\delta^{2}}{\sigma_{T_{0}}^{8}}\log^{-1}\Big(\frac{\sigma_{T_{0}}^{8}}{\delta^{2}}\Big),\quad n\gtrsim\frac{1}{\delta^{d+8}}\log^{\frac{d+8}{2}}\Big(\frac{\sigma_{T_{0}}^{8}}{\delta^{2}}\Big),$
we find
$\mathrm{KL}(q_{T_{0}}\|\widehat{q}_{T-T_{0}})\leq\frac{C\delta^{2}}{\sigma_{T_{0}}^{8}}\log\Big(\frac{1}{\delta}\log\Big(\frac{\sigma_{T_{0}}^{8}}{\delta^{2}}\Big)\Big).$
Finally, using Pinker’s inequality completes the proof. ∎
E.2  Convergence rate of the controllable diffusion models
Theorem 5.6.
Suppose Assumptions 1, 2, and 3 hold. Let $\varepsilon\in(0,1)$. Set the hypothesis classes $\{{\mathscr{H}}_{T-kh}\}_{k=0}^{K=1}$ as (5.2) with the same depth $L$ and number of non-zero parameters $S$ as Theorem 5.3. Let $\widehat{q}_{T-T_{0}}$ be the marginal density of $\widehat{\mathbf{Z}}_{T-T_{0}}^{\leftarrow}$ defined in (4.8), and let $({\mathcal{M}}\circ{\mathcal{T}}_{R})_{\sharp}\widehat{q}_{T-T_{0}}$ defined as (4.9). Then it follows that
(E.9) ${\mathbb{E}}\Big[{\mathcal{W}}_{2}^{2}(q_{0},({\mathcal{M}}\circ{\mathcal{T}}_{R})_{\sharp}\widehat{q}_{T-T_{0}})\Big]\leq C\varepsilon\log^{3}\Big(\frac{1}{\varepsilon}\Big).$
provided that the truncation radius $R$, the terminal time $T$, the step size $h$, the number of steps $K$, the error of reference score $\varepsilon_{\mathrm{ref}}$, the number of samples $n$ for Doob’s matching, and the early-stopping time $T_{0}$ are set, respectively, as
$R\asymp\log^{\frac{1}{2}}\Big(\frac{1}{\varepsilon}\Big),\quad T\asymp\log\Big(\frac{1}{\varepsilon^{2}}\Big),\quad K\gtrsim\frac{1}{\varepsilon^{4}}\log^{2}\Big(\frac{1}{\varepsilon^{2}}\Big),\quad h\lesssim\varepsilon^{4}\log^{-1}\Big(\frac{1}{\varepsilon^{2}}\Big)$
$\varepsilon_{\mathrm{ref}}^{2}\lesssim\varepsilon^{2}\log^{-1}\Big(\frac{1}{\varepsilon^{2}}\Big),\quad n\gtrsim\frac{1}{\varepsilon^{3(d+8)}}\log^{\frac{d+8}{2}}\Big(\frac{1}{\varepsilon^{2}}\Big).$
Here $C$ is a constant depending only on $d$, $\bar{B}$, and $\underline{B}$.
Proof of Theorem 5.6.
According to the triangular inequality, we have
(E.10) ${\mathcal{W}}_{2}^{2}(q_{0},({\mathcal{M}}\circ{\mathcal{T}}_{R})_{\sharp}\widehat{q}_{T-T_{0}})$ $=3\underbrace{{\mathcal{W}}_{2}^{2}(q_{0},{\mathcal{M}}_{\sharp}q_{T_{0}})}_{\text{(i)}}+3\underbrace{{\mathcal{W}}_{2}^{2}({\mathcal{M}}_{\sharp}q_{T_{0}},({\mathcal{M}}\circ{\mathcal{T}}_{R})_{\sharp}q_{T_{0}})}_{\text{(ii)}}$
$\quad+3\underbrace{{\mathcal{W}}_{2}^{2}(({\mathcal{M}}\circ{\mathcal{T}}_{R})_{\sharp}q_{T_{0}},({\mathcal{M}}\circ{\mathcal{T}}_{R})_{\sharp}\widehat{q}_{T-T_{0}})}_{\text{(iii)}}.$
Here the term (i) represents the early-stopping error, the term (ii) represents the truncation error, while the term (iii) represents the error of controllable diffusion models (4.8). In the rest of the proof, we bound these three errors, respectively.
Step 1. Bound the term (i) in (E.10). To estimate the 2-Wasserstein distance between the target distribution $q_{0}$ and the scaled early-stopping distribution ${\mathcal{M}}_{\sharp}q_{T_{0}}$, we begin by producing a coupling of them. Let $\mathbf{Z}_{0}\sim q_{0}$, and let ${\bm{\varepsilon}}\sim{\mathcal{N}}({\mathbf{0}},\mathbf{I}_{d})$ be independent of $\mathbf{Z}_{0}$. Define $\widetilde{\mathbf{Z}}_{T_{0}}\coloneq\mathbf{Z}_{0}+\sigma_{T_{0}}\mu_{T_{0}}^{-1}{\bm{\varepsilon}}$. It is apparent that $\widetilde{\mathbf{Z}}_{T_{0}}\sim{\mathcal{M}}_{\sharp}q_{T_{0}}$. Then
(E.11) ${\mathcal{W}}_{2}^{2}(q_{0},{\mathcal{M}}_{\sharp}q_{T_{0}})\leq{\mathbb{E}}\big[\|\mathbf{Z}_{0}-\widetilde{\mathbf{Z}}_{T_{0}}\|_{2}^{2}\big]=\frac{\sigma_{T_{0}}^{2}}{\mu_{T_{0}}^{2}}{\mathbb{E}}\big[\|{\bm{\varepsilon}}\|_{2}^{2}\big]=\frac{d\sigma_{T_{0}}^{2}}{\mu_{T_{0}}^{2}}.$
Step 2. Bound the term (ii) in (E.10). Let $\mathbf{Z}_{T_{0}}\sim q_{T_{0}}$. According to the definition of the truncation operator ${\mathcal{T}}_{R}$, the joint law of $(\mu_{T_{0}}^{-1}\mathbf{Z}_{T_{0}},\mu_{T_{0}}^{-1}\mathbf{Z}_{T_{0}}{\mathbbm{1}}_{B({\mathbf{0}},R)}(\mathbf{Z}_{T_{0}}))$ is a coupling of ${\mathcal{M}}_{\sharp}q_{T_{0}}$ and $({\mathcal{M}}\circ{\mathcal{T}}_{R})_{\sharp}q_{T_{0}}$. Therefore,
${\mathcal{W}}_{2}^{2}({\mathcal{M}}_{\sharp}q_{T_{0}},({\mathcal{M}}\circ{\mathcal{T}}_{R})_{\sharp}q_{T_{0}})$ $\leq{\mathbb{E}}\big[\|\mu_{T_{0}}^{-1}\mathbf{Z}_{T_{0}}-\mu_{T_{0}}^{-1}\mathbf{Z}_{T_{0}}{\mathbbm{1}}_{B({\mathbf{0}},R)}(\mathbf{Z}_{T_{0}})\|_{2}^{2}\big]$
$=\frac{1}{\mu_{T_{0}}^{2}}\int\|\mathbf{z}-\mathbf{z}{\mathbbm{1}}_{B({\mathbf{0}},R)}(\mathbf{z})\|_{2}^{2}q_{T_{0}}(\mathbf{z})\,\mathrm{d}\mathbf{z}$
$=\frac{1}{\mu_{T_{0}}^{2}}\int\|\mathbf{z}\|_{2}^{2}{\mathbbm{1}}_{{\mathbb{R}}^{d}-B({\mathbf{0}},R)}(\mathbf{z})q_{T_{0}}(\mathbf{z})\,\mathrm{d}\mathbf{z}$
$\leq\frac{1}{\mu_{T_{0}}^{2}}{\mathbb{E}}^{\frac{1}{2}}\big[\|\mathbf{Z}_{T_{0}}\|_{2}^{4}\big]\operatorname{Pr}^{\frac{1}{2}}\{\|\mathbf{Z}_{T_{0}}\|_{2}>R\}$
$\lesssim\frac{1}{\mu_{T_{0}}^{2}}d2^{d+1}\exp\Big(-\frac{R^{2}}{4d\mu_{T_{0}}^{2}+8\sigma_{T_{0}}^{2}}\Big),$
where the second ineq holds from Cauchy-Schwarz inequality, and the last inequality is due to Lemma G.3 and Corollary G.2. By setting $R^{2}=(4d\mu_{T_{0}}^{2}+8\sigma_{T_{0}}^{2})\log(\varepsilon^{-1})$, we have
(E.12) ${\mathcal{W}}_{2}^{2}({\mathcal{M}}_{\sharp}q_{T_{0}},({\mathcal{M}}\circ{\mathcal{T}}_{R})_{\sharp}q_{T_{0}})\lesssim\frac{d2^{d}}{\mu_{T_{0}}^{2}}\varepsilon.$
Step 3. Bound the term (iii) in (E.10). Let $\mathbf{Z}_{T_{0}}^{R}\sim({\mathcal{T}}_{R})_{\sharp}q_{T_{0}}$ and $\widehat{\mathbf{Z}}_{T_{0}}^{R}\sim({\mathcal{T}}_{R})_{\sharp}\widehat{q}_{T-T_{0}}$ be optimal coupled. This means
(E.13) ${\mathcal{W}}_{2}^{2}(({\mathcal{T}}_{R})_{\sharp}q_{T_{0}},({\mathcal{T}}_{R})_{\sharp}\widehat{q}_{T-T_{0}})={\mathbb{E}}\big[\|\mathbf{Z}_{T_{0}}^{R}-\widehat{\mathbf{Z}}_{T_{0}}^{R}\|_{2}^{2}\big].$
On the other hand, $\mu_{T_{0}}^{-1}\mathbf{Z}_{T_{0}}^{R}\sim({\mathcal{M}}\circ{\mathcal{T}}_{R})_{\sharp}q_{T_{0}}$ and $\mu_{T_{0}}^{-1}\widehat{\mathbf{Z}}_{T_{0}}^{R}\sim({\mathcal{M}}\circ{\mathcal{T}}_{R})_{\sharp}\widehat{q}_{T-T_{0}}$. Hence,
${\mathcal{W}}_{2}^{2}(({\mathcal{M}}\circ{\mathcal{T}}_{R})_{\sharp}q_{T_{0}},({\mathcal{M}}\circ{\mathcal{T}}_{R})_{\sharp}\widehat{q}_{T-T_{0}})$
(E.14) $\leq{\mathbb{E}}\big[\|\mu_{T_{0}}^{-1}\mathbf{Z}_{T_{0}}^{R}-\mu_{T_{0}}^{-1}\widehat{\mathbf{Z}}_{T_{0}}^{R}\|_{2}^{2}\big]=\frac{1}{\mu_{T_{0}}^{2}}{\mathcal{W}}_{2}^{2}(({\mathcal{T}}_{R})_{\sharp}q_{T_{0}},({\mathcal{T}}_{R})_{\sharp}\widehat{q}_{T-T_{0}}),$
where the equality holds from (E.13). Then using Villani [2009, Theorem 6.15] and the data processing inequality, we have
${\mathcal{W}}_{2}^{2}(({\mathcal{T}}_{R})_{\sharp}q_{T_{0}},({\mathcal{T}}_{R})_{\sharp}\widehat{q}_{T-T_{0}})$ $=2R^{2}\|({\mathcal{T}}_{R})_{\sharp}q_{T_{0}}-({\mathcal{T}}_{R})_{\sharp}\widehat{q}_{T-T_{0}}\|_{\mathrm{TV}}$
(E.15) $\leq 2R^{2}\|q_{T_{0}}-\widehat{q}_{T-T_{0}}\|_{\mathrm{TV}}.$
Combining (E.14) and (E.15) yields
${\mathcal{W}}_{2}^{2}(({\mathcal{M}}\circ{\mathcal{T}}_{R})_{\sharp}q_{T_{0}},({\mathcal{M}}\circ{\mathcal{T}}_{R})_{\sharp}\widehat{q}_{T-T_{0}})$ $\leq\frac{2R^{2}}{\mu_{T_{0}}^{2}}\|q_{T_{0}}-\widehat{q}_{T-T_{0}}\|_{\mathrm{TV}}$
$\leq\frac{2R^{2}}{\mu_{T_{0}}^{2}}\frac{C^{\prime}\varepsilon^{3}}{\sigma_{T_{0}}^{4}}$
(E.16) $\leq\frac{2(4d\mu_{T_{0}}^{2}+8\sigma_{T_{0}}^{2})\log(\varepsilon^{-1})}{\mu_{T_{0}}^{2}}\frac{C^{\prime}\varepsilon^{3}}{\sigma_{T_{0}}^{4}}\log^{2}\Big(\frac{1}{\varepsilon}\Big),$
where $C$ is a constant depending only on $d$, $\bar{B}$, and $\underline{B}$, and the second inequality holds from Corollary E.1 with $\delta=\varepsilon^{3}$.
Step 4. Conclusion. Substituting (E.11), (E.12), and (E.16) into (E.10) yields
${\mathcal{W}}_{2}^{2}(q_{0},({\mathcal{M}}\circ{\mathcal{T}}_{R})_{\sharp}\widehat{q}_{T-T_{0}})$ $\lesssim\frac{d\sigma_{T_{0}}^{2}}{\mu_{T_{0}}^{2}}+\frac{d2^{d}}{\mu_{T_{0}}^{2}}\varepsilon+\frac{2(4d\mu_{T_{0}}^{2}+8\sigma_{T_{0}}^{2})\log(\varepsilon^{-1})}{\mu_{T_{0}}^{2}}\frac{C^{\prime}\varepsilon^{3}}{\sigma_{T_{0}}^{4}}\log^{2}\Big(\frac{1}{\varepsilon}\Big)$
$\leq C\Big\{\sigma_{T_{0}}^{2}+\varepsilon+\frac{\varepsilon^{3}}{\sigma_{T_{0}}^{4}}\log^{3}\Big(\frac{1}{\varepsilon}\Big)\Big\},$
where $C$ is a constant depending only on $d$, $\bar{B}$, and $\underline{B}$. Letting $\sigma_{T_{0}}^{2}\asymp\varepsilon$, i.e., $T_{0}\asymp\varepsilon$, completes the proof. ∎
Appendix F Derivations in Section 5.4
Proposition 5.7.
Suppose Assumptions 4 and 2 hold. Then for any $t\in(0,T)$ and $\mathbf{x}\in{\mathbb{R}}^{d}$, we have
(F.1) $h_{t}^{*}(\mathbf{x})=\bar{h}_{t}^{*}(\mathbf{P}^{\top}\mathbf{x})\coloneq{\mathbb{E}}[w(\mathbf{P}\bar{\mathbf{X}}_{T}^{\leftarrow})\mid\bar{\mathbf{X}}_{t}^{\leftarrow}=\mathbf{P}^{\top}\mathbf{x}].$
Proof of Proposition 5.7.
According to Assumption 4, a particle $\mathbf{X}_{0}$ following $p_{0}$ satisfies
$\mathbf{X}_{0}\stackrel{{\scriptstyle\,\mathrm{d}}}{{=}}\mathbf{P}\bar{\mathbf{X}}_{0},\quad\bar{\mathbf{X}}_{0}\sim\bar{p}_{0}.$
We first establish the relations between $p_{t}$ and $\bar{p}_{t}$. It is straightforward that
$p_{t}(\mathbf{x})$ $=\int\varphi_{d}(\mathbf{x};\mu_{t}\mathbf{x}_{0},\sigma_{t}^{2}\mathbf{I}_{d})p_{0}(\mathbf{x}_{0})\,\mathrm{d}\mathbf{x}_{0}$
$=\int\varphi_{d}(\mathbf{x};\mu_{t}\mathbf{x}_{0},\sigma_{t}^{2}\mathbf{I}_{d})\Big(\int\delta_{\mathbf{P}\bar{\mathbf{x}}_{0}}(\mathbf{x}_{0})\bar{p}_{0}(\bar{\mathbf{x}}_{0})\,\mathrm{d}\bar{\mathbf{x}}_{0}\Big)\,\mathrm{d}\mathbf{x}_{0}$
$=\int\Big(\int\varphi_{d}(\mathbf{x};\mu_{t}\mathbf{x}_{0},\sigma_{t}^{2}\mathbf{I}_{d})\delta_{\mathbf{P}\bar{\mathbf{x}}_{0}}(\mathbf{x}_{0})\,\mathrm{d}\mathbf{x}_{0}\Big)\bar{p}_{0}(\bar{\mathbf{x}}_{0})\,\mathrm{d}\bar{\mathbf{x}}_{0}$
$=\int\varphi_{d}(\mathbf{x};\mu_{t}\mathbf{P}\bar{\mathbf{x}}_{0},\sigma_{t}^{2}\mathbf{I}_{d})\bar{p}_{0}(\bar{\mathbf{x}}_{0})\,\mathrm{d}\bar{\mathbf{x}}_{0}$
$=(2\pi\sigma_{t}^{2})^{-\frac{d}{2}}\int\exp\Big(-\frac{\|\mathbf{x}-\mu_{t}\mathbf{P}\bar{\mathbf{x}}_{0}\|_{2}^{2}}{2\sigma_{t}^{2}}\Big)\bar{p}_{0}(\bar{\mathbf{x}}_{0})\,\mathrm{d}\bar{\mathbf{x}}_{0}$
$=(2\pi\sigma_{t}^{2})^{-\frac{d}{2}}\int\exp\Big(-\frac{\|(\mathbf{I}_{d}-\mathbf{P}\mathbf{P}^{\top})\mathbf{x}+\mathbf{P}\mathbf{P}^{\top}\mathbf{x}-\mu_{t}\mathbf{P}\bar{\mathbf{x}}_{0}\|_{2}^{2}}{2\sigma_{t}^{2}}\Big)\bar{p}_{0}(\bar{\mathbf{x}}_{0})\,\mathrm{d}\bar{\mathbf{x}}_{0}$
$=(2\pi\sigma_{t}^{2})^{-\frac{d}{2}}\int\exp\Big(-\frac{\|(\mathbf{I}_{d}-\mathbf{P}\mathbf{P}^{\top})\mathbf{x}\|_{2}^{2}+\|\mathbf{P}\mathbf{P}^{\top}\mathbf{x}-\mu_{t}\mathbf{P}\bar{\mathbf{x}}_{0}\|_{2}^{2}}{2\sigma_{t}^{2}}\Big)\bar{p}_{0}(\bar{\mathbf{x}}_{0})\,\mathrm{d}\bar{\mathbf{x}}_{0}$
$=(2\pi\sigma_{t}^{2})^{-\frac{d}{2}}\exp\Big(-\frac{\|(\mathbf{I}_{d}-\mathbf{P}\mathbf{P}^{\top})\mathbf{x}\|_{2}^{2}}{2\sigma_{t}^{2}}\Big)\int\exp\Big(-\frac{\|\mathbf{P}^{\top}\mathbf{x}-\mu_{t}\bar{\mathbf{x}}_{0}\|_{2}^{2}}{2\sigma_{t}^{2}}\Big)\bar{p}_{0}(\bar{\mathbf{x}}_{0})\,\mathrm{d}\bar{\mathbf{x}}_{0}$
(F.2) $=\exp\Big(-\frac{\|(\mathbf{I}_{d}-\mathbf{P}\mathbf{P}^{\top})\mathbf{x}\|_{2}^{2}}{2\sigma_{t}^{2}}\Big)\bar{p}_{t}(\mathbf{P}^{\top}\mathbf{x}),$
where the seventh equality invokes the fact that $(\mathbf{I}_{d}-\mathbf{P}\mathbf{P}^{\top})\mathbf{x}$ is orthogonal to $\mathbf{P}\mathbf{P}^{\top}\mathbf{x}-\mu_{t}\mathbf{P}\bar{\mathbf{x}}_{0}$, the eighth equality is due to $\|\mathbf{P}\mathbf{v}\|_{2}=\|\mathbf{v}\|_{2}$ for each $\mathbf{v}\in{\mathbb{R}}^{d}$, and the last equality used (5.5). Then by a similar argument as the density, we find
$h_{T-t}^{*}(\mathbf{x})={\mathbb{E}}[w(\mathbf{X}_{0})\mid\mathbf{X}_{t}=\mathbf{x}]$
$=\frac{1}{p_{t}(\mathbf{x})}\int w(\mathbf{x}_{0})\varphi_{d}(\mathbf{x};\mu_{t}\mathbf{x}_{0},\sigma_{t}^{2}\mathbf{I}_{d})p_{0}(\mathbf{x}_{0})\,\mathrm{d}\mathbf{x}_{0}$
$=\frac{1}{p_{t}(\mathbf{x})}\int w(\mathbf{P}\bar{\mathbf{x}}_{0})\varphi_{d}(\mathbf{x};\mu_{t}\mathbf{P}\bar{\mathbf{x}}_{0},\sigma_{t}^{2}\mathbf{I}_{d})\bar{p}_{0}(\bar{\mathbf{x}}_{0})\,\mathrm{d}\bar{\mathbf{x}}_{0}$
$=\frac{1}{p_{t}(\mathbf{x})}(2\pi\sigma_{t}^{2})^{-\frac{d}{2}}\int w(\mathbf{P}\bar{\mathbf{x}}_{0})\exp\Big(-\frac{\|\mathbf{x}-\mu_{t}\mathbf{P}\bar{\mathbf{x}}_{0}\|_{2}^{2}}{2\sigma_{t}^{2}}\Big)\bar{p}_{0}(\bar{\mathbf{x}}_{0})\,\mathrm{d}\bar{\mathbf{x}}_{0}$
$=\frac{1}{p_{t}(\mathbf{x})}(2\pi\sigma_{t}^{2})^{-\frac{d}{2}}\int w(\mathbf{P}\bar{\mathbf{x}}_{0})\exp\Big(-\frac{\|(\mathbf{I}_{d}-\mathbf{P}\mathbf{P}^{\top})\mathbf{x}+\mathbf{P}\mathbf{P}^{\top}\mathbf{x}-\mu_{t}\mathbf{P}\bar{\mathbf{x}}_{0}\|_{2}^{2}}{2\sigma_{t}^{2}}\Big)\bar{p}_{0}(\bar{\mathbf{x}}_{0})\,\mathrm{d}\bar{\mathbf{x}}_{0}$
$=\frac{1}{p_{t}(\mathbf{x})}\exp\Big(-\frac{\|(\mathbf{I}_{d}-\mathbf{P}\mathbf{P}^{\top})\mathbf{x}\|_{2}^{2}}{2\sigma_{t}^{2}}\Big)\int w(\mathbf{P}\bar{\mathbf{x}}_{0})\varphi_{d}(\mathbf{P}^{\top}\mathbf{x};\mu_{t}\bar{\mathbf{x}}_{0},\sigma_{t}^{2}\mathbf{I}_{d})\bar{p}_{0}(\bar{\mathbf{x}}_{0})\,\mathrm{d}\bar{\mathbf{x}}_{0}$
$=\frac{1}{\bar{p}_{t}(\mathbf{P}^{\top}\mathbf{x})}\int w(\mathbf{P}\bar{\mathbf{x}}_{0})\varphi_{d}(\mathbf{P}^{\top}\mathbf{x};\mu_{t}\bar{\mathbf{x}}_{0},\sigma_{t}^{2}\mathbf{I}_{d})\bar{p}_{0}(\bar{\mathbf{x}}_{0})\,\mathrm{d}\bar{\mathbf{x}}_{0}$
$={\mathbb{E}}[w(\mathbf{P}\bar{\mathbf{X}}_{0})\mid\bar{\mathbf{X}}_{t}=\mathbf{P}^{\top}\mathbf{x}]={\mathbb{E}}[w(\mathbf{P}\bar{\mathbf{X}}_{T}^{\leftarrow})\mid\bar{\mathbf{X}}_{T-t}^{\leftarrow}=\mathbf{P}^{\top}\mathbf{x}],$
where the second and the eighth equalities are due to Bayes’ rule, the seventh equality holds from (F.2). This completes the proof. ∎
Proposition 5.8.
Suppose Assumptions 4 and 2 hold. Then for all $t\in(0,T)$ and $\bar{\mathbf{x}}\in{\mathbb{R}}^{d^{*}}$, the following bounds hold:
  1. (i)
    $\underline{B}\leq\bar{h}_{t}^{*}(\bar{\mathbf{x}})\leq\bar{B}$;
  2. (ii)
    $\max_{1\leq k\leq d}|D_{k}\bar{h}_{t}^{*}(\bar{\mathbf{x}})|\leq 2\sigma_{T-t}^{-2}\bar{B}$; and
  3. (iii)
    $\max_{1\leq k,\ell\leq d}|D_{k\ell}^{2}\bar{h}_{t}^{*}(\bar{\mathbf{x}})|\leq 6\sigma_{T-t}^{-4}\bar{B}$,
where $D_{k}$ and $D_{k\ell}^{2}$ denote the first-order and second-order partial derivatives with respect to the input coordinates, respectively.
Proof of Proposition 5.8.
By the simialr argument as Lemmas C.1, C.4, and C.5, we conclude the desired results. ∎
By a similar argument as Lemma D.8, we have the following approximation error bounds for low-dimensional Doob’s $h$-function $\bar{h}_{t}^{*}$ (5.7).
Lemma F.1 (Approximation error).
Suppose Assumptions 4 and 2 hold. Let $R\geq 1$, and let the hypothesis class ${\mathscr{H}}_{t}$ be defined as (5.8) with $L\leq C\log N$ and $S\leq N^{d^{*}}$, then
$\|h_{t}-\bar{h}_{t}^{*}\|_{L^{2}(p_{T-t})}^{2}$ $\leq C\frac{\bar{B}^{2}\log^{4}N}{\sigma_{T-t}^{8}N^{4}},$
$\|\nabla h_{t}-\nabla\bar{h}_{t}^{*}\|_{L^{2}(p_{T-t})}^{2}$ $\leq C\frac{\bar{B}^{2}\log^{2}N}{\sigma_{T-t}^{8}N^{2}}.$
provided that $R^{2}=(4d\mu_{t}^{2}+8\sigma_{t}^{2})\log N^{4}$, where $C$ is a constant only depending on $d^{*}$.
Theorem 5.9.
Suppose Assumptions 4 and 2 hold. Let $t\in(0,T)$. Set the hypothesis class ${\mathscr{H}}_{t}$ as
(F.3) ${\mathscr{H}}_{t}\coloneq\left\{h_{t}\in N(L,S):\begin{aligned} &\sup_{\mathbf{x}\in{\mathbb{R}}^{d}}h_{t}(\mathbf{x})\leq\bar{B},\penalty 10000\ \inf_{\mathbf{x}\in{\mathbb{R}}^{d}}h_{t}(\mathbf{x})\geq\underline{B},\\ &\max_{1\leq k\leq d}\sup_{\mathbf{x}\in{\mathbb{R}}^{d}}|D_{k}h_{t}(\mathbf{x})|\leq 2\sigma_{T-t}^{-2}\bar{B}\end{aligned}\right\},$
where $L={\mathcal{O}}(\log n)$ and $S={\mathcal{O}}(n^{\frac{d^{*}}{d^{*}+8}})$. Let $\widehat{h}_{t}^{\lambda}$ be the gradient-regularized empirical risk minimizer defined as (4.6), and let $h_{t}^{*}$ be the Doob’s $h$-function defined as (3.10). Then the following inequality holds:
${\mathbb{E}}\Big[\|\nabla\log\widehat{h}_{t}^{\lambda}-\nabla\log h_{t}^{*}\|_{L^{2}(p_{T-t})}^{2}\Big]\leq C\sigma_{T-t}^{-8}n^{-\frac{2}{d^{*}+8}}\log^{4}n,$
provided that the regularization parameter $\lambda$ is set as $\lambda={\mathcal{O}}(n^{-\frac{2}{d^{*}+8}})$, where $C$ is a constant depending only on $d^{*}$, $\bar{B}$, and $\underline{B}$.
Proof of Theorem 5.9.
Using the same arguments as the proof of Theorem 5.3 and applying Lemma F.1 completes the proof. ∎
Appendix G Auxilary Lemmas
Lemma G.1.
Suppose Assumption 1 holds. Let $\mathbf{X}_{t}\sim p_{t}$. Then for each $\xi>0$,
$\operatorname{Pr}\big\{\|\mathbf{X}_{t}\|_{2}\geq\xi\big\}\leq 2^{d+1}\exp\Big(-\frac{\xi^{2}}{4d\mu_{t}^{2}+8\sigma_{t}^{2}}\Big).$
Proof of Lemma G.1.
According to Assumption 1, we have
(G.1) ${\mathbb{E}}\Big[\exp\Big(\frac{\|\mu_{t}\mathbf{X}_{0}\|_{2}^{2}}{2d\mu_{t}^{2}}\Big)\Big]\leq 2.$
Let ${\bm{\varepsilon}}\sim N({\mathbf{0}},\mathbf{I}_{d})$. Then it follows that
${\mathbb{E}}\Big[\exp\Big(\frac{\|\sigma_{t}{\bm{\varepsilon}}\|_{2}^{2}}{4\sigma_{t}^{2}}\Big)\Big]$ $={\mathbb{E}}\Big[\exp\Big(\frac{\|{\bm{\varepsilon}}\|_{2}^{2}}{4}\Big)\Big]$
$=(2\pi)^{-\frac{d}{2}}\int\exp\Big(\frac{\|{\bm{\varepsilon}}\|_{2}^{2}}{4}\Big)\exp\Big(-\frac{\|{\bm{\varepsilon}}\|_{2}^{2}}{2}\Big)\,\mathrm{d}{\bm{\varepsilon}}$
(G.2) $=(2\pi)^{-\frac{d}{2}}\int\exp\Big(-\frac{\|{\bm{\varepsilon}}\|_{2}^{2}}{4}\Big)\,\mathrm{d}{\bm{\varepsilon}}\leq 2^{d}.$
Notice that $\mathbf{X}_{t}\stackrel{{\scriptstyle\,\mathrm{d}}}{{=}}\mu_{t}\mathbf{X}_{0}+\sigma_{t}{\bm{\varepsilon}}$, where $\mathbf{X}_{0}\sim p_{0}$ and ${\bm{\varepsilon}}\sim N({\mathbf{0}},\mathbf{I}_{d})$ are independent. Therefore,
${\mathbb{E}}\Big[\exp\Big(\frac{\|\mathbf{X}_{t}\|_{2}^{2}}{4d\mu_{t}^{2}+8\sigma_{t}^{2}}\Big)\Big]$ $={\mathbb{E}}\Big[\exp\Big(\frac{\|\mu_{t}\mathbf{X}_{0}+\sigma_{t}{\bm{\varepsilon}}\|_{2}^{2}}{4d\mu_{t}^{2}+8\sigma_{t}^{2}}\Big)\Big]$
$\leq{\mathbb{E}}\Big[\exp\Big(\frac{\|\mu_{t}\mathbf{X}_{0}\|_{2}^{2}}{2d\mu_{t}^{2}+4\sigma_{t}^{2}}+\frac{\|\sigma_{t}{\bm{\varepsilon}}\|_{2}^{2}}{2d\mu_{t}^{2}+4\sigma_{t}^{2}}\Big)\Big]$
$\leq{\mathbb{E}}\Big[\exp\Big(\frac{\|\mu_{t}\mathbf{X}_{0}\|_{2}^{2}}{2d\mu_{t}^{2}+4\sigma_{t}^{2}}\Big)\Big]{\mathbb{E}}\Big[\exp\Big(\frac{\|\sigma_{t}{\bm{\varepsilon}}\|_{2}^{2}}{2d\mu_{t}^{2}+4\sigma_{t}^{2}}\Big)\Big]$
(G.3) $\leq{\mathbb{E}}\Big[\exp\Big(\frac{\|\mu_{t}\mathbf{X}_{0}\|_{2}^{2}}{2d\mu_{t}^{2}}\Big)\Big]{\mathbb{E}}\Big[\exp\Big(\frac{\|\sigma_{t}{\bm{\varepsilon}}\|_{2}^{2}}{4\sigma_{t}^{2}}\Big)\Big]\leq 2^{d+1},$
where the the first inequality follows from Cauchy-Schwarz inequality, the second inequality holds from the independence of $\mathbf{X}_{0}$ and $\bm{\varepsilon}$}, and the last inequality is due to (G.1) and (G.2). Then we aim to bound the tail probability. For each $\xi>0$, we have
$\operatorname{Pr}\big\{\|\mathbf{X}_{t}\|_{2}\geq\xi\big\}$ $=\operatorname{Pr}\Big\{\frac{\|\mathbf{X}_{t}\|_{2}^{2}}{4d\mu_{t}^{2}+8\sigma_{t}^{2}}\geq\frac{\xi^{2}}{4d\mu_{t}^{2}+8\sigma_{t}^{2}}\Big\}$
$=\operatorname{Pr}\Big\{\exp\Big(\frac{\|\mathbf{X}_{t}\|_{2}^{2}}{4d\mu_{t}^{2}+8\sigma_{t}^{2}}\Big)\geq\exp\Big(\frac{\xi^{2}}{4d\mu_{t}^{2}+8\sigma_{t}^{2}}\Big)\Big\}$
$\leq\exp\Big(-\frac{\xi^{2}}{4d\mu_{t}^{2}+8\sigma_{t}^{2}}\Big){\mathbb{E}}\Big[\exp\Big(\frac{\|\mathbf{X}_{t}\|_{2}^{2}}{4d\mu_{t}^{2}+8\sigma_{t}^{2}}\Big)\Big]$
$\leq 2^{d+1}\exp\Big(-\frac{\xi^{2}}{4d\mu_{t}^{2}+8\sigma_{t}^{2}}\Big),$
where the first inequality invokes Markov’s inequality, and the last inequality is due to (G.3). This completes the proof. ∎
Corollary G.2.
Suppose Assumptions 1 and 2 hold. Let $\mathbf{Z}_{t}\sim q_{t}$. Then for each $\xi>0$,
$\operatorname{Pr}\big\{\|\mathbf{Z}_{t}\|_{2}\geq\xi\big\}\leq 2^{d+1}\exp\Big(-\frac{\xi^{2}}{4d\mu_{t}^{2}+8\sigma_{t}^{2}}\Big).$
Proof of Corollary G.2.
Under Assumptions 1 and 2, $\mathrm{supp}(q_{0})=\mathrm{supp}(p_{0})$. The same argument as Lemma G.1 completes the proof. ∎
Lemma G.3.
Suppose Assumptions 1 and 2 hold. Let $\mathbf{Z}_{t}\sim q_{t}$. Then for each $\xi>0$,
${\mathbb{E}}\big[\|\mathbf{Z}_{t}\|_{2}^{4}\big]\lesssim d^{2}.$
Proof of Lemma G.3.
Let ${\bm{\varepsilon}}\sim N({\mathbf{0}},\mathbf{I}_{d})$. It is straightforward that
${\mathbb{E}}\big[\|{\bm{\varepsilon}}\|_{2}^{4}\big]=4\varGamma\Big(\frac{d+4}{2}\Big)\varGamma\Big(\frac{d}{2}\Big)\leq(d+4)^{2}.$
Since $\mathbf{Z}_{t}\stackrel{{\scriptstyle\,\mathrm{d}}}{{=}}\mu_{t}\mathbf{Z}_{0}+\sigma_{t}{\bm{\varepsilon}}$ with $\mathbf{Z}_{0}\sim q_{0}$ independent of $\bm{\varepsilon}$}, it follows from the triangular inequality that
${\mathbb{E}}\big[\|\mathbf{Z}_{t}\|_{2}^{4}\big]$ $\leq 8\mu_{t}^{4}{\mathbb{E}}\big[\|\mathbf{Z}_{0}\|_{2}^{4}\big]+8\sigma_{t}^{4}{\mathbb{E}}\big[\|{\bm{\varepsilon}}\|_{2}^{4}\big]$
$\leq 8(d^{2}+(d+4)^{2}),$
where we used the fact that $\mu_{t},\sigma_{t}\leq 1$, and $\mathrm{supp}(q_{0})=\mathrm{supp}(p_{0})$ under Assumptions 1 and 2. This completes the proof. ∎